### v4

# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

Abstract: The embrace of globalization and protectionism among economies has ebbed and flowed over the past few decades. These fluctuations call for quantitative analytics to help countries improve their trade policies. Changing attitudes toward globalization also imply that the best trade policies vary over time and are country-specific. We argue that the imports and exports of all economies constitute a counterbalanced network where conflict and cooperation are two sides of the same coin. Quantitative competitiveness is then formulated for each country using a network counterbalance equilibrium. A country could improve its relative strength in the network by embracing globalization, protectionism, trade collaboration, or conflict. This paper presents some necessary conditions for globalization and trade wars, evaluates their side effects, derives national bargaining powers, identifies appropriate targets for conflict or collaboration, and recommends fair resolutions for trade conflicts. Data and events from 2000 to 2019 support these conditions.
Keywords:  globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma
JEL Code: C71, C78, F11, F13, F15, O24
1 Introduction
De-globalization and a trade war are two levels of economic conflict; the latter directly involves only two trade partners, whereas the former represents one country’s disintegration from the global economy. Nevertheless, both could adversely affect global economic activity. This paper applies a unified game-theoretic approach to formulate simple but decisive rules to mitigate either type of conflict. We focus on import and export activities, ignoring political, territorial, ideological, military, and cultural conflicts. However, geopolitical data have already been incorporated into trade behaviors and volumes. From this perspective, a country’s primary purpose in engaging in trade wars or globalization is to improve its competitiveness and trade balance in the trade system. Also, the engagement affects the future geopolitical landscape.
To analytically examine trade wars and globalization, we face several daunting challenges. First, there are always conflicts and cooperation in international trade. While the trade of goods and services results from bilateral collaboration, a country’s growing strength could weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study could be too impotent to capture the complexity of trade relationships. The second challenge is adequately constructing a high-level description of the interactive trade system because exports and imports have already reflected the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility or social welfare functions in this conflicting and cooperative context. Next, policy variables, such as imports and exports as production’s percentages, should be linked to these economic objectives. So, countries could apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, it is fruitful to bring data into the analytics to explain past events and find optimal solutions for the near future. The solutions could identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization. In the literature, studies on globalization often capitalize on econometric analysis, which uses historical data to estimate unknown parameters. However, the estimated effects could hardly catch up with the recent changes or seize both cooperation and non-cooperation.
We measure the relative strengths of the countries in the network by the solution to the counterbalance equilibrium, defined by Hu and Shapley, (2003), and use it as an objective in our analysis. The solution, called authority distribution, integrates both cooperation and competition inside the global economic system. The data-driven distribution is a social welfare function and an endogenously weighted average of utility functions. Moreover, it measures the competitive advantages of the countries and their long-term influences on the whole system. Also, this competitiveness brings extra welfare beyond the economy of scale and comparative advantage. The additional advantage, called the Matthew effect, is another driving force for competition after comparative advantages.
This paper attempts to provide strategic solutions for policymakers and ignores how an individual firm or economic agent improves firm’s competitiveness through research and development. Maximizing national competitive strength is a purview of government policymakers rather than trade companies, which often capitalize on comparative advantages. A country is not merely an aggregate of firms, and the interests of trade companies only sometimes align with those of policymakers, who also consider the side effects of trade, such as the shrinkage of domestic labor markets and the overseas shift of industries. So, our studies neglect the firm-based micro foundation and maximize a social welfare function, pivoting domestic production decision-making. However, the competitiveness measure is also a weighted average of comparative advantages that considerably reflect the profits of domestic firms engaging in international trade and the welfare of consumers.
When starting a trade war against a trade partner, a country raises tariffs or creates other barriers to increase competitiveness. However, the result depends on the counterpart’s reaction. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in a conflict or cooperation. On the other hand, the reaction may also benchmark their production ratio to avoid trade deficits. Using the threshold and the benchmark ratios, we derive a Nash bargaining solution for countries with heterogeneous preferences between competitiveness and trade balance, and the solution generates consistent national bargaining power. It provides a solution to resolve bilateral trade frictions.
A country with no specific targets to confront could choose between further globalization or protectionism. By minor perturbations in its overall exports, we calibrate the effect on its competitiveness from the counterbalance equilibrium. The sign of the effect tells if the country should go for more globalization or not. There are many debates over economic globalization in the literature, and Stiglitz (2002, 2017) documented their detailed viewpoints. Different viewpoints, assumptions, and considerations could reach opposite conclusions about globalization for a country. Our data-driven approach ignores arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time. By analyzing the ComTrade data of the United Nations (2021) for the years 2000 through 2019, we find that protectionism could have been a better choice for the USA and the United Kingdom but a worse option for China, Germany, Japan, and Russia if they targeted competitiveness and trade balances only. Data from 2017 also show that trade wars could be an option for the USA to maintain its competitiveness. However, adverse side effects on third parties and allies were also tangible. According to our theorem of impossibility trilemma, it has to reduce imports if it aims to lower its national debt by earning trade surpluses and also preserve its competitiveness.
In the following exposition, we first introduce the competitiveness objective in Section 2. Then, Section 3 formulates a necessary condition for a country to launch a trade war or form a trade partnership to optimize its objectives. After that, a bargaining solution is derived in Section 4 to resolve conflicts or share profits. Next, Section 5 discusses a similar condition for globalization or protectionism. Then, in an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, we discuss several extensions and limitations of this framework in Section 7. Our exposition is self-contained and the proofs are in the Appendix.
2 Competitive Advantage in the International Trade System
Based on Adam Smith’s economy of scale and David Ricardo’s comparative advantage, if all production factors were free of flow, then, hopefully, global production would be maximized. However, not all countries share the same well-off status; some succeed over some time, whereas others fail during the same period. Even the most powerful economies have failed in human history, and there needs to be more mechanisms to enforce the fairness of profit sharing in international trade. Thus, a country has to compete for its share and, at the same time, cooperate with others to maximize global production. Both competition and cooperation are essential in international trade.
Domestic production is paramount for policymakers, whereas profit maximization is for firms or producers. Production with zero or slightly negative profit means nothing to firm owners but may still be an option for the government to uplift employment. Production maximization, for example, results in less unemployment, while labor may be cut or resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
We regard the international trade system as a game on a network in which countries act as nodes and trades as directional edges. The trade volumes flow along the edges, and imports and exports move in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or antiglobalization are often framed to shield domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of the trade counterparts.
This section introduces a new type of competitiveness as an objective and social welfare function for the countries in the network. Competitiveness is a relative measure, and we use a constant-sum game to derive it. Absolute production and total trade volume, thus, may give limited information on the economy’s competitive position. The game mocks the in-and-out flows over the network; without data change, these steady flows reach a static equilibrium due to long-run accumulation. The equilibrium is the competitiveness vector, which describes the utilities of the countries’ trade activities. We detail the steps toward the data-based utility and social welfare functions, which are unfamiliar in the trade literature.
2.1 Long-Run Influence in International Trade System
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\not=j$, let $P_{ij}$ be the fraction of country $i$’s production, which exports to country $j$. The exports include all services and goods, such as final goods and intermediate goods, both of which may also contain parts made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are attributable directly to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\not=i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $j\in\mathcal{N}$. We place all the fractions $P_{ij}$ into a $n\times n$ stochastic matrix $P=[P_{ij}]$.
We look for a utility function which determines country $i$’s production or expected production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
$v_{i}(S)\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{j\in S}P_{ij},$ (1)
together with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. So, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. It is normalized since $v_{i}(\mathcal{N})=1$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game and its Shapley value is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$ (Shapley, 1953). Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern (1953) utility function, i.e., $P_{ij}$ is $j$’s utility in the game of $(\mathcal{N},v_{i})$. So, countries have different preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. However, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse judgments of every country in $\mathcal{N}$. It rates the set of countries by $\mathcal{N}$ as a whole.
The element $P_{ij}$ is a type of pro rata power when country $i$ makes its production. The multilinear extension of $v_{i}$ is defined by
$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
where $x_{j}\in[0,1]$ is the probability with which $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also means $j$’s progress in completing its imports from $i$. In addition, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For an annual production, for example, each country finishes $1/12=8.33\%$ of its imports from $i$ in any month. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s influence or power when country $i$ completes its production from 00 to $1$. According to Owen, (1972), the integral equals the Shapley value for the coalitional game, i.e.,
$\int_{0}^{1}\bigtriangledown\tilde{v}_{i}(x,\cdots,x)\mathrm{d}x=(P_{i1},P_{i2},\cdots,P_{in})^{\top}$
where $\top$” denotes vector or matrix transpose.
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them together. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. They are primarily payments for past, current, or future exports. Trade is about exchanging goods and services; as the medium of exchange, monetary flow makes transactions possible in modern international trade. Nevertheless, payments alone could not fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Also, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. Besides, a country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and low-risky investment instruments, so foreign governments or companies must earn extra U.S. dollars from their trades to put into their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for a fictitious example, the chip and design software of the phone could come from the USA and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
in which $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs all countries in $\mathcal{N}$.
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. As a result, $P^{t}$ converges only if $t$ is large enough for all the direct impacts to travel from the top of global economic chains down to the very bottom. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.
2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, called authority distribution in Hu and Shapley, (2003), also solves the following counterbalance equilibrium
$\pi=\pi P$ (3)
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and $\pi_{i}\geq 0$. There exist many counterbalanced systems in the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties — even the most competitive economy needs both cooperation and competition. Between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Though unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). An econometric analysis, for example, could hardly capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From the noncooperative aspect, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase of $\pi_{j}$ may imply a decrease of $\pi_{i}$. So, in theory, there are $n(n-1)/2$ potential trade conflicts, either small or large, in the trade system. Most of them are minor trade disagreements or disputes, so the “war” metaphor is hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Also, the increase of $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initialized by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively.
From the cooperative aspect, country $i$ should help $j$ improve $\pi_{j}$ whenever $P_{ji}>0$ because $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}P_{ji}$ (4)
from Eq. (3). Also, Eq. (3) implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ in which
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}\left(P^{t}\right)_{ji}$ (5)
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. So, Eq. (5) further implies that $i$ should help all other countries in the trade system, including the poorest and the least competitive ones, to enhance their $\pi_{j}$. Thus, the double-edged counterbalance implies that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is how much it should assist or contest $j$ without sacrificing its national interests and even increasing $\pi_{i}$.
[…]
Figure I: Dynamics of authority flow for country $i$ and $j$.
We can physically visualize $\pi_{i}$ as a container or pool in the dynamics of counterbalance equilibrium, in which comparative advantages are the primary driving force for the flows. On the one hand, in the inflow equation Eq. (4), $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, other things being equal. It also derives more authority from countries on which it has a sizable direct influence (i.e., $P_{ji}$ is large), other things being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$ and so, authority flows from $j$ to $i$ alongside the same column. On the other hand, $i$ also contributes to other countries as seen in the following outflow equation
$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\not=i}\pi_{k}P_{kj}.$
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, other things being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries have a break-even point when
$\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ takes more from or gives more to country $j$, respectively. For country $i$, a simple implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\not=j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\not=j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ still maintains an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
2.3 $\pi_{i}$ as Competitive Advantage
Country $i$’s competitive advantage refers to its ability to do better than its competitors in the trade system. For any product exported from country $i$ to $j$, $i$’s competitors include all countries other than itself. So, its competitive advantage extends from bilateral country relations to a relationship with the whole system — it is the ability to beat the best of all other countries in the $j$th market for any $j\in\mathcal{N}$. On the contrary, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over a specific trade partner on particular services or products; it is a bilateral relationship. According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, and a low cost of capital. These factors, imported from the outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker — in Eq. (4), $\pi_{j}P_{ji}$ contains $j$’s return of competitiveness to $\pi_{i}$. In the above smartphone example, the competitive factors include the chips and patents from the USA, the cheap labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means a solid competitiveness to occupy the $j$th market and win a lion share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$ because domestic products also compete with country $j$’s imports when the consumers make their consumption choices, and they have no home bias. We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. The weights are not exogenously assigned but endogenously determined. Also, more weight should be placed on a more competitive reference market than on a less competitive one because it is more challenging to occupy a competitive market than a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. So, both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
Theorem 1
If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
The national competitiveness $\pi_{i}$ is not the sum of the individual competitiveness of the country’s domestic firms, though they are positively associated. In the endogenous weighting, worldwide consumers unconsciously participate in the evaluation of competitiveness, and all products are involved. Essentially, countries unknowingly judge countries themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is, to a great extent, endogenously determined by the distribution of firm productivity in individual countries. Nevertheless, government policies can affect the distribution without significantly mitigating the welfare of international trade by switching imports or exports from one country to another. It can also enhance firm productivity through tax exemption, subsidies, and heavy tariffs on foreign products.
Also, $\pi$ is a von Neumann-Morgenstern Bergson social welfare function (e.g., Harsanyi, 1955; Hammond, 1987) since it is a positively weighted sum of von Neumann-Morgenstern individual utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$; the weights are $\pi$ itself. Thus, the weighted utility function $\pi$ is not determined by how much production factors are consumed but by how much the consumers prefer the products, judged regardless of national interests or preferences by individual countries. Pareto optimality is satisfied when weighing these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant.
However, the welfare function $\pi$ departs significantly from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors but implied from the counterbalance equilibrium because indirect exports also act upon the final products or services. Any exported goods and services from one country to another have different utilities for the exporting and importing countries. Therefore, each country has its unique utility function, and a single utility function or linear regression cannot capture the diversity of all countries. Also, we use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages these factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models. Thus, no unknown parameters are to be estimated, and no residuals exist to differentiate the data from a parametric model. According to Barney and Felin, (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Because of broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network. Also, unjustified assumptions and overly mathematization could mislead the conclusions and policy implications (e.g., Romer, 2015).
The index $\pi_{i}$ contains an inalienable part besides absorbing spillover effects through $P$. There are $t-1$ occurrences of $P_{ii}$ in Eq. (2), which largely accounts for $\pi_{i}P_{ii}$ in Eq. (4) as $t\to\infty$. A global value chain generally takes only finite steps, but it also contains indefinite domestic loops when we trace all the added values of the final products. In the smartphone example, the manufacturing labor’s education, skills, and experience are ultimately achieved domestically over many years, but all somehow contribute to the final product. Also, many USA workers develop the patents and design software domestically for an extended period. These inner loops affect other countries but not vice versa; this part of competitiveness never flows in from external. Thus, social infrastructure and political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that help do business all contribute to national competitiveness (e.g., Delgado et al., 2012). Also, raw oil fields, natural gas reserves, rare minerals, and patented technologies do not come from external sources and remain with $\pi_{i}P_{ii}$ as the nation’s core competitive factors. Refraining exports of rare earth elements or licensed foreign usage of patented technologies, for example, could destroy one or more global value chains.
For a specific global value chain, on the other end of the inalienable parts is the design and assembly of the final product. Product leadership bridges the market, the consumers, strategic direction, technology innovation, sales, and production factors together; this leadership is another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability curtail the risk of costs.
Finally, Eq. (2) implies a solid positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. Nations sell goods that they have in abundance and buy goods that are in short supply domestically. With great abundance and shortage, country $i$ could have vast nonzero elements in the $i$th column and row of the matrix $P$, resulting in astronomical nonzero items summed in Eq. (2). A large economy also produces a high differentiation of goods and services, which bear location advantage and minimal transaction costs for domestic consumers. Free of tariffs and overseas shipping, they also contain a price advantage. For these reasons, $P_{ii}$ is often the largest in the $i$th row of $P$, with or without the home bias. Therefore, a massive economy with a comprehensively integrated industrial system (such as China and the USA) is likely, but not necessary, to have a large value in $\pi$. A country with a single industrial structure is vulnerable to brutal competition, and its $\pi_{i}$ is fragile and volatile when battling with the USA for competitiveness.
Therefore, we use $\pi_{i}$ as an objective for country $i$ when making trade policies. It quantifies the economy’s depth and interplays with breadth; it also incorporates cooperation and non-cooperation with others. It is endogenously extracted from the comparative advantages of all products and services. As each country has a unique location, natural and human resources, and economic systems, a single utility function, which presumes one stereotype for all economies, may be inadequate to capture this diversity. In our study, we recognize that each country, as is its trade stance with partners, is different. Each adds inclusiveness value to the global economy, of which trade between nations is a mainstay.
2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness, i.e.,
$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\not=i}\pi_{j}P_{ji},$ (7)
in which the proportion $P_{ii}$ is generally larger than .7.7.7.7. Thus, $\pi$ has the Matthew effect of accumulated advantage, sometimes summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies have substantial competitiveness compared to less powerful ones, revealing a nonlinear relation between $\pi_{i}$ and $i$’s production $g_{i}$. For example, the USA’s $\pi_{i}$ could be much higher than Japan’s, compared to their production ratio. The amplification also comes from the USA’s relative advantages over other countries, compared to Japan’s. In Eq. (7) with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$,
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\not=\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\not=\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ comes from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, in general, for other countries $j$. Alternatively, the Matthew effect is largely explained by the preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among the countries according to how competitive they already are.
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. So, it has a quadratic growth when $\tilde{g}_{i}$ soars linearly and in a linear regression,
$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i}$ (8)
for some positive $c_{1}$ and $c_{2}$. Therefore, when dealing with a rapidly thriving country $j$, $i$ should mitigate its conflicts with $j$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say,
$g_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i},\quad\mathrm{for}\ \mathrm{some}\ c_{4}>0,$
then $g_{i}$ would be accelerated quadratically, giving rise to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $g_{i}$ would grow exponentially to their maximum capacity. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
Granting extra welfare to competitive economies is another major force that drives trade flows over the network. So, an economic superpower would make all efforts to maintain its superpower status when facing challenges from emerging competitors; maintaining its advantage in $\pi$ avoids the domino effect, the opposite of the Matthew effect. As Copeland (2014) says, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Thus, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. Accordingly, an emerging superpower’s aspiration to assume its rightful place as a superpower would face serious challenges. World War I and II were fiery combats between existing and emerging superpowers, resetting new world orders.
The welfare of competitive advantage leads to an agglomeration effect, which clusters all countries around a handful of economic superpowers. For example, country $i$ considers certain imports from either country $j$ or $k$ where $g_{j}$ is significantly larger than $g_{k}$, other things being equal. If $i$ selects $j$, by Eq. (4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
$\pi_{j}\Delta P_{ji}=\pi_{j}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j}{g_{j}}=\frac{\pi_{j}}{g_{j}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j)$
where $\Delta P_{ji}$ is the change on $P_{ji}$ due to the imports. The approximation ignores the ripple effect from the shock. If $i$ selects $k$, then the competitiveness country $i$ derives from $k$ in the imports is approximately
$\pi_{k}\Delta P_{ki}=\pi_{k}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k}{g_{k}}=\frac{\pi_{k}}{g_{k}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k).$
Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$ and, so, country $i$ would likely choose $j$, other things being equal. Otherwise, $k$ could counter the offer by lowering its price of exports to $(\pi_{k}g_{j})/(\pi_{j}g_{k})$ of $j$’s price. In either case, a small country (such as $k$) tends to fear the emergence of a neighboring superpower (such as $j$) that shares similar location advantage and resource endowment.
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It is a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal such that domestic and foreign goods can compete for the domestic market, and the eigenvector $\pi$ measures the competitiveness. The same domestic products are used domestically or overseas. Hu, (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; so, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Besides, $P$ normalizes the scales of economies such that each row has a unit sum. As a result, $P$ has the largest row eigenvalue $1$, which is common to all stochastic matrices, and its unit-sum eigenvector $\pi$, which is specific to $P$.
3 Bilateral Trade War for Competitiveness
Expecting a positive change to $\pi_{i}$, country $i$ considers a trade war against its trade partner $j$. Gaming on the matrix $P$, $i$’s first-move action could change the element $P_{ji}$ — the portion of $j$’s production exported to $i$. Raising tariffs or decreasing import quotas upon $j$, for example, reduces imports from country $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume that the bilateral conflict involves only the two counterparties without directly sacrificing or benefiting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. Nevertheless, the change on $P_{ji}$ or $P_{ij}$ eventually ramifies over the global value chains in multiple steps, and all third parties get involved indirectly. The simplicity could be far from realistic in the 2018 China-USA trade war as the Biden administration also urged its allies to ban Chinese companies from buying advanced chips and chip-making equipment (e.g., New York Times, 2022).
As a reprisal, country $j$ could react by decreasing $P_{ij}$ or $P_{ji}$ sooner or later. Without the reprisal, the war dies out at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages, but it could also bring in further sanctions from $i$ when $j$ shows a sign of weakness. Overreaction to the hostile change on $P_{ji}$ or $P_{ij}$ may also severely hurt $\pi_{j}$. After the Trump administration began setting tariffs and other trade barriers on China in January 2018, for example, China responded in April 2018 by imposing tariffs on 128 products imported from America (cf., Washington Post, 2018). Say, in retaliation for a $\Delta P_{ji}$ change on $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s $\Delta P_{ij}$ change on $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. So, $\lambda_{ij}=1/\lambda_{ji}$ in the veil of any first-mover advantage. We assume that $\lambda_{ji}>0$ to align the actions from both countries. Thus, ignoring the temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact on $\pi$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; so the reduced exports will be consumed domestically, and the modified $P$ remains as a stochastic matrix. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into a $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
The coefficient $\lambda_{ji}$ often depends on many exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. In general, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. For a specific context with a set of fair assumptions, however, many researchers in the literature have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). In the following exposition, we first assume the coefficient is an exogenous constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
We introduce a few notations for the next five theorems and their corollaries. Let $I_{n}$ be the $n\times n$ identity matrix and let $\pi_{{}_{-i}}$ be the transpose of $\pi$ with $\pi_{i}$ removed. The square matrix $Z_{i}$ is the transpose of $P$ with its $i$th row and $i$th column removed. Also, the column vector $\alpha_{i}$ extracts the $i$th row from $P$ and then drops its $i$th element. As usual, $\vec{1}_{n}$ and $\vec{0}_{n}$ are the $n\times 1$ column vectors with all ones and zeros, respectively. For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has one for its $j$th element and zeros elsewhere. Finally, the vector $\gamma_{ji}$ takes the values of $e_{j}$ and then removes its $i$th element. So, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
Theorem 2
With the above setting and notations, for any $j\not=i$,
$\hskip 8.5359pt\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}},$ (9)
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}},$ (10)
and
$\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (11)
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain
$\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
Using Theorem 2, we have the following Corollary 1, already implicated in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
Corollary 1
When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
Besides, since
$(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 2. In particular, when $\lambda_{ji}=0$ in Eq. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Also, Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
Corollary 2
For any $i\not=j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
By varying $j$ in $\mathcal{N}$, we could apply Eq. (9) to identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change on $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$ regardless of the sign of $\mathrm{d}P_{ji}$.
In search for the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
$\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}=\frac{\mathrm{d}\pi_{i}/\pi_{i}}{\mathrm{d}P_{ji}/P_{ji}}=\frac{P_{ji}}{\pi_{i}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}.$ (12)
[…]
$\mathop{\mathrm{argmax}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
as $i$’s best trade partner and
$\mathop{\mathrm{argmin}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
as $i$’s best competitor. In Eq. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his best competitor’s best competitor, according to Corollary 2. Thus, $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
Eq. (11) is the side effects on all other countries due to $i$’s change on $P_{ji}$, including $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. With the perturbation on $P_{ji}$, some countries gain, and others lose their competitiveness; the gain or loss may be even more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Though $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not the same magnitude, so one movement could be significant, whereas the other remains insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ unlikely exercises these extreme actions because $\pi_{i}$’s gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been set up to tame trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
4 A Nash Bargaining Solution for Reprisal Coefficients $\Lambda$
When other factors also play a role, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to start a conflict action against $j$. Among them, the most primary are trade deficits. This section studies a bilateral bargaining solution for $\lambda_{ji}$ when trade balance is also an objective. If small economies are willing to dodge direct competition with large ones, the vast majority of the $n(n-1)/2$ conflicts are resolved or mitigated by the bargaining solution.
4.1 Trade Balance
For a given period, country $i$’s trade surplus from $j$ means $j$’s trade deficit of the same magnitude from $i$, and vice versa. With a trade surplus, job opportunities are created in the country, decreasing its unemployment rate and expanding its economy of scale. It also improves the economy’s creditworthiness because it can use the surplus to pay down its foreign debt, if any. A country’s surplus could result from its increasing competitive advantage in tradable goods and services when it beats the competitors in the counterpart’s market. Opinions on the impacts of trade deficits, either bilateral or overall, are diverse and sometimes controversial, although huge deficits can undermine the sustainability of the economy. In the short run, however, trade deficits help nations avoid shortages of goods and mitigate related economic problems, such as inflation and poverty. Implications of trade deficits also depend on their impacts on national security and how they are financed. No nation would completely ignore trade deficits or competitiveness; otherwise, an one-sided emphasis will not last long.
There is a benchmark choice of $\lambda_{ji}$ for net trade balances: $\lambda_{ji}=g_{j}/g_{i}$, the ratio between their productions. In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$ while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$. If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these two countries, as claimed by Theorem 3. This condition is for instant balance change. In aggregate value over a specific time period, there is no trade surplus and deficit between these two countries if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$ because $g_{i}P_{ij}$ is the exports from $i$ to $j$ and $g_{j}P_{ji}$ the exports from $j$ to $i$. For $i$ to have a trade surplus with $j$, we need $g_{i}\lambda_{ji}\Delta P_{ji}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain this trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but hurts $\pi_{i}$ — with overly excessive exports from $i$ to $j$ in the $i$th row of $P$, country $i$ transfers additional production power to $j$, reducing $\pi_{i}$. So $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports even though $i$ loses fewer. This could eventually adversely affect the productions of both countries in the future so that most countries would avoid mutually damaging stalemates. In general, the parameter $\lambda_{ji}$ is determined by both players. However, the sign and size of $\Delta P_{ji}$ are often decided by one only, especially if it has overwhelming power over its counterpart.
Theorem 3
$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
To ease the trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change on $\Delta P_{ji}$. On the other hand, if they look for a balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ removes any trade deficits in the future. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for a long-run rivalry due to $\pi$’s involvement in all countries and all industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and relieve relevant problems such as unemployment and national debt. According to Pozsar, (2022), for example, the 2018 China-USA trade war started in this way: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but built up massive national debt, accumulated through wars on terrors, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China wanted to build a global 5G network and make cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is a proper choice for $\lambda_{ji}$.
We study a general case when a superpower country $i$ faces imminent challenges from an emerging economy $j$. By the Matthew effect, likely $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This can happen only when $\Delta P_{ij}<0$ when both countries lose comparative advantages. Figure II illustrates this trilemma, in which up to two policy objectives are possible for country $i$. If it adopts a position on edge $b$, for example, then it upturns comparative advantage and earns a trade surplus, but gives up competitiveness unless it absorbs competitiveness from other countries. Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned with each other and also do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas $j$’s at another one. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries enjoy escalating economy of scale and diminishing marginal cost of production. According to the proof of Theorem 4, country $j$ could achieve these three objectives listed in the theorem. This seems unfair to country $i$. However, it is conditional on the Matthew effect, which is somewhat “biased” toward $\pi_{i}$. Therefore, Theorem 4 is indeed a redress to the effect.
Theorem 4 (Impossibility Trilemma)
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not achieve all the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
Figure II: Impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
4.2 The Bargaining Solution
We consider a compromise solution between trade balance and competitiveness because a small change $\Delta P_{ji}$ would hurt or benefit both countries. In the trade system, some countries prefer competitiveness while others prefer improved trade balance (either reducing deficit or increasing surplus). Their preferences affect the choice of $\lambda_{ji}$, which holds $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If the countries are indifferent between competitiveness and trade balance, then the middle point seems to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this middle-point solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Otherwise, whenever countries $i$ and $j$ have opposite priorities between enhancing competitiveness and improving trade balance, we have a bargaining problem and apply Nash, (1950) bargaining to find a cooperative solution, which leverages between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ and also binds the condition $\lambda_{ji}\lambda_{ij}=1$.
Let
$p\overset{\mathrm{def}}{=\joinrel=}\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}\quad\mathrm{and}\quad q\overset{\mathrm{def}}{=\joinrel=}\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}$ (15)
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. Without $i$’s bargaining, $\lambda_{ji}$ can secure $p$ and $q$ by Corollary 1 or Theorem 3. Similarly, without $j$’s bargaining, $\lambda_{ij}$ can secure $1/p$ and $1/q$ because
$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}$
and
$\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
We have two scenarios when country $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. So $u_{j}(x)=x$ is an utility function for $j$ because $u_{j}(q)=q>p=u_{j}(p)$. Similarly, $u_{i}(x)=x$ is an utility function for $i$ since $u_{i}(1/p)=1/p>1/q=u_{i}(1/q)$. These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation. By Eq. (15), $p$ and $1/q$ are their respective status quo utilities, obtained if one decides not to bargain with the other player. Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (16)
[…]
Figure III: The Nash bargaining solution $\lambda_{ji}^{*}$ for Eq. (16) and (17), respectively.
In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$. Their utility functions are $u_{j}(x)=q-x$ and $u_{i}(x)=1/p-x$, respectively, and the status quo is 00 for both. The maximization problem becomes
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(q-\lambda_{ji}\right)\left(\frac{1}{p}-\lambda_{ij}\right).$ (17)
The right plot of Figure III illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$ which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$. In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (17) by
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (18)
in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively. To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss. The final outcome is Eq. (18). Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq. (18) and, so, in Eq. (17).
Theorem 5 answers the maximization problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is the same for both preference scenarios, so it also works when country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. In Eq. (16), (17) and (18), both countries have a symmetric role and the same prior bargaining power. This is an axiom of the Nash bargaining. The solution remains unchanged under any positive affine transformation of the utility functions, according to another axiom of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. If $g_{j}<g_{i}$, for example, then $\pi_{j}/\pi_{i}<g_{j}/g_{i}$ is more likely to occur than $\pi_{j}/\pi_{i}\geq g_{j}/g_{i}$, according to the effect. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation fills the gap between these two ratios without changing the likelihood.
Theorem 5
The Nash bargaining solution to Eq. (16), (17), or (18) is
$\left\{\begin{array}[]{rcl}\lambda_{ji}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},\hskip 14.22636pt\mathrm{and}\\ \lambda_{ij}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}\end{array}\right.$
for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
The solution of $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$, in general, relies on the choice of status quo points. In Eq. (16), for example, both players could intensify or weaken their status quo positions. If the intensification or weakening is symmetrical with both players, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 3. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
Corollary 3
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$ in Eq. (16), (17), or (18), the solution is still $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
In summary, when country $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{cl}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}&\mathrm{otherwise.}\end{array}\right.$ (19)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb. Country $j$ should avoid competing for a higher $\pi_{j}$ with a significantly larger-sized economy $i$, resulting in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, no matter trade surplus or deficit. This was probably China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$. As a result, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, (1968)’s power transition theory).
Finally, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ in Eq. (19) are also Nash bargaining solutions for $\lambda_{ji}$ because
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (20)
and
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (21)
In Eq. (20) and (21), both countries battle for better competitiveness and net trade balance, respectively. At the solution points, the objective functions are zero and no country gets extra welfare, already implied in Corollary 1 and Theorem 3.
4.3 Posterior Bargaining Power
In a real bargaining context, non-economic factors (such as military power, population, culture, tradition, legislation, and location) and alternative optimal criteria could considerably compromise the symmetry axiom and hence deviate the bargaining solutions substantially different from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$. Thus, the countries likely have unequal prior bargaining power in Theorem 5, and we introduce a posterior bargaining power that updates the prior one.
To extend the bilateral Nash solution in Theorem 5 to multilateral, we define
$\beta_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{\sqrt{\pi_{i}g_{i}}}{\sum\limits_{j\in\mathcal{N}}\sqrt{\pi_{j}g_{j}}}$
as country $i$’s posterior bargaining power or national capability in the trade system. It is normalized such that all bargaining power sums to one and $\beta_{i}$ can be compared over time. Also, it measures bargaining power in multinational negotiations. When allocating votes in forming or reforming a multinational alliance, for example, using the power could avoid any bilateral dialogs and discussions.
Clearly, $\lambda^{*}_{ij}=\beta_{i}/\beta_{j}$ for any $i,j\in\mathcal{N}$. The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ construes a linear ordering for all countries and holds a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of that of $i$ over $j$ and that of $j$ over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. So, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when the companies in country $i$ rely on inputs from its partner countries. The more the companies depend on the partners, the less the country’s power to coerce concessions from the partners. From the consumption aspect, the bargaining power increases when the country consumes more imported goods and services by influencing the production of its counterparts. Between the determinants, $g_{i}$ is the country’s breadth of economy, and $\pi_{i}$ is predominantly the depth of the economy since the matrix $P$ in Eq. (3) already normalizes the production volumes. However, the positive correlation between them implies a risk of domino effect in which a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, which incurs further reduction of $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop of $g_{i}$.
5 Economic Globalization
Globalization refers to removing the barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the United Kingdom’s Brexit in 2020 was a clear rejection of globalization. On the contrary, a protectionism measure could raise tariffs on all imported goods, set import quotas for all other countries, and enact stricter government regulations on all imports. In 2018, for example, the Trump administration imposed punishing tariffs on all steel and aluminum imports from all other countries (e.g., Washington Post, 2018).
Globalization creates winners and losers within a nation but it does not automatically compensate the losers. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. Again, the objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments on the whole matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, say, with a small change $\Delta P_{ii}$ on $P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we make the following assumption: whenever $P$ has a change in the $i$th column, the other columns change proportionally to ensure each row still has a unit sum. Thus, the change $\Delta P_{ii}$ on $P_{ii}$ provokes the adjustments $\Delta P$ on $P$ in three levels of fallout. First, for any $j\not=i$, $P_{ij}$’s proportional response to $\Delta P_{ii}$ is
$\Delta P_{ij}=-\frac{P_{ij}}{1-P_{ii}}\Delta P_{ii},$
and the $i$th row of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{P_{i1}}{1-P_{ii}},\cdots,-\frac{P_{i,i-1}}{1-P_{ii}},1,-\frac{P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{P_{in}}{1-P_{ii}}\right),$
which has a zero sum and the adjustments are proportional to the $i$th row of $P$. Secondly, the change $\Delta P_{ij}$ causes
$\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}=-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}$
due to country $j$’s reprisal reaction. Therefore, the $i$th column of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{\lambda_{i1}P_{i1}}{1-P_{ii}},\cdots,-\frac{\lambda_{i,i-1}P_{i,i-1}}{1-P_{ii}},1,-\frac{\lambda_{i,i+1}P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{\lambda_{in}P_{in}}{1-P_{ii}}\right)^{\top}.$
Lastly, for any $t\not=i$, $P_{jt}$’s proportional response to $\Delta P_{ji}$ is
$\Delta P_{jt}=-\frac{P_{jt}}{1-P_{ji}}\Delta P_{ji}=\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}.$
The sum of the $j$th row in $\Delta P$ is
$-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}+\sum\limits_{t\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}=\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\left[-1+\sum\limits_{t\not=i}\frac{P_{jt}}{1-P_{ji}}\right]\Delta P_{ii}=0.$
We define the matrix $M$ by
$\left[\begin{array}[]{ccccccc}\frac{\lambda_{i1}P_{i1}P_{11}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1,i-1}}{1-P_{1i}}&-\lambda_{i1}P_{i1}&\frac{\lambda_{i1}P_{i1}P_{1,i+1}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1n}}{1-P_{1i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i-1}}{1-P_{i-1,i}}&-\lambda_{i,i-1}P_{i,i-1}&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i+1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,n}}{1-P_{i-1,i}}\\ -P_{i1}&\cdots&-P_{i,i-1}&1-P_{ii}&-P_{i,i+1}&\cdots&-P_{in}\\ \frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i-1}}{1-P_{i+1,i}}&-\lambda_{i,i+1}P_{i,i+1}&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i+1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,n}}{1-P_{i+1,i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{in}P_{in}P_{n1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{n,i-1}}{1-P_{ni}}&-\lambda_{in}P_{in}&\frac{\lambda_{in}P_{in}P_{n,i+1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{nn}}{1-P_{ni}}\end{array}\right].$
Then $\Delta P=M\Delta P_{ii}/(1-P_{ii})$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M/(1-P_{ii})$ as $\Delta P_{ii}\to 0$. In $M$, there are $22n-$ 2 negative numbers, all in the $i$th row or column; the others are positive. We also let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
Theorem 6
With the above setting and notations, we have
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (22)
and
$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (23)
Thus, the decision rule on globalization could be like this: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and to prevent major internal trade disputes. Also, location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (23) measures the side effects of $i$’s protectionism on the rest of the world. In general, some would benefit from it, whereas others would not.
Country $i$’s optimal level of globalization maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$; the first order condition is necessary at the optimal level. This can simply test if $\pi M_{i}$ is a zero vector or not, according to the first part of Corollary 4. Given $\lambda_{ji}$ for all $i$ and $j$ in $\mathcal{N}$, Eq. (22) and (23) for all $i\in\mathcal{N}$ define a dynamical system that describes the instant change of $\pi$ with respect to the diagonal of $P$. Given the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on $P$. In such an algorithm, for example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% at the same time; and this succession continues until the matrix $P$ converges. At the potential convergence, the final $P$ and $\pi$ reach a new counterbalance equilibrium in Eq. (24). However, $P$ needs no more adjustments if $\Lambda$ satisfies certain desirable properties, such as balanced bilateral trade according to Theorem 3 and the second part of Corollary 4. Also, $P$ needs no adjustments if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all countries fight for competitiveness in every bilateral conflict. At the convergence, any tiny perturbation on $P_{ii}$ would have a nonsignificant effect on $\pi$, discouraging $i$ from making small changes. Thus, the diagonal of $P$ is relatively stable levels of globalization in these two special cases. Yet, even if a point satisfies all the first-order conditions, it is optimal for some countries only and not every country maximizes its $\pi_{i}$ at this point. Besides, the derivatives in Eq. (22) and (23) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. For these reasons, it might be of limited practical use to seek a far-reached solution to the first-order conditions, even if it exists. Nevertheless, Theorem 6 provides each country with a right direction to steer.
Corollary 4
1. $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$; 2. If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
$\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$ (24)
where $\bigotimes$ is the element-wise matrix product; as two special cases, Eq. (24) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ for all $i$ and $j$ in $\mathcal{N}$, or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (22) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}^{+}$ or $j\in\mathcal{N}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$; and the positive and negative contributions are partially offset. Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as
$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (25)
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is likely better than a pure globalization or protectionism strategy because the right-hand side of Eq. (25) is less than or equal to
$-\sum\limits_{j\in\mathcal{N}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$, respectively.
6 Empirical Study
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$ for the years 2000 to 2019. The bilateral trade data come from the ComTrade database of the United Nations (2021). We use the import and export data not for consumer goods from the World Bank’s WITS Database (2021). Due to the rationality of importers and exporters, the data have largely reflected political, territorial, ideological, cultural, war, national security and other geopolitical considerations, as well as transaction and shipping costs, resource endowment, industrial distribution, and location advantages. For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. During this period, there were a subprime mortgage crisis between 2007 and 2010 and a European debt crisis between 2009 and 2011. All these events heavily impacted relevant countries. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the United Kingdom, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
First, Figure IV plots $\pi_{i}$ from 2000 to 2019. China and Russia gained, and Canada remained almost unchanged while six others lost. Specifically, the USA’s $\pi_{i}$ steadily grew before 2003; it then sharply dropped till 2011 and slowly recovered. Overall, it dropped by 12.7%. China’s $\pi_{i}$ sharply increased during the two crises but almost leveled off elsewhere. Canada and Russia showed their resilience in $\pi$ because of rising prices of commodities in these turmoiling years. The four Western European countries suffered significantly from the debt crisis, and their $\pi_{i}$ continued falling till 2015; manufacturing-focused Germany started falling in 2000 while three others from 2008. Japan’s $\pi_{i}$ plummeted nearly half in the first eight years but finally stopped this momentum between 2007 and 2014. It suffered a domino effect from 2000 to 2007 when China and South Korea challenged its automobile and electric industries. Beyond these nine countries, the others overall became more competitive with their total competitiveness improved by 22.7%, as seen in Figure VI. Many southern asian-, southeast asian-, and middle east countries had tripled their competitiveness including, but not limited to, India, Indonesia, Iraq, United Arab Emirates, and Vietnam.
Also, the Matthew effect was clearly demonstrated. For example, the USA and China’s production ratio was $1.344$ in 2018, but their $\pi_{i}$ ratio was $3.070$. The effect can also be evidenced by Figure VI, in which $\pi_{i}$ has a strong positive correlation with $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
Figure V: The sum of $\pi_{i}$ for all other countries (2000-2019).
Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling $g_{i}$ (2000-2019).
Table I: Country $j$’s direct contribution to $\pi_{i}$ for 2017 and 2019 (in percentage)
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA $\pi_{i}$
CAN $\frac{76.647}{77.390}$ $\frac{1.2635}{1.2732}$ $\frac{0.5390}{0.5179}$ $\frac{0.3222}{0.2781}$ $\frac{0.5922}{0.5728}$ $\frac{0.2877}{0.2607}$ $\frac{0.4329}{0.4768}$ $\frac{0.0410}{0.0355}$ $\frac{14.492}{13.916}$$\frac{.025965}{.026125}$
CHN $\frac{0.3952}{0.3439}$ $\frac{85.584}{84.126}$ $\frac{0.8233}{0.8565}$ $\frac{0.3967}{0.3931}$ $\frac{0.3423}{0.3479}$ $\frac{0.1878}{0.2052}$ $\frac{1.3131}{1.4868}$ $\frac{0.3110}{0.3095}$ $\frac{1.5849}{2.1950}$$\frac{.093905}{.086596}$
DEU $\frac{0.2088}{0.1682}$ $\frac{1.2344}{1.2456}$ $\frac{67.501}{67.410}$ $\frac{2.2469}{2.3993}$ $\frac{1.8879}{2.0095}$ $\frac{1.4928}{1.5634}$ $\frac{0.4265}{0.3507}$ $\frac{0.3737}{0.5379}$ $\frac{3.2504}{3.0282}$$\frac{.040521}{.040105}$
FRA $\frac{0.2345}{0.2171}$ $\frac{0.7032}{0.6617}$ $\frac{3.5649}{3.5687}$ $\frac{74.940}{75.190}$ $\frac{1.9053}{1.8352}$ $\frac{1.6149}{1.5794}$ $\frac{0.2216}{0.2111}$ $\frac{0.1656}{0.2247}$ $\frac{2.4124}{2.3005}$$\frac{.030196}{.030772}$
GBR $\frac{0.6234}{0.6356}$ $\frac{0.9550}{0.9428}$ $\frac{2.2098}{2.4071}$ $\frac{1.5773}{1.6503}$ $\frac{75.873}{75.862}$ $\frac{0.6921}{0.7259}$ $\frac{0.3716}{0.4077}$ $\frac{0.1986}{0.1736}$ $\frac{4.0273}{3.5605}$$\frac{.037385}{.036390}$
ITA $\frac{0.1418}{0.1419}$ $\frac{1.0159}{0.9305}$ $\frac{3.2895}{3.1651}$ $\frac{2.3107}{2.3567}$ $\frac{1.3228}{1.2695}$ $\frac{74.945}{75.662}$ $\frac{0.1958}{0.2123}$ $\frac{0.4274}{0.4525}$ $\frac{1.7055}{1.5402}$$\frac{.019532}{.020507}$
JPN $\frac{0.3297}{0.3118}$ $\frac{2.0275}{1.9912}$ $\frac{0.5242}{0.4869}$ $\frac{0.3184}{0.3041}$ $\frac{0.4209}{0.3614}$ $\frac{0.2218}{0.2139}$ $\frac{87.560}{87.198}$ $\frac{0.1581}{0.1924}$ $\frac{2.2315}{1.9753}$$\frac{.043971}{.045859}$
RUS $\frac{0.1182}{0.1099}$ $\frac{2.6727}{2.2351}$ $\frac{1.7330}{1.7824}$ $\frac{0.9775}{1.0709}$ $\frac{0.7020}{0.7266}$ $\frac{0.8945}{0.8129}$ $\frac{0.6044}{0.5044}$ $\frac{80.225}{79.511}$ $\frac{1.5211}{1.6313}$$\frac{.010657}{.012509}$
USA $\frac{1.3878}{1.4971}$ $\frac{0.8107}{1.0322}$ $\frac{0.4527}{0.4749}$ $\frac{0.2580}{0.2740}$ $\frac{0.5935}{0.5643}$ $\frac{0.1733}{0.1802}$ $\frac{0.3514}{0.3897}$ $\frac{0.0368}{0.0464}$ $\frac{92.142}{91.369}$$\frac{.293880}{.272142}$
* The numerators are for 2019 and the denominators for 2017.
[…]
Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
Figure VII plots the posterior bargaining power from 2000 to 2019. It has similar but smoother patterns as Figure IV. Due to rapid production growth, China’s bargaining power steadily rose, including in 2005 and 2015, when it had troughs in $\pi_{i}$. Also, its $\beta_{i}>\pi_{i}$, whereas the USA’s $\beta_{i}<\pi_{i}$ in each year, showing $\beta_{i}$’s mitigation on the Matthew effect in $\pi_{i}$. Additionally, Russia’s bargaining power sharply dropped after the Russo-Georgian War in 2008 and its annexation of Crimea in 2014. Figure VIII compares the bargaining power versus the USA. The four western European countries’ bargaining power against the USA slightly increased after introducing the Eurozone in 1999 and modestly decreased after 2008. On average, they lost 15.7% of their bargaining power to the USA. Besides, China and Russia improved their relative power multiple times in the first 15 years and remained constant after that. Meanwhile, Japan lost about 50%. Modestly immune to the crises, Canada’s relative bargaining strength sharply rose during the crises and quickly returned to its original level due to its strong alignment with the USA.
Figure VIII: Bargaining power versus the USA (2000-2019).
Table II: $1,000\times\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{-25.78}{-5.78}$ $\frac{-5.49}{-0.37}$ $\frac{-1.47}{0.29}$ $\frac{-0.72}{2.07}$ $\frac{-2.35}{-0.16}$ $\frac{-4.87}{-5.35}$ $\frac{-0.62}{-0.53}$ $\frac{15.75}{63.78}$
CHN $\frac{6.58}{6.90}$ $\frac{7.39}{16.62}$ $\frac{5.36}{7.76}$ $\frac{6.05}{8.46}$ $\frac{2.14}{5.25}$ $\frac{11.56}{31.43}$ $\frac{0.89}{-2.48}$ $\frac{45.19}{65.48}$
DEU $\frac{1.18}{0.14}$ $\frac{-7.73}{-6.05}$ $\frac{6.49}{3.43}$ $\frac{10.07}{5.30}$ $\frac{1.29}{0.18}$ $\frac{0.10}{-5.13}$ $\frac{-1.86}{-4.59}$ $\frac{23.54}{15.06}$
FRA $\frac{0.98}{-0.19}$ $\frac{-8.51}{-5.16}$ $\frac{-13.75}{-6.39}$ $\frac{5.54}{2.67}$ $\frac{-4.20}{-2.57}$ $\frac{-0.90}{-3.67}$ $\frac{-2.08}{-3.10}$ $\frac{12.90}{9.72}$
GBR $\frac{0.73}{-0.96}$ $\frac{-16.47}{-7.66}$ $\frac{-18.11}{-7.01}$ $\frac{-5.31}{-2.28}$ $\frac{-4.64}{-2.58}$ $\frac{-3.39}{-7.65}$ $\frac{-2.38}{-1.99}$ $\frac{8.25}{10.81}$
ITA $\frac{1.01}{0.14}$ $\frac{-8.29}{-5.79}$ $\frac{-3.66}{-0.48}$ $\frac{6.12}{3.66}$ $\frac{7.17}{4.41}$ $\frac{-0.21}{-2.97}$ $\frac{-2.65}{-6.84}$ $\frac{12.51}{11.97}$
JPN $\frac{4.31}{3.52}$ $\frac{-26.28}{-14.31}$ $\frac{-0.28}{4.76}$ $\frac{2.01}{3.72}$ $\frac{4.33}{6.66}$ $\frac{0.35}{2.27}$ $\frac{-1.77}{-2.08}$ $\frac{30.88}{69.74}$
RUS $\frac{1.52}{4.49}$ $\frac{-5.95}{3.48}$ $\frac{8.91}{56.17}$ $\frac{9.85}{24.14}$ $\frac{9.05}{26.41}$ $\frac{5.10}{25.58}$ $\frac{2.78}{6.37}$ $\frac{24.25}{56.37}$
USA $\frac{-1.78}{-5.56}$ $\frac{-48.89}{-17.13}$ $\frac{-12.58}{-6.53}$ $\frac{-4.11}{-2.91}$ $\frac{-3.01}{-2.77}$ $\frac{-4.26}{-2.95}$ $\frac{-10.14}{-23.53}$ $\frac{-2.18}{-1.84}$
* The numerators are for 2017 and the denominators for 2000.
Next, Table II lists the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for 2000 and 2017, calculated from Eq. (9) and (12). If the USA had cut 1% of its imports from China in 2017, for example, its $\pi_{i}$ would have increased by .04889%. While the USA may have had reasons to launch trade wars against all eight others, its top target was China in 2017 and Japan in 2000, according to the numerators and denominators of the last row. As seen from the last column, China and Japan would have been the two largest beneficiaries if they had further collaborated with the USA in 2017 and 2000, respectively. However, comparing the numerators and denominators in the column, the benefits were considerably reduced from 2000 to 2017. Also, starting trade friction from any of these eight countries against the USA would be a loss to the country. Additionally, any other country’s trade war against Russia would be nearly fruitless to increase its $\pi_{i}$, since both numerators and denominators are small in the eighth column. Also, the numbers symmetric with the diagonal have opposite signs, verifying Corollary 2.
Table III: $1,000\times\frac{\mathrm{d}\log\pi_{j}}{\mathrm{d}\log P_{ii}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{52.87}{-188.64}$ $\frac{-16.74}{-9.79}$ $\frac{5.29}{17.08}$ $\frac{4.44}{17.11}$ $\frac{-6.47}{13.46}$ $\frac{5.43}{15.50}$ $\frac{-9.91}{3.11}$ $\frac{3.34}{4.84}$ $\frac{-10.59}{-10.07}$
CHN $\frac{195.55}{35.76}$ $\frac{-2324.27}{-1564.33}$ $\frac{106.47}{31.54}$ $\frac{180.96}{36.26}$ $\frac{232.75}{40.17}$ $\frac{126.87}{31.74}$ $\frac{-81.64}{-33.03}$ $\frac{-135.91}{-2.35}$ $\frac{186.27}{38.81}$
DEU $\frac{38.68}{39.59}$ $\frac{-10.75}{-52.20}$ $\frac{-253.88}{-139.10}$ $\frac{-38.50}{-44.43}$ $\frac{0.50}{-0.04}$ $\frac{-45.19}{-47.50}$ $\frac{14.09}{7.10}$ $\frac{-37.45}{-205.30}$ $\frac{41.38}{44.19}$
FRA $\frac{21.86}{20.97}$ $\frac{-9.83}{-22.47}$ $\frac{-69.45}{-62.18}$ $\frac{-41.92}{-12.66}$ $\frac{-17.07}{-22.86}$ $\frac{-67.71}{-56.16}$ $\frac{1.87}{1.19}$ $\frac{-43.68}{-97.76}$ $\frac{22.93}{25.18}$
GBR $\frac{15.92}{-5.81}$ $\frac{-12.59}{-23.38}$ $\frac{-67.31}{-54.42}$ $\frac{-49.89}{-44.88}$ $\frac{164.52}{145.52}$ $\frac{-42.47}{-29.07}$ $\frac{-7.26}{-13.58}$ $\frac{-33.94}{-117.05}$ $\frac{17.78}{26.54}$
ITA $\frac{18.26}{19.31}$ $\frac{1.06}{-15.52}$ $\frac{-48.34}{-45.02}$ $\frac{-36.50}{-36.91}$ $\frac{-1.16}{-0.96}$ $\frac{-257.39}{-190.00}$ $\frac{4.94}{0.75}$ $\frac{-32.92}{-135.50}$ $\frac{20.83}{23.82}$
JPN $\frac{66.38}{212.93}$ $\frac{-191.11}{-612.51}$ $\frac{35.78}{163.77}$ $\frac{56.09}{212.97}$ $\frac{61.37}{201.14}$ $\frac{46.13}{200.28}$ $\frac{-726.94}{-1518.40}$ $\frac{-19.37}{117.48}$ $\frac{69.31}{236.68}$
RUS $\frac{12.12}{7.85}$ $\frac{-20.29}{-11.83}$ $\frac{-16.42}{4.56}$ $\frac{4.27}{7.57}$ $\frac{8.86}{8.30}$ $\frac{-3.98}{3.36}$ $\frac{-3.65}{3.75}$ $\frac{-832.77}{-2139.12}$ $\frac{12.34}{8.00}$
USA $\frac{-142.03}{-638.28}$ $\frac{-567.12}{-794.20}$ $\frac{-449.08}{-239.99}$ $\frac{-308.90}{-168.36}$ $\frac{-283.70}{-260.03}$ $\frac{-279.33}{-167.92}$ $\frac{-411.84}{-742.34}$ $\frac{-410.17}{-643.02}$ $\frac{834.90}{640.50}$
* The numerators are for 2017 and the denominators are for 2000.
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eq. (22), (23), and (12). From the diagonal, globalization significantly helped China, Germany, Italy, Japan, and Russia. It was a good choice for Canada in 2000 but a bad one in 2017. The effects on France were insignificant because of its small numerator and denominator. The USA and the United Kingdom would be anti-globalization advocates; this explains recent events such as the Trump administration’s withdrawal from international organizations and the United Kingdom’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans and immigrants demanded more inexpensive goods from China, uplifting trade deficits and national debts. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000. The side effects on third parties are off the diagonal. For example, the USA’s further globalization benefited all eight other countries, and so its protectionism would meet serious resistance from its partners. China was the largest beneficiary from the USA’s globalization in both years. Canada suffered the least with the USA’s protectionism in 2017; its further globalization would have slightly profited the USA, according to the first row. The last column implies that further globalization from seven other countries would hurt the USA’s $\pi_{i}$; Japan and China topped the list in 2000 and 2017, respectively. Their hurt declined by 70.2% and ascended by 380%, respectively, in the 17 years. In addition, the second row helps to explain the USA and its western allies’ objection to China’s BRI-alike initiatives, intensified by $4.74$ times from 2000 to 2017 according to the ratios of their numerators and denominators.
Finally, we evaluate the side effects on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is calculated from Eq. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the six other countries were slightly negatively affected in 2018 and 2019. Therefore, these countries would likely take no formal action to prevent the war. Also, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second best would be the United Kingdom for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Probably realizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term battle.
Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
Year CAN DEU FRA GBR ITA JPN RUS CHN USA
2018 -26.114 7.103 3.013 .444 6.661 13.639 15.097 105.808 -53.960
2019 -21.815 3.223 2.133 .039 5.324 12.392 14.894 93.744 -45.954
* Country $i$ is for the USA, $j$ for China, and $k$ for others.
The cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
$\frac{\pi_{j}}{\pi_{i}}=\frac{.092618}{.284327}<\frac{g_{j}}{g_{i}}=\frac{17494.79}{23511.17}.$
For any positive $\mathrm{d}P_{ji}$, therefore, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was
[…]
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. So if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Also, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
7 Discussions
Economic globalization and trade wars have emerged as broad public interests because of global competition among countries, large-scale restructuring of labor markets, and international distribution of added value. We argue that import and export data have already reflected these issues. Therefore, we attempt to extract a country’s competitive strength and bargaining power from the data, identify its best strategies for dealing with trade conflicts and cooperation, and determine if it should move more or less toward globalization. The employed network- and game-theoretical methodologies focus on contemporaneously interactive relationships among economies and are free of parameters and pre-set utility functions. The network game approaches could be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
The forgoing account, however, passes over several vital factors that deserve further investigation. First, counterbalance disequilibrium may exist when trades are notably imbalanced. This could happen in many ways. For example, when country $i$ uses borrowed money to consume $j$’s exports, its consumption still has power over $j$’s production. However, the power comes not only from $i$, but also from the money lender, which has some control over $i$’s consumption. If the lender happens to be $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ underestimated. Secondly, unlike an economic superpower, an emerging economy may focus on its absolute growth of exports instead of competitive advantages. During China’s primitive capital accumulation in the 1980s and 1990s, for example, earning extra US dollars was necessary to finance it industrial modernization. Thus, maximizing $\pi_{i}$ may be a low priority for an emerging economy. The time dimension, absent from the current research, is required to formulate growth. Thirdly, we have yet to include economic globalization’s capital, information, and technology sides. They interplay with goods and services, but their data are not as easily accessible as goods and services. Also, several data issues may cast doubt on the results of the empirical study. There are many missing values in the service data, and country $i$’s reported imports from $j$ are often inconsistent with $j$’s reported exports to $i$. Next, the integration of regional economies merits further attention. Theorem 2 could be used to find the suitable candidates for the integration. If country $i$ belongs to several RTAs, however, $\Delta P_{ii}$ may not be proportionally offset by other countries in Theorem 6. Lastly, one could add other variables to the demand-sided counterbalance Eq. (3), especially those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. Any factors directly impacting $i$’s production could be counted in the $i$th row of $P$. Though consumption may still dominate the fractions, not all productions are driven by consumption.
One could extend the framework in several directions. First, between a bilateral trade war and anti-globalization, a country could target a select group of partners without directly affecting others. For example, as Table II indicates, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan. Secondly, though the dynamical system of Eq. (9)–(11) or Eq. (22)–(23) assumes a fixed $P$, we could include a time dimension to $P$. With the time horizon, we could discuss a trade war’s effects on inflation, $\pi_{i}$’s dependence on its historical values, and the improvement of trade deficits. The time dimension could also improve the forecast capacity of the model. Intertemporal delay often exists when production happens in one year but consumption in the following years. If the production takes two years, then the previous year’s imports of intermediate goods could be included in the current year’s completed production. Next, we could let $\lambda_{ji}$ depend on other determinants besides the competitiveness and production ratios. An industrial or geopolitical analysis could shed light on the choice. For example, as Japan and South Korea rivaled for automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was proper for $\lambda_{ji}$. In the same year, no industry could be a battlefield for China and the USA, so $g_{j}/g_{i}$ seemed a natural choice for $\lambda_{ji}$. Also, in the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954), the distance between trade partners is a vital determinant of their trade flows. Thus, one may discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$. Econometric analyses could also bring additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is larger or less than 00, we may say $\pi_{i}$ is over- or under-valued, respectively. There are other regression forms and $(\pi_{i},g_{i})$ could be panel data. Inclusion of additional regressors, such as debt to GDP ratio, to the regression helps better understand $\pi$ by explanatory data. Finally, anticipating the side effects of $i$’s trade war against $j$, country $j$ could seek cooperation with third parties, which would substantially suffer from the conflict. However, adding more countries to both sides may result in a world war of trade, just like the Cold War between 1947 and 1991.
Reference
Appendix
A1. Proof of Theorem 1
We multiply $\vec{1}_{n}$ on both right sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we then obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$. Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
A2. Proof of Theorem 2
We apply matrix calculus with the restriction of Eq. (3) to prove the theorem. When making a small shock or perturbation $\Delta P$ to $P$ in Eq. (3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation of
$\pi+\Delta\pi=(\pi+\Delta\pi)[P+\Delta P]$ (A.1)
subject to $\Delta P\vec{1}_{n}=\vec{0}_{n}$ and $\Delta\pi\vec{1}_{n}=0$. After subtracting $\pi=\pi P$ from Eq. (A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is
$\Delta\pi[I_{n}-P]\approx\pi\Delta P.$ (A.2)
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{j}$, and $\pi_{-i}$. Accordingly, $P$ have three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the same row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$.
Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\left[I_{n}-P\right]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\left(\vec{0}^{\top}_{i-1},-\lambda_{ji}\pi_{i}+\pi_{j},\vec{0}^{\top}_{j-i-1},\lambda_{ji}\pi_{i}-\pi_{j},\vec{0}^{\top}_{n-j}\right)$ (A.3)
where the right-hand side has all zeros except the $i$th and the $j$th elements.
Also, we partition the transpose of $P$ as
$P^{\top}=\left[\begin{array}[]{ccccc}H_{1}&\eta_{1}&H_{2}&\eta_{2}&H_{3}\\ \mu_{1}&\boxed{P_{ii}}&\mu_{2}&\boxed{P_{ji}}&\mu_{3}\\ H_{4}&\eta_{3}&H_{5}&\eta_{4}&H_{6}\\ \mu_{4}&\boxed{P_{ij}}&\mu_{5}&\boxed{P_{jj}}&\mu_{6}\\ H_{7}&\eta_{5}&H_{8}&\eta_{6}&H_{9}\end{array}\right]$ (A.4)
where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$.
To solve $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ from Eq. (A.3), we write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq. (A.3) as
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&\vec{0}_{i-1}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&-\lambda_{ji}\pi_{i}+\pi_{j}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&\vec{0}_{j-i-1}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\lambda_{ji}\pi_{i}-\pi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&\vec{0}_{n-j}\end{array}\right].$
Since all rows of Eq (A.5), except for the first one, add up to a zero vector, we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{0}&\vec{0}_{j-i-1}^{\top}&\boxed{1}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$ (A.6)
to the left side of Eq. (A.5). Because Eq. (A.6) is the $(n+1)\times(n+1)$ identity matrix except for $n$ ones in the $(i+1)$st row, the multiplication adds the rows of Eq. (A.5) (except for the first row) to its $(i+1)$st row. Consequently, the $(i+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged.
After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$ (A.7)
To make the matrix in Eq. (A.7) lower-triangular, we multiply the following non-singular matrix
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.8)
to the left side of Eq. (A.7) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.9),
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.9),
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
and thus
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}\right]\\ &=&(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
Similarly, after we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{1}&\vec{0}_{j-i-1}^{\top}&\boxed{0}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$
to the left side of Eq. (A.5), the $(j+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$ (A.10)
We next multiply the following non-singular matrix
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
to the left side of Eq. (A.10) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}&\vec{0}_{n-1}^{\top}&(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.11),
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}}.$
A3. Proof of Corollary 2
First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices. Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative. Therefore, by Eq. (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$. Also,
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\lambda_{ji}\mathrm{d}P_{ji}}=\frac{1}{\lambda_{ji}^{2}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0.$
A4. Proof of Theorem 4
When $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, there are three possible intervals and two values for $\lambda_{ji}$: $0<\lambda_{ji}<\pi_{j}/\pi_{i}$, $\lambda_{ji}=\pi_{j}/\pi_{i}$, $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$, $\lambda_{ji}=g_{j}/g_{i}$, or $g_{j}/g_{i}<\lambda_{ji}<\infty$. The small change $\Delta P_{ji}$ could be positive or negative, and $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$ has the same sign. Using Theorems 2 and 3, Table V lists all possible outcomes for these ten scenarios of $(\lambda_{ji},\Delta P_{ji})$.
Table V: Scenarios of $(\lambda_{ji},\Delta P_{ji})$ and their impacts on country $i$’s $\pi_{i}$ and net trade with $j$
$0<\lambda_{ji}<\frac{\pi_{j}}{\pi_{i}}$ $\lambda_{ji}=\frac{\pi_{j}}{\pi_{i}}$ $\frac{\pi_{j}}{\pi_{i}}<\lambda_{ji}<\frac{g_{j}}{g_{i}}$ $\lambda_{ji}=\frac{g_{j}}{g_{i}}$ $\frac{g_{j}}{g_{i}}<\lambda_{ji}<\infty$
$\Delta P_{ji}>0$ $\pi_{i}\uparrow$ $\Delta\pi_{i}=0$ $\pi_{i}\downarrow$ $\pi_{i}\downarrow\downarrow$ $\pi_{i}\downarrow\downarrow\downarrow$
---- - - --- - -- zero net $+$
$\Delta P_{ji}<0$ $\pi_{i}\downarrow$ $\Delta\pi_{i}=0$ $\pi_{i}\uparrow$ $\pi_{i}\uparrow\uparrow$ $\pi_{i}\uparrow\uparrow\uparrow$
+++++++ + + +++++ + $+$ zero net --
$+$ and -- for net trade surplus and deficit, respectively.
Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, $+$, or --.
$\Delta\pi_{j}$ and $\Delta\pi_{i}$ have opposite signs and country $i$ and $j$’s net trade balances have opposite signs.
From the first row, country $i$ could not increase $\pi_{i}$ and achieve a trade surplus with country $j$ at the same time. When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$. In these ten scenarios, on the one hand, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes. On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
A5. Proof of Theorem 5
Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (16), the Nash bargaining solution is
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-p\right)\left(\frac{1}{x}-\frac{1}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{q}+\frac{p}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Also by Eq. (17) or (18),
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-q\right)\left(\frac{1}{x}-\frac{1}{p}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{p}+\frac{q}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
A6. Proof of Corollary 3
When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Letting $x=\lambda_{ji}$, the bargaining solution to Eq. (16) becomes
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-cp\right)\left(\frac{1}{x}-\frac{c}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{cx}{q}+\frac{cp}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
In the above, we need $c>0$ and $cp\leq q/c$, i.e., $0<c\leq\sqrt{q/p}$.
After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (17) or (18) becomes
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(\frac{q}{c}-x\right)\left(\frac{1}{cp}-\frac{1}{x}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{q}{cx}+\frac{x}{cp}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Finally, it is easy to check that $\sqrt{pq}\in[cp,q/c]$ for any $0<c\leq\sqrt{q/p}$.
A7. Proof of Theorem 6
With the change $\Delta P_{ii}$ on $P_{ii}$, we have $\Delta P=\frac{\Delta P_{ii}}{1-P_{ii}}M$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\frac{M}{1-P_{ii}}$ as $\Delta P_{ii}\to 0$. We write $\frac{M}{1-P_{ii}}$ as $\frac{M}{1-P_{ii}}=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $\frac{M}{1-P_{ii}}$. Then $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$. We divide Eq. (A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi\frac{M}{1-P_{ii}}=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.12)
With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.12) is
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&(\pi\xi_{1},\cdots,\pi\xi_{i-1})^{\top}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&\pi\xi_{i}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&(\pi\xi_{i+1},\cdots,\pi\xi_{j-1})^{\top}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\pi\xi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&(\pi\xi_{j+1},\cdots,\pi\xi_{n})^{\top}\end{array}\right].$
As the sum of the last column of (A.13) $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we multiply the nonsingular matrix of Eq. (A.6) to the left side of Eq. (A.13) to make the $(i+1)$st row of Eq. (A.13) to zeros. After dropping the $(i+1)$st row in Eq. (A.13) and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$ (A.14)
We next multiply the matrix of Eq. (A.8) to the left side of Eq. (A.14) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\frac{1}{1-P_{ii}}\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$
Therefore, by the first row of Eq. (A.15),
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.15),
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}$
and thus
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
A8. Proof of Corollary 4
Part 1: If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eq. (22) and Eq. (23). Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
Conversely, if $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$. Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq. (22), we have
$\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}=0.$ (A.16)
We plug Eq. (A.16) into Eq. (23) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get
$\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
Therefore, $\left(\pi M_{i}\right)^{\top}=(1-P_{ii})\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
Part 2: Using the definition of $M_{i}$, $\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that
$\sum\limits_{j\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}P_{it}=0$
for all $t\not=i$. We sum the above equation over all $t\not=i$ to get
[…]
which leads to
$\sum\limits_{j\not=i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.17)
and
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\bigotimes P\right]_{ji}^{\top}=\pi_{i}.$
When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$.
Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.17) is clearly valid and so is Eq. (24). If $\lambda_{ij}=P_{ji}/P_{ij}$, then Eq. (4) implies Eq. (A.17) and, therefore, Eq. (24) also holds.

### v5

@@
-GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE
-Abstract: The embrace of globalization and protectionism among economies has ebbed and flowed over the past few decades.
-These fluctuations call for quantitative analytics to help countries improve their trade policies.
-Changing attitudes toward globalization also imply that the best trade policies vary over time and are country-specific.
-We argue that the imports and exports of all economies constitute a counterbalanced network where conflict and cooperation are two sides of the same coin.
-Quantitative competitiveness is then formulated for each country using a network counterbalance equilibrium.
-A country could improve its relative strength in the network by embracing globalization, protectionism, trade collaboration, or conflict.
-This paper presents some necessary conditions for globalization and trade wars, evaluates their side effects, derives national bargaining powers, identifies appropriate targets for conflict or collaboration, and recommends fair resolutions for trade conflicts.
-Data and events from 2000 to 2019 support these conditions.
-Keywords:  globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma
-JEL Code: C71, C78, F11, F13, F15, O24
+
+Abstract
+This paper explores the exciting dynamics of international trade, focusing on the strategic balance between competition and cooperation in the global trade network.
+It highlights how competitive advantages, rather than comparative advantages, drive trade conflicts and deglobalization.
+By using the balance of power, the paper introduces a quantitative measure of competitiveness that serves as a common goal for all countries, alongside trade balance.
+It then examines how nations can boost their competitive strength and trade balance through globalization, protectionism, trade collaboration, or conflict.
+The study provides practical insights for policymakers on managing trade relationships, resolving conflicts, and determining the right level of globalization by analyzing trade data.
+The analysis is supported by a comparison between theory-driven quantitative evidence and historical events, using real-world trade data from 2000 to 2019.
+keywords:
+globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
+Declarations of interests: I have nothing to declare.
+Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
+Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author never used any Generative AI and AI-assisted technologies.
 1 Introduction
-De-globalization and a trade war are two levels of economic conflict; the latter directly involves only two trade partners, whereas the former represents one country’s disintegration from the global economy.
-Nevertheless, both could adversely affect global economic activity.
-This paper applies a unified game-theoretic approach to formulate simple but decisive rules to mitigate either type of conflict.
-We focus on import and export activities, ignoring political, territorial, ideological, military, and cultural conflicts.
-However, geopolitical data have already been incorporated into trade behaviors and volumes.
-From this perspective, a country’s primary purpose in engaging in trade wars or globalization is to improve its competitiveness and trade balance in the trade system.
-Also, the engagement affects the future geopolitical landscape.
-To analytically examine trade wars and globalization, we face several daunting challenges.
-First, there are always conflicts and cooperation in international trade.
-While the trade of goods and services results from bilateral collaboration, a country’s growing strength could weaken its trade partner’s position.
-Thus, a purely cooperative or non-cooperative study could be too impotent to capture the complexity of trade relationships.
-The second challenge is adequately constructing a high-level description of the interactive trade system because exports and imports have already reflected the comparative advantages of specific goods and services.
-The system is a directional network from which we aim to quantitatively formulate the economies’ utility or social welfare functions in this conflicting and cooperative context.
-Next, policy variables, such as imports and exports as production’s percentages, should be linked to these economic objectives.
-So, countries could apply optimization methods to choose appropriate policies to achieve their economic goals.
-Finally, it is fruitful to bring data into the analytics to explain past events and find optimal solutions for the near future.
-The solutions could identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
-In the literature, studies on globalization often capitalize on econometric analysis, which uses historical data to estimate unknown parameters.
-However, the estimated effects could hardly catch up with the recent changes or seize both cooperation and non-cooperation.
-We measure the relative strengths of the countries in the network by the solution to the counterbalance equilibrium, defined by Hu and Shapley, (2003), and use it as an objective in our analysis.
-The solution, called authority distribution, integrates both cooperation and competition inside the global economic system.
-The data-driven distribution is a social welfare function and an endogenously weighted average of utility functions.
-Moreover, it measures the competitive advantages of the countries and their long-term influences on the whole system.
-Also, this competitiveness brings extra welfare beyond the economy of scale and comparative advantage.
-The additional advantage, called the Matthew effect, is another driving force for competition after comparative advantages.
-This paper attempts to provide strategic solutions for policymakers and ignores how an individual firm or economic agent improves firm’s competitiveness through research and development.
-Maximizing national competitive strength is a purview of government policymakers rather than trade companies, which often capitalize on comparative advantages.
-A country is not merely an aggregate of firms, and the interests of trade companies only sometimes align with those of policymakers, who also consider the side effects of trade, such as the shrinkage of domestic labor markets and the overseas shift of industries.
-So, our studies neglect the firm-based micro foundation and maximize a social welfare function, pivoting domestic production decision-making.
-However, the competitiveness measure is also a weighted average of comparative advantages that considerably reflect the profits of domestic firms engaging in international trade and the welfare of consumers.
-When starting a trade war against a trade partner, a country raises tariffs or creates other barriers to increase competitiveness.
-However, the result depends on the counterpart’s reaction.
-We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in a conflict or cooperation.
-On the other hand, the reaction may also benchmark their production ratio to avoid trade deficits.
-Using the threshold and the benchmark ratios, we derive a Nash bargaining solution for countries with heterogeneous preferences between competitiveness and trade balance, and the solution generates consistent national bargaining power.
-It provides a solution to resolve bilateral trade frictions.
+Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production would be maximized if all production factors could move freely.
+However, not all countries are equally prosperous; some succeed while others struggle.
+Even powerful economies have failed in the past, highlighting the need for fair profit sharing in international trade.
+Therefore, countries must compete for their share while cooperating to maximize global production.
+Both competition and cooperation are essential in international trade.
+This paper offers fair and strategic solutions for each nation within the trade network.
+Analyzing trade wars and globalization presents several challenges.
+First, international trade involves both conflict and cooperation.
+While voluntary trade of goods and services results from bilateral collaboration, a country’s growing strength can weaken its trade partner’s position.
+Thus, a purely cooperative or non-cooperative study may not capture the complexity of trade relationships.
+The second challenge is creating a high-level description of the interactive trade system that focuses on competitive advantage, as exports and imports already reflect the comparative advantages of specific goods and services.
+The system is a directional network from which we aim to quantitatively formulate the economies’ utility and social welfare functions in this conflicting and cooperative context.
+Policy variables, such as imports and exports as percentages of production, should be linked to these economic objectives.
+Countries can then apply optimization methods to choose appropriate policies to achieve their economic goals.
+Finally, incorporating data into the analytics can explain past events and find optimal solutions for the near future.
+The solutions can identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
+We view the international trade system as a game on a network where countries act as nodes and trades as directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016).
+Trade volumes flow along the edges, with imports and exports moving in opposite directions.
+In the multilateral network, a unilateral action could achieve unpredictable or even opposite results.
+Trade wars or anti-globalization efforts are often framed to protect domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping measures.
+These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005).
+Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of trade counterparts.
+For each country, the strengths and weaknesses lie in its domestic products and foreign imports and exports, forming a coalitional game.
+The Shapley value (1953) of the game provides a revealed cardinal preference or utility function for the individual country (Roth, 1977) and a power index in completing domestic production.
+The common battlefield for competition and cooperation is a social welfare function, which weights the individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987).
+These weights are chosen to balance the powers within the network (Hu and Shapley, 2003).
+The social welfare function measures the relative strengths of each country and serves as an objective in our analysis.
+It integrates both cooperation and competition within the global economic system, assessing the competitive advantages of countries and their long-term impacts on the entire system.
+This competitiveness generates additional welfare beyond economies of scale.
+The extra advantage, known as the Matthew effect, is another driving force for collaboration beyond comparative advantages.
+De-globalization and trade wars represent two levels of economic conflict.
+The former involves a country’s withdrawal from the global economy, while the latter involves direct conflict between two trade partners.
+Both can negatively affect global economic activity.
+We apply a unified game-theoretic approach to formulate simple yet decisive rules to mitigate either type of conflict.
+Given that competitiveness has a constant sum in the game, the pursuit of competitiveness becomes the primary reason for initiating a trade war or anti-globalization action, while comparative advantages promote globalization and trade collaboration.
+This study focuses on import and export activities, considering that geopolitical factors may already be incorporated into trade behaviors and volumes.
+Given the rationality of importers and exporters, trade data largely reflect political, territorial, ideological, cultural, war, national security, and other geopolitical considerations, alongside transaction and shipping costs, resource endowment, industrial distribution, and location advantages.
+From this perspective, a country’s primary objective in engaging in trade wars or de-globalization is to enhance its trade balance and competitiveness, as indicated by the data-driven social welfare function.
+The objective of this paper is to offer strategic solutions for policymakers to enhance national competitive strength, rather than focusing solely on the competitiveness of individual firms.
+Government policymakers, rather than trade companies, are responsible for improving national competitiveness, as trade companies often exploit comparative advantages.
+A nation is not merely a collection of firms, and the interests of trade companies do not always align with those of policymakers, who must also consider the negative impacts of trade, such as the contraction of domestic labor markets and the relocation of industries overseas.
+Therefore, our study disregards production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead maximizes the social welfare function, emphasizing domestic production decision-making by global consumers.
+However, the measure of competitiveness is also a weighted average of comparative advantages, weighted by the preferences of consumers, whether affluent or impoverished.
+When a country initiates a trade war against a trade partner, it raises tariffs or imposes other barriers to enhance competitiveness.
+However, the outcome depends on the partner’s response.
+We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in conflict or cooperation.
+Additionally, the reaction may benchmark their production ratio to avoid trade deficits.
+Utilizing the threshold and benchmark ratios, we derive Nash bargaining solutions (1950), generating consistent national bargaining power and providing methods to resolve bilateral trade frictions.
 A country with no specific targets to confront could choose between further globalization or protectionism.
-By minor perturbations in its overall exports, we calibrate the effect on its competitiveness from the counterbalance equilibrium.
-The sign of the effect tells if the country should go for more globalization or not.
-There are many debates over economic globalization in the literature, and Stiglitz (2002, 2017) documented their detailed viewpoints.
-Different viewpoints, assumptions, and considerations could reach opposite conclusions about globalization for a country.
-Our data-driven approach ignores arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security.
+By making minor adjustments to its overall exports, we calibrate the effect on its competitiveness.
+The sign of the effect indicates whether the country should pursue further globalization.
+The analyses of trade wars and globalization also shed light on forming multilateral trade agreements.
+The solutions are data-driven, relying solely on trade and production data.
+For a specific context with a set of fair assumptions, many researchers have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005).
+There are also numerous debates over economic globalization in the literature, with various detailed viewpoints documented in Stiglitz (2002, 2017).
+Different perspectives, assumptions, and considerations can lead to opposing conclusions about globalization for a country.
+Our data-driven approach disregards arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security.
 Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them.
 Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time.
-By analyzing the ComTrade data of the United Nations (2021) for the years 2000 through 2019, we find that protectionism could have been a better choice for the USA and the United Kingdom but a worse option for China, Germany, Japan, and Russia if they targeted competitiveness and trade balances only.
-Data from 2017 also show that trade wars could be an option for the USA to maintain its competitiveness.
-However, adverse side effects on third parties and allies were also tangible.
-According to our theorem of impossibility trilemma, it has to reduce imports if it aims to lower its national debt by earning trade surpluses and also preserve its competitiveness.
+We test our new theories by analyzing ComTrade data from the United Nations (2021) for the years 2000 through 2019, comparing theory-derived results with major economic events during the same period.
+Our findings suggest that protectionism could have been a more advantageous strategy for the USA and the United Kingdom (UK), but a less favorable option for China, Germany, Japan, and Russia, particularly if their objectives were centered on competitiveness.
+Data from 2017 indicate that trade wars might be a viable strategy for the USA to maintain its competitiveness and reduce its trade deficits, and we assess the impacts of the China–USA trade war in 2018 and 2019.
+However, we also observe significant adverse side effects on third parties.
+According to our theorem of impossibility trilemma, the USA must reduce imports if it aims to lower its national debt by generating trade surpluses while preserving its competitiveness.
+Our contribution to the trade literature lies in introducing a novel strategic and comprehensive framework for analyzing cooperation and competition within the global economic system.
+Our analyses eschew unnecessary assumptions and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015).
+Unlike traditional studies, our approach does not rely on unknown parameters and formulates a quantitative measure of competitiveness for each country using the balance of power.
+This measure encompasses overall strength, including both direct and indirect influences.
+The study also derives national bargaining powers and offers strategic solutions for policymakers to manage trade relationships effectively.
 In the following exposition, we first introduce the competitiveness objective in Section 2.
-Then, Section 3 formulates a necessary condition for a country to launch a trade war or form a trade partnership to optimize its objectives.
-After that, a bargaining solution is derived in Section 4 to resolve conflicts or share profits.
-Next, Section 5 discusses a similar condition for globalization or protectionism.
-Then, in an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019.
-Finally, we discuss several extensions and limitations of this framework in Section 7.
-Our exposition is self-contained and the proofs are in the Appendix.
-2 Competitive Advantage in the International Trade System
-Based on Adam Smith’s economy of scale and David Ricardo’s comparative advantage, if all production factors were free of flow, then, hopefully, global production would be maximized.
-However, not all countries share the same well-off status; some succeed over some time, whereas others fail during the same period.
-Even the most powerful economies have failed in human history, and there needs to be more mechanisms to enforce the fairness of profit sharing in international trade.
-Thus, a country has to compete for its share and, at the same time, cooperate with others to maximize global production.
-Both competition and cooperation are essential in international trade.
-Domestic production is paramount for policymakers, whereas profit maximization is for firms or producers.
-Production with zero or slightly negative profit means nothing to firm owners but may still be an option for the government to uplift employment.
-Production maximization, for example, results in less unemployment, while labor may be cut or resourced overseas to maximize shareholders’ profits.
-Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors to the firms.
+Section 3 formulates a necessary condition for a country to initiate a trade war or form a trade partnership to optimize its objectives.
+Section 4 derives Nash bargaining solutions to resolve bilateral conflicts or share profits.
+Section 5 discusses a similar condition for globalization or protectionism.
+In an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019.
+Finally, Section 7 discusses several extensions and limitations of this framework.
+Our exposition is self-contained, and the proofs are in the Appendix.
+2 A Network-based Social Welfare Function
+This section introduces a novel type of competitiveness and social welfare function for countries within the trade network.
+Competitiveness is a relative measure derived from a constant-sum game.
+Absolute production and total trade volume may provide limited information regarding an economy’s competitive position.
+The game simulates the inflows and outflows within the network; without change, these steady flows reach a static equilibrium due to long-term accumulation.
+The equilibrium represents the competitiveness vector, which describes the utilities of the countries’ positions in the network game.
+We detail the steps toward the data-based social welfare function and its unique properties, which are unfamiliar in the trade literature.
+Domestic production is paramount for policymakers, whereas profit maximization is crucial for firms or producers.
+Production with zero or slightly negative profit may be insignificant to firm owners but can still be an option for the government to boost employment.
+For example, production maximization results in less unemployment, while labor may be resourced overseas to maximize shareholders’ profits.
+Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors and the distribution of profit to the firms.
 Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
-We regard the international trade system as a game on a network in which countries act as nodes and trades as directional edges.
-The trade volumes flow along the edges, and imports and exports move in opposite directions.
-In the multilateral network, a unilateral action could achieve unpredictable or even opposite results.
-Trade wars or antiglobalization are often framed to shield domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping.
-These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005).
-Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of the trade counterparts.
-This section introduces a new type of competitiveness as an objective and social welfare function for the countries in the network.
-Competitiveness is a relative measure, and we use a constant-sum game to derive it.
-Absolute production and total trade volume, thus, may give limited information on the economy’s competitive position.
-The game mocks the in-and-out flows over the network; without data change, these steady flows reach a static equilibrium due to long-run accumulation.
-The equilibrium is the competitiveness vector, which describes the utilities of the countries’ trade activities.
-We detail the steps toward the data-based utility and social welfare functions, which are unfamiliar in the trade literature.
-2.1 Long-Run Influence in International Trade System
 For countries in the network, we describe their aggregate trade activities with a square matrix $P$.
-Assume there are $n$ countries labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$.
-For any $i,j\in\mathcal{N}$ with $i\not=j$, let $P_{ij}$ be the fraction of country $i$’s production, which exports to country $j$.
-The exports include all services and goods, such as final goods and intermediate goods, both of which may also contain parts made by third parties or country $j$.
-Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as domestically produced final products.
-Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are attributable directly to the production, such as raw materials, intermediate goods, and crude oil.
+Assume there are $n$ countries, labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$.
+For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ be the fraction of country $i$’s production that exports to country $j$.
+The exports include all services and goods, including intermediate goods; they may contain intermediate goods made by third parties or country $j$.
+Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as all domestically produced final products.
+Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are directly attributable to the production, such as raw materials, intermediate goods, and crude oil.
 It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$.
@@
 Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$.
-We use $P_{ii}=1-\sum\limits_{j\not=i}P_{ij}\geq 0$ for the non-exporting fraction.
-Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $j\in\mathcal{N}$.
-We place all the fractions $P_{ij}$ into a $n\times n$ stochastic matrix $P=[P_{ij}]$.
-We look for a utility function which determines country $i$’s production or expected production from a demand- or consumption-side perspective.
+We use $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ for the non-exporting fraction.
+Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $i\in\mathcal{N}$.
+We place all the fractions $P_{ij}$ into an $n\times n$ stochastic matrix $P=[P_{ij}]$.
+Each row of $P$ defines an individual utility function that determines a specific country’s production from a demand- or consumption-side perspective.
 For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
 $v_{i}(S)\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{j\in S}P_{ij},$ (1)
-together with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$.
-So, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$.
-It is normalized since $v_{i}(\mathcal{N})=1$.
-Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game and its Shapley value is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$ (Shapley, 1953).
-Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern (1953) utility function, i.e., $P_{ij}$ is $j$’s utility in the game of $(\mathcal{N},v_{i})$.
-So, countries have different preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$.
-However, their preferences change as $i$ varies in $\mathcal{N}$.
-In the matrix $P$, therefore, each row is an individual utility function whereas column countries are the arguments or social states of the function.
-A social welfare function generally combines these individual utility functions into one, accounting for the diverse judgments of every country in $\mathcal{N}$.
-It rates the set of countries by $\mathcal{N}$ as a whole.
-The element $P_{ij}$ is a type of pro rata power when country $i$ makes its production. The multilinear extension of $v_{i}$ is defined by
+with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$.
+Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$.
+Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$.
+Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern utility function (1953), i.e., $P_{ij}$ measures an appraisal agent’s satisfaction when they play country $j$’s role in the game of $(\mathcal{N},v_{i})$.
+The agent is neutral with respect to ordinary and strategic risks detailed in Roth (1977).
+Thus, countries have cardinal preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$.
+Yet, their preferences change as $i$ varies in $\mathcal{N}$.
+In the matrix $P$, therefore, each row is an individual utility function, whereas column countries are the arguments or social states of the function.
+A social welfare function generally combines these individual utility functions into one, accounting for the diverse value judgments of countries in $\mathcal{N}$.
+It rates the set of countries in $\mathcal{N}$ as a whole and serves as a common platform for the countries to compete and collaborate.
+2.1 Long-Run Influence in the Trade System
+With the normalization (i.e., $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$) and monotonicity of $v_{i}$, $P_{ij}$ bears much similarity to the Shapley-Shubik power index (1954), measuring country $j$’s chance of pivoting in forming the result $v_{i}(\mathcal{N})$.
+Indeed, $P_{ij}$ is a type of pro rata power when country $i$ makes its full production.
+The multilinear extension of $v_{i}$ is a function defined by:
 $\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
-where $x_{j}\in[0,1]$ is the probability with which $j$ participates in the consumption-driven production.
+where $x_{j}\in[0,1]$ is the probability with which country $j$ participates in the consumption-driven production.
 Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$.
 When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production.
-Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also means $j$’s progress in completing its imports from $i$.
-In addition, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$.
+Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also represents $j$’s progress in completing its imports from $i$.
+Additionally, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$.
 Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$.
-For an annual production, for example, each country finishes $1/12=8.33\%$ of its imports from $i$ in any month.
-Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s influence or power when country $i$ completes its production from 00 to $1$.
-According to Owen, (1972), the integral equals the Shapley value for the coalitional game, i.e.,
+For example, if country $j$ completes 25% of its consumption, all other countries also finish their 25% at the same time.
+Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s overall influence or power that propels country $i$ to complete the production.
+According to Owen (1972), the integral equals the Shapley value for the coalitional game, i.e.,
 $\int_{0}^{1}\bigtriangledown\tilde{v}_{i}(x,\cdots,x)\mathrm{d}x=(P_{i1},P_{i2},\cdots,P_{in})^{\top}$
-where $\top$” denotes vector or matrix transpose.
+where $\top$” denotes vector or matrix transpose. Therefore, $P$ contains $n\times n$ quantitative powers, and the balance of power in $P$ refers to a state of stability.
 From a supply-side perspective, inputs to the production function in Eq.
@@
 However, it is impossible to compile a complete list of these variables.
-Their effects are also time-varying and country-specific; each product has a unique way of combining them together.
+Their effects are also time-varying and country-specific; each product has a unique way of combining them.
 For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products.
-They are primarily payments for past, current, or future exports.
-Trade is about exchanging goods and services; as the medium of exchange, monetary flow makes transactions possible in modern international trade.
-Nevertheless, payments alone could not fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only.
+These assets are primarily payments for past, current, or future exports.
+Trade involves exchanging goods and services; and monetary flow as the medium of exchange makes transactions possible in modern international trade.
+Nevertheless, payments alone cannot fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only.
 Zero payments may occur for intrafirm transactions of exports within a multinational company.
-Also, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively.
-Besides, a country can accumulate trade deficits if it is a highly desirable destination for foreign investment.
+Additionally, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively.
+A country can accumulate trade deficits if it is a highly desirable destination for foreign investment.
 For example, U.S.
-Treasury Bonds are lucrative and low-risky investment instruments, so foreign governments or companies must earn extra U.S.
-dollars from their trades to put into their foreign investment portfolios.
+Treasury Bonds are lucrative and risk-free investment instruments after the Cold War, so foreign governments or companies must earn extra U.S.
+dollars from their trades to add to their foreign investment portfolios.
 The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system.
 In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere.
-When China exports a smartphone to the USA, for a fictitious example, the chip and design software of the phone could come from the USA and the camera from Korea.
+When China exports a smartphone to the USA, for example, the chip and design software of the phone could come from the USA, and the camera from Korea.
 Japan could manufacture the camera screen using US patents.
@@
 $\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
-in which $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$.
-It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs all countries in $\mathcal{N}$.
+where $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$.
+It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs through all countries in $\mathcal{N}$.
 Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely.
@@
 The direct influences in $P$ pass through interlinked global value chains and there are millions of them.
-As a result, $P^{t}$ converges only if $t$ is large enough for all the direct impacts to travel from the top of global economic chains down to the very bottom.
+Consequently, $P^{t}$ converges only if $t$ is large enough.
 The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012).
@@
 2.2 Mixed Cooperation and Noncooperation
-The $1\times n$ row vector $\pi$, called authority distribution in Hu and Shapley, (2003), also solves the following counterbalance equilibrium
+The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also solves the following counterbalance equilibrium:
 $\pi=\pi P$ (3)
-subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and $\pi_{i}\geq 0$.
-There exist many counterbalanced systems in the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
+subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$.
+Numerous counterbalanced systems exist in both the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
 The counterbalance in Eq.
-(3) implies $\pi$’s mixed cooperative and noncooperative properties — even the most competitive economy needs both cooperation and competition.
-Between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years.
-Though unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147).
-An econometric analysis, for example, could hardly capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
-From the noncooperative aspect, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase of $\pi_{j}$ may imply a decrease of $\pi_{i}$.
-So, in theory, there are $n(n-1)/2$ potential trade conflicts, either small or large, in the trade system.
-Most of them are minor trade disagreements or disputes, so the “war” metaphor is hyperbolic in these cases.
+(3) implies $\pi$’s mixed cooperative and noncooperative properties.
+This is a distinctive feature between comparative and competitive advantages; while both promote collaborations, competitive advantage may bring trade wars and deglobalization.
+For instance, between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years.
+Although unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147).
+Econometric analysis often struggles to capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
+From a non-cooperative perspective, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$.
+Thus, theoretically, there are $n(n-1)/2$ potential trade conflicts, either small or large, within the trade system.
+Most of these are minor trade disagreements or disputes, making the “war” metaphor hyperbolic in these cases.
 However, it is worth mitigating even minor disputes before they become problematic or catastrophic.
-Also, the increase of $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initialized by changes in the matrix $P$.
-In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively.
-From the cooperative aspect, country $i$ should help $j$ improve $\pi_{j}$ whenever $P_{ji}>0$ because $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in
+Additionally, the increase in $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initiated by changes in the matrix $P$.
+In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively, for bilateral trade wars and globalization.
+From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, as $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in:
 $\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}P_{ji}$ (4)
-from Eq. (3). Also, Eq. (3) implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ in which
+from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t},$ where
 $\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}\left(P^{t}\right)_{ji}$ (5)
 for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$.
-So, Eq.
-(5) further implies that $i$ should help all other countries in the trade system, including the poorest and the least competitive ones, to enhance their $\pi_{j}$.
-Thus, the double-edged counterbalance implies that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$.
-The trade-off is how much it should assist or contest $j$ without sacrificing its national interests and even increasing $\pi_{i}$.
+Thus, Eq.
+(5) further implies that country $i$ should assist all other countries in the trade system, including the poorest and least competitive ones, to enhance their $\pi_{j}$.
+Hence, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$.
+The trade-off is determining how much it should assist or contest $j$ without sacrificing its national interests and potentially increasing $\pi_{i}$.
 […]
 Figure I: Dynamics of authority flow for country $i$ and $j$.
-We can physically visualize $\pi_{i}$ as a container or pool in the dynamics of counterbalance equilibrium, in which comparative advantages are the primary driving force for the flows.
-On the one hand, in the inflow equation Eq.
-(4), $i$ accumulates or absorbs authority from other countries through its direct influences.
-It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, other things being equal.
-It also derives more authority from countries on which it has a sizable direct influence (i.e., $P_{ji}$ is large), other things being equal.
-In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$ and so, authority flows from $j$ to $i$ alongside the same column.
-On the other hand, $i$ also contributes to other countries as seen in the following outflow equation
-$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\not=i}\pi_{k}P_{kj}.$
-The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, other things being equal.
+We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary driving forces for the flows.
+In the inflow equation Eq.
+(4), country $i$ accumulates or absorbs authority from other countries through its direct influences.
+It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, all else being equal.
+Additionally, it derives more authority from countries on which it has a significant direct influence (i.e., $P_{ji}$ is large), all else being equal.
+In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$, and thus, authority flows from $j$ to $i$ along the same column.
+Conversely, country $i$ also contributes to other countries, as illustrated in the following outflow equation:
+$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\neq i}\pi_{k}P_{kj}.$
+The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal.
 In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
-In the accounting of $\pi$, the two countries have a break-even point when
+In the accounting of $\pi$, the two countries reach a break-even point when:
 $\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
-Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ takes more from or gives more to country $j$, respectively.
-For country $i$, a simple implication of Eq.
+Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either takes more authority from or gives more to country $j$, respectively.
+For country $i$, a straightforward implication of Eq.
 (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively.
@@
 However, the outflow from $i$ to $j$ does not necessarily match the inflow.
-With involvements of third parties, the inflows and outflows meet at the equilibrium Eq.
+With the involvements of third parties, the inflows and outflows meet at the equilibrium Eq.
 (3) for $i$, i.e.
-$\pi_{i}P_{ij}+\sum\limits_{k\not=j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\not=j}\pi_{k}P_{ki}=\pi_{i}$.
+$\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}$.
 For the same reason, it could be overly restrictive for all countries to balance their bilateral trades.
-For example, country $i$ still maintains an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
-2.3 $\pi_{i}$ as Competitive Advantage
-Country $i$’s competitive advantage refers to its ability to do better than its competitors in the trade system.
-For any product exported from country $i$ to $j$, $i$’s competitors include all countries other than itself.
-So, its competitive advantage extends from bilateral country relations to a relationship with the whole system — it is the ability to beat the best of all other countries in the $j$th market for any $j\in\mathcal{N}$.
-On the contrary, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over a specific trade partner on particular services or products; it is a bilateral relationship.
-According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, and a low cost of capital.
-These factors, imported from the outside or manufactured domestically, are assembled in the final products.
-Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker — in Eq.
-(4), $\pi_{j}P_{ji}$ contains $j$’s return of competitiveness to $\pi_{i}$.
-In the above smartphone example, the competitive factors include the chips and patents from the USA, the cheap labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea.
+For example, country $i$ can still maintain an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
+Finally, Eq.
+(3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$.
+It represents a type of eigenvector centrality, a subclass of network centrality.
+In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987).
+Our $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market, and the eigenvector $\pi$ measures competitiveness (see Section 2.3).
+The same domestic products are used both domestically or overseas.
+Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal.
+Furthermore, $P$ normalizes the scales of economies such that each row has a unit sum.
+Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
+2.3 $\pi$ as Competitive Advantages
+The competitive advantage of country $i$ refers to its ability to outperform its competitors within the trade system.
+For any product exported from country $i$ to country $j$, the competitors of country $i$ include all other countries.
+Thus, its competitive advantage extends from bilateral country relations to a relationship with the entire system — it is the ability to surpass all other countries in the $j$th market for any $j\in\mathcal{N}$.
+In contrast, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over the trade partner’s opportunity cost; it is a bilateral relationship.
+Therefore, globalization may have made competitive advantages more prevalent than comparative advantages.
+According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, land, market affordability, and low cost of capital.
+These factors, imported from outside or manufactured domestically, are assembled in the final products.
+Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker.
+In the smartphone example, the competitive factors include the chips and patents from the USA, the labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea.
 Various studies have also described qualitative strategies to enhance firm competitiveness.
@@
 To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus.
-Denote $\tau_{j}$ the $j$th row of $P$.
+Denote $\tau_{j}$ as the $j$th row of $P$.
 Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production.
 So, it measures other countries’ aggregate comparative advantages over country $j$.
-Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means a solid competitiveness to occupy the $j$th market and win a lion share of $j$’s imports.
+Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means solid competitiveness to occupy the $j$th market and win a lion’s share of $j$’s imports.
 If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum.
-Yet, the $j$th value is not excluded from $\tau_{j}$ because domestic products also compete with country $j$’s imports when the consumers make their consumption choices, and they have no home bias.
+Yet, the $j$th value is not excluded from $\tau_{j}$, because domestic products also compete with country $j$’s imports when consumers make their consumption choices, and they have no home bias.
 We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
-The weights are not exogenously assigned but endogenously determined.
-Also, more weight should be placed on a more competitive reference market than on a less competitive one because it is more challenging to occupy a competitive market than a less competitive one.
+These individual cardinal utilities are not comparable because it is more challenging to occupy a competitive market than a less competitive one.
+We make them comparable by weighting, and the weights are not exogenously assigned but endogenously determined.
+More weight should be placed on a more competitive reference market than on a less competitive one.
 Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$.
 Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$.
-So, both the weighted score and the weight vector quantify the competitiveness of the countries.
+Both the weighted score and the weight vector quantify the competitiveness of the countries.
 To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$.
@@
 If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
-The national competitiveness $\pi_{i}$ is not the sum of the individual competitiveness of the country’s domestic firms, though they are positively associated.
-In the endogenous weighting, worldwide consumers unconsciously participate in the evaluation of competitiveness, and all products are involved.
-Essentially, countries unknowingly judge countries themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats.
-The structure of the global trade network is, to a great extent, endogenously determined by the distribution of firm productivity in individual countries.
-Nevertheless, government policies can affect the distribution without significantly mitigating the welfare of international trade by switching imports or exports from one country to another.
-It can also enhance firm productivity through tax exemption, subsidies, and heavy tariffs on foreign products.
-Also, $\pi$ is a von Neumann-Morgenstern Bergson social welfare function (e.g., Harsanyi, 1955; Hammond, 1987) since it is a positively weighted sum of von Neumann-Morgenstern individual utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$; the weights are $\pi$ itself.
-Thus, the weighted utility function $\pi$ is not determined by how much production factors are consumed but by how much the consumers prefer the products, judged regardless of national interests or preferences by individual countries.
-Pareto optimality is satisfied when weighing these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$.
-The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant.
-However, the welfare function $\pi$ departs significantly from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014).
-It is not explicitly expressed in a specific functional form of production factors but implied from the counterbalance equilibrium because indirect exports also act upon the final products or services.
-Any exported goods and services from one country to another have different utilities for the exporting and importing countries.
-Therefore, each country has its unique utility function, and a single utility function or linear regression cannot capture the diversity of all countries.
-Also, we use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages these factors into final products and how differently it trades the products with others.
-It is precisely derived from the data and uses no parametric models.
-Thus, no unknown parameters are to be estimated, and no residuals exist to differentiate the data from a parametric model.
-According to Barney and Felin, (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors.
-Because of broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges.
+In the endogenous weighting, worldwide consumers unconsciously participate in evaluating national competitiveness, encompassing all products and services.
+Essentially, countries as a whole unknowingly assess themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats.
+The structure of the global trade network is largely determined endogenously by the distribution of firm productivity within individual countries.
+Nevertheless, government policies can influence this distribution without significantly affecting the welfare of international trade by shifting imports or exports between countries.
+Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing heavy tariffs on foreign products.
+However, national competitiveness $\pi_{i}$ is not merely the sum of the individual competitiveness of domestic firms; firm competitiveness are weighted in $\pi$ by consumers, wealthy or poor, domestic or overseas.
+Affordability thus plays a vital role in $\pi$.
+The positive weights imply a positive correlation between firm and national competitiveness.
+Additionally, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
+Thus, the weighted utility function $\pi$ is determined not by the consumption of production factors or the added value, but by consumer preferences, judged independently of national interests or preferences.
+Pareto optimality is achieved when weighting these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$.
+The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant, such as scaling $\tau_{i}$ by $g_{i}$.
+Harsanyi (1955) and Hammond (1987) provide additional ethical, risk, and utilitarianism connotations for this type of social welfare function.
+With the weights being $\pi$ itself, $\pi$ exhibits some special properties, discussed in Section 2.4.
+However, the social welfare function $\pi$ significantly departs from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014).
+It is not explicitly expressed in a specific functional form of production factors and total factor productivity but implied from the counterbalance equilibrium, as indirect imports also impact the final products or services.
+Exported goods and services have different utilities for exporting and importing countries.
+Therefore, each country has a unique utility function, and a single utility function cannot capture the diversity of all countries.
+We use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages production factors into final products and how differently it trades the products with others.
+It is precisely derived from the data and uses no parametric models, thus no unknown parameters need to be estimated, and no residuals exist to differentiate the data from a parametric model.
+Additionally, according to Barney and Felin (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors.
+Due to broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges.
 With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network.
-Also, unjustified assumptions and overly mathematization could mislead the conclusions and policy implications (e.g., Romer, 2015).
-The index $\pi_{i}$ contains an inalienable part besides absorbing spillover effects through $P$.
-There are $t-1$ occurrences of $P_{ii}$ in Eq.
-(2), which largely accounts for $\pi_{i}P_{ii}$ in Eq.
-(4) as $t\to\infty$.
-A global value chain generally takes only finite steps, but it also contains indefinite domestic loops when we trace all the added values of the final products.
-In the smartphone example, the manufacturing labor’s education, skills, and experience are ultimately achieved domestically over many years, but all somehow contribute to the final product.
-Also, many USA workers develop the patents and design software domestically for an extended period.
-These inner loops affect other countries but not vice versa; this part of competitiveness never flows in from external.
-Thus, social infrastructure and political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that help do business all contribute to national competitiveness (e.g., Delgado et al., 2012).
-Also, raw oil fields, natural gas reserves, rare minerals, and patented technologies do not come from external sources and remain with $\pi_{i}P_{ii}$ as the nation’s core competitive factors.
-Refraining exports of rare earth elements or licensed foreign usage of patented technologies, for example, could destroy one or more global value chains.
-For a specific global value chain, on the other end of the inalienable parts is the design and assembly of the final product.
-Product leadership bridges the market, the consumers, strategic direction, technology innovation, sales, and production factors together; this leadership is another type of competitiveness.
+The index $\pi_{i}$ contains a unique component, $\pi_{i}P_{ii}$, which does not originate from external source, as indicated by Eq.
+(4).
+This encompasses, but is not limited to, raw oil fields, natural gas reserves, rare minerals, and patented technologies, all of which remain integral to the nation’s core competitive factors.
+A global value chain typically involves finite steps but also includes indefinite domestic loops when tracing all the added values of the final products.
+For instance, in the smartphone example, the education, skills, and experience of manufacturing labor are ultimately developed domestically over many years, contributing to the final product.
+Additionally, numerous U.S.
+workers develop patents and design software domestically for extended periods.
+These inner loops affect other countries but not vice versa; this aspect of competitiveness never flows in from external sources.
+Consequently, social infrastructure, political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that facilitate business all contribute to national competitiveness (e.g., Delgado et al., 2012).
+For a specific global value chain, at the other end of the unique component is the design and assembly of the final product.
+Product leadership bridges the market, consumers, strategic direction, technology innovation, sales, and production factors together; this leadership represents another type of competitiveness.
 Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods.
 Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing.
-Technological advancements reduce labor costs, and public safety and social stability curtail the risk of costs.
-Finally, Eq.
-(2) implies a solid positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy.
-Nations sell goods that they have in abundance and buy goods that are in short supply domestically.
-With great abundance and shortage, country $i$ could have vast nonzero elements in the $i$th column and row of the matrix $P$, resulting in astronomical nonzero items summed in Eq.
+Technological advancements reduce labor costs, and public safety and social stability mitigate the risk of costs.
+Eq.
+(2) indicates a robust positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy.
+After adjusting for the sizes of $g_{i}$ in $P$, the interaction effects between competitiveness and economies of scale remain, thereby conferring some cost advantages.
+Nations export goods that are abundant and import goods that are scarce domestically.
+With significant abundance and shortage, country $i$ could have substantial nonzero elements in the $i$th column and row of the matrix $P$, resulting in considerable nonzero items summed in Eq.
 (2).
-A large economy also produces a high differentiation of goods and services, which bear location advantage and minimal transaction costs for domestic consumers.
-Free of tariffs and overseas shipping, they also contain a price advantage.
-For these reasons, $P_{ii}$ is often the largest in the $i$th row of $P$, with or without the home bias.
-Therefore, a massive economy with a comprehensively integrated industrial system (such as China and the USA) is likely, but not necessary, to have a large value in $\pi$.
-A country with a single industrial structure is vulnerable to brutal competition, and its $\pi_{i}$ is fragile and volatile when battling with the USA for competitiveness.
-Therefore, we use $\pi_{i}$ as an objective for country $i$ when making trade policies.
-It quantifies the economy’s depth and interplays with breadth; it also incorporates cooperation and non-cooperation with others.
-It is endogenously extracted from the comparative advantages of all products and services.
-As each country has a unique location, natural and human resources, and economic systems, a single utility function, which presumes one stereotype for all economies, may be inadequate to capture this diversity.
-In our study, we recognize that each country, as is its trade stance with partners, is different.
-Each adds inclusiveness value to the global economy, of which trade between nations is a mainstay.
+A large economy also generates a high differentiation of goods and services, which offer location advantages and minimal transaction costs for domestic consumers.
+Free from tariffs and overseas shipping, these goods also possess a price advantage.
+Consequently, $P_{ii}$ is often the largest in the $i$th row of $P$, irrespective of home bias.
+Therefore, a vast economy with a comprehensively integrated industrial system (such as China and the USA) is likely, though not necessarily, to have a large value in $\pi$.
+A country with a single industrial structure is susceptible to intense competition, and its $\pi_{i}$ is fragile and volatile when competing with the USA for competitiveness.
+Hence, we use $\pi_{i}$ as an objective for country $i$ when formulating trade policies.
+It quantifies the economy’s depth and its interplay with breadth, encompassing both cooperation and non-cooperation with other nations.
+This measure is endogenously derived from the comparative advantages of all products and services.
+Given that each country has unique locations, natural and human resources, and economic systems, a single utility function that assumes one stereotype for all economies may be insufficient to capture this diversity.
+Our study acknowledges that each country, along with its trade stance with partners, is distinct.
+Each contributes inclusiveness value to the global economy, where international trade is a cornerstone.
 2.4 The Matthew Effect in $\pi$
-In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness, i.e.,
-$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\not=i}\pi_{j}P_{ji},$ (7)
-in which the proportion $P_{ii}$ is generally larger than .7.7.7.7.
-Thus, $\pi$ has the Matthew effect of accumulated advantage, sometimes summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies have substantial competitiveness compared to less powerful ones, revealing a nonlinear relation between $\pi_{i}$ and $i$’s production $g_{i}$.
-For example, the USA’s $\pi_{i}$ could be much higher than Japan’s, compared to their production ratio.
-The amplification also comes from the USA’s relative advantages over other countries, compared to Japan’s.
+In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness:
+$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\neq i}\pi_{j}P_{ji},$ (7)
+where the proportion $P_{ii}$ is generally larger than .7.7.7.7.
+This illustrates $\pi$’s Matthew effect of accumulated advantage, often summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies exhibit substantial competitiveness compared to less powerful ones, revealing a quadratic relationship between $\pi_{i}$ and $g_{i}$.
+For instance, the USA’s $\pi_{i}$ could be significantly higher than Japan’s, relative to their production ratio.
+The amplification also stems from the USA’s relative advantages over other countries compared to Japan’s.
 In Eq.
-(7) with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$,
-$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\not=\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\not=\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
-the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ comes from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, in general, for other countries $j$.
-Alternatively, the Matthew effect is largely explained by the preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among the countries according to how competitive they already are.
+(7), with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$, respectively,
+$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
+the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$.
+Alternatively, the Matthew effect is largely explained by preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among countries according to their existing competitiveness.
 The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$.
 If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$.
-So, it has a quadratic growth when $\tilde{g}_{i}$ soars linearly and in a linear regression,
-$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i}$ (8)
-for some positive $c_{1}$ and $c_{2}$.
-Therefore, when dealing with a rapidly thriving country $j$, $i$ should mitigate its conflicts with $j$ before the counterpart becomes too strong to compromise.
-Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say,
-$g_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i},\quad\mathrm{for}\ \mathrm{some}\ c_{4}>0,$
-then $g_{i}$ would be accelerated quadratically, giving rise to $\pi_{i}$’s quartic growth.
-Continuing this iteration, eventually, both $\pi_{i}$ and $g_{i}$ would grow exponentially to their maximum capacity.
+Thus, it exhibits quadratic growth when $\tilde{g}_{i}$ increases linearly, and in a linear regression:
+$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i},$ (8)
+for some $c_{1}>0,c_{2}>0$, where $\epsilon_{i}$ is the residual term.
+Therefore, when dealing with a rapidly thriving country $i$, country $j$ should mitigate its conflicts with $i$ before the counterpart becomes too strong to compromise.
+Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say:
+$\tilde{g}_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i}$
+for some $c_{4}>0$, then $\tilde{g}_{i}$ would be accelerated quadratically, leading to $\pi_{i}$’s quartic growth.
+Continuing this iteration, eventually, both $\pi_{i}$ and $\tilde{g}_{i}$ would grow exponentially to their maximum capacities.
 Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
-Granting extra welfare to competitive economies is another major force that drives trade flows over the network.
-So, an economic superpower would make all efforts to maintain its superpower status when facing challenges from emerging competitors; maintaining its advantage in $\pi$ avoids the domino effect, the opposite of the Matthew effect.
-As Copeland (2014) says, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Thus, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$.
-Accordingly, an emerging superpower’s aspiration to assume its rightful place as a superpower would face serious challenges.
-World War I and II were fiery combats between existing and emerging superpowers, resetting new world orders.
-The welfare of competitive advantage leads to an agglomeration effect, which clusters all countries around a handful of economic superpowers.
-For example, country $i$ considers certain imports from either country $j$ or $k$ where $g_{j}$ is significantly larger than $g_{k}$, other things being equal.
-If $i$ selects $j$, by Eq.
-(4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
-$\pi_{j}\Delta P_{ji}=\pi_{j}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j}{g_{j}}=\frac{\pi_{j}}{g_{j}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j)$
-where $\Delta P_{ji}$ is the change on $P_{ji}$ due to the imports.
-The approximation ignores the ripple effect from the shock.
-If $i$ selects $k$, then the competitiveness country $i$ derives from $k$ in the imports is approximately
-$\pi_{k}\Delta P_{ki}=\pi_{k}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k}{g_{k}}=\frac{\pi_{k}}{g_{k}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k).$
-Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$ and, so, country $i$ would likely choose $j$, other things being equal.
-Otherwise, $k$ could counter the offer by lowering its price of exports to $(\pi_{k}g_{j})/(\pi_{j}g_{k})$ of $j$’s price.
-In either case, a small country (such as $k$) tends to fear the emergence of a neighboring superpower (such as $j$) that shares similar location advantage and resource endowment.
-Finally, Eq.
-(3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$.
-It is a type of eigenvector centrality, a subclass of network centrality.
-In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987).
-Our $P$ has a nonzero diagonal such that domestic and foreign goods can compete for the domestic market, and the eigenvector $\pi$ measures the competitiveness.
-The same domestic products are used domestically or overseas.
-Hu, (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; so, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal.
-Besides, $P$ normalizes the scales of economies such that each row has a unit sum.
-As a result, $P$ has the largest row eigenvalue $1$, which is common to all stochastic matrices, and its unit-sum eigenvector $\pi$, which is specific to $P$.
+Additionally, the welfare associated with competitive advantage leads to an agglomeration effect, clustering all countries around a select few of economic superpowers.
+According to Corollary 1, imports from a large economy are preferred over imports from a smaller economy, all else being equal.
+Consequently, the smaller economy should counter this preference by lowering its export prices.
+Meanwhile, a third party could impose different tariffs on large and small countries to ensure that the same imported products have uniform prices in the third country.
+Furthermore, the large market, characterized by numerous producers and consumers where no single entity can dominate, may be favored by the third party for its market competitiveness.
+In either scenario, a small country tends to fear the emergence of a neighboring superpower that shares similar location advantages and resource endowments.
+Corollary 1
+Assume the Matthew effect. If $g_{j}>g_{k}$, then the third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
+Thus, providing additional welfare to competitive economies is another significant force driving trade flows across the network, alongside comparative advantage.
+However, trade wars and protectionism are more likely due to competitive advantages rather than comparative advantages, which benefit all trade partners.
+Accordingly, an economic superpower would exert all efforts to maintain its status when facing challenges from emerging competitors; preserving its advantage in $\pi$ prevents the domino effect, which is contrary to the Matthew effect.
+As Copeland (2014) states, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$.
+An emerging superpower’s aspiration to assume its rightful place would encounter significant challenges.
+World War I and II were intense conflicts between existing and emerging superpowers, leading to the establishment of new world orders.
 3 Bilateral Trade War for Competitiveness
-Expecting a positive change to $\pi_{i}$, country $i$ considers a trade war against its trade partner $j$.
-Gaming on the matrix $P$, $i$’s first-move action could change the element $P_{ji}$ — the portion of $j$’s production exported to $i$.
-Raising tariffs or decreasing import quotas upon $j$, for example, reduces imports from country $j$ and decreases $P_{ji}$.
+Anticipating a significant positive change to $\pi_{i}$, country $i$ considers initiating a trade war against its trade partner $j$.
+Gaming on the matrix $P$, country $i$’s first move could alter the element $P_{ji}$ — the portion of $j$’s production exported to $i$.
+For instance, raising tariffs or decreasing import quotas on $j$ reduces imports from $j$ and decreases $P_{ji}$.
 Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$.
-For simplicity, we assume that the bilateral conflict involves only the two counterparties without directly sacrificing or benefiting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$.
-Nevertheless, the change on $P_{ji}$ or $P_{ij}$ eventually ramifies over the global value chains in multiple steps, and all third parties get involved indirectly.
-The simplicity could be far from realistic in the 2018 China-USA trade war as the Biden administration also urged its allies to ban Chinese companies from buying advanced chips and chip-making equipment (e.g., New York Times, 2022).
-As a reprisal, country $j$ could react by decreasing $P_{ij}$ or $P_{ji}$ sooner or later.
-Without the reprisal, the war dies out at inception, and country $i$ wins.
-No reaction or unconditional surrender could avoid further damages, but it could also bring in further sanctions from $i$ when $j$ shows a sign of weakness.
-Overreaction to the hostile change on $P_{ji}$ or $P_{ij}$ may also severely hurt $\pi_{j}$.
-After the Trump administration began setting tariffs and other trade barriers on China in January 2018, for example, China responded in April 2018 by imposing tariffs on 128 products imported from America (cf., Washington Post, 2018).
-Say, in retaliation for a $\Delta P_{ji}$ change on $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$.
-Similarly, in retaliation for $i$’s $\Delta P_{ij}$ change on $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$.
-So, $\lambda_{ij}=1/\lambda_{ji}$ in the veil of any first-mover advantage.
-We assume that $\lambda_{ji}>0$ to align the actions from both countries.
-Thus, ignoring the temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact on $\pi$.
-Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; so the reduced exports will be consumed domestically, and the modified $P$ remains as a stochastic matrix.
-We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into a $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
-The coefficient $\lambda_{ji}$ often depends on many exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services.
-In general, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks.
-For a specific context with a set of fair assumptions, however, many researchers in the literature have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005).
-In the following exposition, we first assume the coefficient is an exogenous constant in the potential trade war.
+For simplicity, we assume the bilateral conflict involves only the two counterparties in $P$ without directly affecting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$.
+However, changes in $P_{ji}$ or $P_{ij}$ eventually impact global value chains in multiple steps, indirectly involving all third parties.
+This simplification may not reflect the complexity of the 2018 China-USA trade war, where the Biden administration urged allies to ban Chinese companies from purchasing advanced chips and chip-making equipment (New York Times, 2022).
+In retaliation, country $j$ might decrease $P_{ij}$ or $P_{ji}$ sooner or later.
+Without retaliation, the war ends at inception, and country $i$ wins.
+No reaction or unconditional surrender could avoid further damages but might also invite additional sanctions from $i$ if $j$ shows weakness.
+Overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ may also severely harm $\pi_{j}$.
+For example, after the Trump administration imposed tariffs and trade barriers on China in January 2018, China responded in April 2018 by imposing tariffs on 128 products imported from the USA (Washington Post, 2018).
+We introduce a matrix $\Lambda$ for retaliation actions.
+In retaliation for a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$.
+Similarly, in retaliation for $i$’s change $\Delta P_{ij}$ in $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$.
+Thus, $\lambda_{ij}=1/\lambda_{ji}$, in the veil of any first-mover advantage.
+We assume $\lambda_{ji}>0$ to align actions from both countries.
+Ignoring temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact.
+We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
+Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; thus, the reduced exports will be consumed domestically, and the modified $P$ remains a stochastic matrix.
+The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services.
+Generally, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks.
+In the following exposition, we first assume the coefficient is an exogenous or predetermined constant in the potential trade war.
 Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts.
@@
 Theorem 2
-With the above setting and notations, for any $j\not=i$,
+With the above setting and notations, for any $j\neq i$,
 $\hskip 8.5359pt\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}},$ (9)
@@
 $\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (11)
-Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain
+Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain:
 $\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
-Using Theorem 2, we have the following Corollary 1, already implicated in Eq.
+Using Theorem 2, we have the following Corollary 2, already implied in Eq.
 (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$.
@@
 When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
-Besides, since
+Additionally, since
 $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
 has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3.
-In particular, when $\lambda_{ji}=0$ in Eq.
+In particular, when $\lambda_{ji}=0$ in Eqs.
 (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$.
-Also, Eq.
+Eq.
 (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
 Corollary 3
-For any $i\not=j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
-By varying $j$ in $\mathcal{N}$, we could apply Eq.
-(9) to identify country $i$’s trade collaborators and competitors.
+For any $i\neq j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
+By varying $j$ in Eq.
+(9), we can identify country $i$’s trade collaborators and competitors.
 In Eq.
@@
 When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$.
-If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change on $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$ regardless of the sign of $\mathrm{d}P_{ji}$.
-In search for the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$.
+If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change in $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the sign of $\mathrm{d}P_{ji}$.
+In search of the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$.
 The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
@@
 […]
-$\mathop{\mathrm{argmax}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
-as $i$’s best trade partner and
-$\mathop{\mathrm{argmin}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
-as $i$’s best competitor.
-In Eq.
+Furthermore, country $i$ would choose:
+$\mathop{\mathrm{argmax}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
+as $i$’s best trade partner and:
+$\mathop{\mathrm{argmin}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
+as $i$’s worst competitor.
+In Eqs.
 (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$.
-However, $i$ is not his best partner’s best partner nor his best competitor’s best competitor, according to Corollary 2.
-Thus, $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched.
+However, $i$ is not his best partner’s best partner nor his worst competitor’s worst competitor, according to Corollary 3.
+Thus, countries $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched.
 For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
 Eq.
-(11) is the side effects on all other countries due to $i$’s change on $P_{ji}$, including $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$.
-With the perturbation on $P_{ji}$, some countries gain, and others lose their competitiveness; the gain or loss may be even more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$.
-Though $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not the same magnitude, so one movement could be significant, whereas the other remains insignificant.
+(11) describes the side effects on all other countries due to country $i$’s change in $P_{ji}$.
+With the perturbation in $P_{ji}$, some countries gain, while others lose their competitiveness; the gain or loss may be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$.
+Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not of the same magnitude.
+One movement could be significant, while the other remains insignificant; it is also possible that both movements are insignificant.
 Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
@@
 However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero.
-In practice, however, $i$ unlikely exercises these extreme actions because $\pi_{i}$’s gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq.
+In practice, however, $i$ is unlikely to exercise these extreme actions because its gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq.
 (4), which could be negligible to $i$.
@@
 For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages.
-For centuries, many norms and standards have been set up to tame trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
-4 A Nash Bargaining Solution for Reprisal Coefficients $\Lambda$
-When other factors also play a role, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to start a conflict action against $j$.
-Among them, the most primary are trade deficits.
-This section studies a bilateral bargaining solution for $\lambda_{ji}$ when trade balance is also an objective.
-If small economies are willing to dodge direct competition with large ones, the vast majority of the $n(n-1)/2$ conflicts are resolved or mitigated by the bargaining solution.
+For centuries, many norms and standards have been established to tame unilateral trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
+4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
+When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$.
+Among these factors, trade deficits are paramount.
+This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.
 4.1 Trade Balance
-For a given period, country $i$’s trade surplus from $j$ means $j$’s trade deficit of the same magnitude from $i$, and vice versa.
-With a trade surplus, job opportunities are created in the country, decreasing its unemployment rate and expanding its economy of scale.
-It also improves the economy’s creditworthiness because it can use the surplus to pay down its foreign debt, if any.
-A country’s surplus could result from its increasing competitive advantage in tradable goods and services when it beats the competitors in the counterpart’s market.
-Opinions on the impacts of trade deficits, either bilateral or overall, are diverse and sometimes controversial, although huge deficits can undermine the sustainability of the economy.
-In the short run, however, trade deficits help nations avoid shortages of goods and mitigate related economic problems, such as inflation and poverty.
-Implications of trade deficits also depend on their impacts on national security and how they are financed.
-No nation would completely ignore trade deficits or competitiveness; otherwise, an one-sided emphasis will not last long.
-There is a benchmark choice of $\lambda_{ji}$ for net trade balances: $\lambda_{ji}=g_{j}/g_{i}$, the ratio between their productions.
-In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$ while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$.
-If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these two countries, as claimed by Theorem 3.
-This condition is for instant balance change.
-In aggregate value over a specific time period, there is no trade surplus and deficit between these two countries if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$ because $g_{i}P_{ij}$ is the exports from $i$ to $j$ and $g_{j}P_{ji}$ the exports from $j$ to $i$.
-For $i$ to have a trade surplus with $j$, we need $g_{i}\lambda_{ji}\Delta P_{ji}>g_{j}\Delta P_{ji}$.
+For a given period, a trade surplus for country $i$ with country $j$ implies a trade deficit of the same magnitude for $j$ with $i$, and vice versa.
+A trade surplus generates job opportunities, reduces unemployment, and expands the economy of scale.
+It also enhances creditworthiness by enabling the country to pay down foreign debt.
+A surplus can result from competitive advantages in tradable goods and services.
+Opinions on trade deficits vary, though large deficits can undermine economic sustainability.
+In the short term, trade deficits help avoid shortages and mitigate issues like inflation and poverty.
+The implications of trade deficits also depend on their impact on national security and how they are financed.
+No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis will not endure.
+There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances, $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3.
+For country $i$ to have a trade surplus with country $j$ due to the change $\Delta P_{ij}$, $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$.
 Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$.
-To sustain this trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but hurts $\pi_{i}$ — with overly excessive exports from $i$ to $j$ in the $i$th row of $P$, country $i$ transfers additional production power to $j$, reducing $\pi_{i}$.
-So $\lambda_{ji}$ should be neither too large nor too small.
+To sustain trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$.
+Excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$.
+Therefore, $\lambda_{ji}$ should be neither too large nor too small.
 However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$.
-In this case, both $i$ and $j$ lose exports even though $i$ loses fewer.
-This could eventually adversely affect the productions of both countries in the future so that most countries would avoid mutually damaging stalemates.
-In general, the parameter $\lambda_{ji}$ is determined by both players.
-However, the sign and size of $\Delta P_{ji}$ are often decided by one only, especially if it has overwhelming power over its counterpart.
+In this case, both $i$ and $j$ lose exports, though $i$ loses fewer.
+This could adversely affect future production, so most countries avoid mutually damaging stalemates.
+Generally, $\lambda_{ji}$ is determined by both players, but the sign and size of $\Delta P_{ji}$ are often decided by one, especially if it has overwhelming power.
 Theorem 3
 $\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
-To ease the trade tension, we have studied two ways to determine $\lambda_{ji}$.
-On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change on $\Delta P_{ji}$.
-On the other hand, if they look for a balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ removes any trade deficits in the future.
-Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for a long-run rivalry due to $\pi$’s involvement in all countries and all industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and relieve relevant problems such as unemployment and national debt.
-According to Pozsar, (2022), for example, the 2018 China-USA trade war started in this way: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but built up massive national debt, accumulated through wars on terrors, significant tax cuts, and the 2007 subprime mortgage crisis.
-Nevertheless, when China wanted to build a global 5G network and make cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort.
+To ease trade tension, we have studied two ways to determine $\lambda_{ji}$.
+On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $\Delta P_{ji}$.
+On the other hand, if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates any future trade deficits.
+Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for long-term rivalry due to $\pi$’s involvement in all countries and industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and alleviate related problems such as unemployment and national debt.
+According to Pozsar (2022), for example, the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis.
+Nevertheless, when China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort.
 In this example, the focal point is competitiveness rather than trade balance.
-Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is a proper choice for $\lambda_{ji}$.
-We study a general case when a superpower country $i$ faces imminent challenges from an emerging economy $j$.
-By the Matthew effect, likely $\pi_{j}/\pi_{i}<g_{j}/g_{i}$.
-Based on Theorem 4, country $i$ cannot have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$.
-This can happen only when $\Delta P_{ij}<0$ when both countries lose comparative advantages.
-Figure II illustrates this trilemma, in which up to two policy objectives are possible for country $i$.
-If it adopts a position on edge $b$, for example, then it upturns comparative advantage and earns a trade surplus, but gives up competitiveness unless it absorbs competitiveness from other countries.
-Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned with each other and also do not completely contradict those of country $j$.
-In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas $j$’s at another one.
-A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries enjoy escalating economy of scale and diminishing marginal cost of production.
-According to the proof of Theorem 4, country $j$ could achieve these three objectives listed in the theorem.
-This seems unfair to country $i$.
-However, it is conditional on the Matthew effect, which is somewhat “biased” toward $\pi_{i}$.
-Therefore, Theorem 4 is indeed a redress to the effect.
+Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is an appropriate choice for $\lambda_{ji}$.
+We examine a scenario where a superpower country $i$ faces imminent challenges from an emerging economy $j$.
+According to the Matthew effect, it is likely that $\pi_{j}/\pi_{i}<g_{j}/g_{i}$.
+Based on Theorem 4, country $i$ cannot simultaneously have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$.
+This situation occurs only when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages.
+Figure II illustrates this trilemma, where up to two policy objectives are achievable for country $i$.
+For instance, if country $i$ adopts a position on edge $b$, it can overturn comparative advantage and earn a trade surplus but sacrifices competitiveness unless it absorbs competitiveness from other countries.
+In the early 2000s, bilateral exports between China and the USA increased as a percentage of their respective production.
+According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
 Theorem 4 (Impossibility Trilemma)
-Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not achieve all the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
+Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
 Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
-4.2 The Bargaining Solution
-We consider a compromise solution between trade balance and competitiveness because a small change $\Delta P_{ji}$ would hurt or benefit both countries.
-In the trade system, some countries prefer competitiveness while others prefer improved trade balance (either reducing deficit or increasing surplus).
-Their preferences affect the choice of $\lambda_{ji}$, which holds $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively.
-If the countries are indifferent between competitiveness and trade balance, then the middle point seems to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$.
-However, this middle-point solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$.
-Otherwise, whenever countries $i$ and $j$ have opposite priorities between enhancing competitiveness and improving trade balance, we have a bargaining problem and apply Nash, (1950) bargaining to find a cooperative solution, which leverages between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ and also binds the condition $\lambda_{ji}\lambda_{ij}=1$.
+Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned and do not completely contradict those of country $j$.
+In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas country $j$’s interests could lie at another point.
+A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries benefit from escalating economies of scale and diminishing marginal costs of production.
+According to the proof of Theorem 4, country $j$ could simultaneously achieve all three objectives listed in the theorem.
+This seems unfair to country $i$, but it is conditional on the Matthew effect, which is somewhat biased toward $\pi_{i}$.
+Therefore, Theorem 4 serves as a redress to the effect.
+4.2 Nash Bargaining Solutions
+This subsection examines a compromise solution between trade balance and competitiveness.
+In the trade system, some countries prioritize competitiveness while others aim to improve trade balance, either by reducing deficits or increasing surpluses.
+Their preferences influence the choice of $\lambda_{ji}$, which maintains $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively.
+If countries are indifferent between competitiveness and trade balance, the midpoint appears to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$.
+However, this midpoint solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$.
+Alternatively, when countries $i$ and $j$ have opposing priorities between enhancing competitiveness and improving trade balance, the Nash bargaining solution (1950) can be employed to find a cooperative solution that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while adhering to the condition $\lambda_{ji}\lambda_{ij}=1$.
 Let
 $p\overset{\mathrm{def}}{=\joinrel=}\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}\quad\mathrm{and}\quad q\overset{\mathrm{def}}{=\joinrel=}\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}$ (15)
-which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$.
-Without $i$’s bargaining, $\lambda_{ji}$ can secure $p$ and $q$ by Corollary 1 or Theorem 3.
-Similarly, without $j$’s bargaining, $\lambda_{ij}$ can secure $1/p$ and $1/q$ because
-$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}$
-and
-$\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
-We have two scenarios when country $i$ and $j$ have opposite preferences on $p$ and $q$.
+which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. By Corollary 2 and Theorem 3, without bargaining, $\lambda_{ji}$ can secure $p$ and $q$, and $\lambda_{ij}$ can secure $1/p$ and $1/q$ because:
+$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}\quad\mathrm{and}\quad\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
+We consider two scenarios when countries $i$ and $j$ have opposite preferences on $p$ and $q$.
 In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$.
-So $u_{j}(x)=x$ is an utility function for $j$ because $u_{j}(q)=q>p=u_{j}(p)$.
-Similarly, $u_{i}(x)=x$ is an utility function for $i$ since $u_{i}(1/p)=1/p>1/q=u_{i}(1/q)$.
-These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation.
+However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$.
 By Eq.
-(15), $p$ and $1/q$ are their respective status quo utilities, obtained if one decides not to bargain with the other player.
-Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem
-$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (16)
-[…]
-Figure III: The Nash bargaining solution $\lambda_{ji}^{*}$ for Eq. (16) and (17), respectively.
+(15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other.
 In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$.
-Their utility functions are $u_{j}(x)=q-x$ and $u_{i}(x)=1/p-x$, respectively, and the status quo is 00 for both.
-The maximization problem becomes
-$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(q-\lambda_{ji}\right)\left(\frac{1}{p}-\lambda_{ij}\right).$ (17)
-The right plot of Figure III illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve.
-They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$ which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
-In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal.
-However, we can rephrase Eq.
-(17) by
-$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (18)
-in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively.
-To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss.
-The final outcome is Eq.
-(18).
-Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq.
-(18) and, so, in Eq.
-(17).
-Theorem 5 answers the maximization problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design.
-The solution is the same for both preference scenarios, so it also works when country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance.
-In Eq.
-(16), (17) and (18), both countries have a symmetric role and the same prior bargaining power.
-This is an axiom of the Nash bargaining.
-The solution remains unchanged under any positive affine transformation of the utility functions, according to another axiom of the bargaining solution.
+Theorem 5 addresses the bargaining problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design.
+The solution is consistent across both preference scenarios, thus it remains effective when either country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance.
+The solution remains unchanged under any positive affine transformation of their utility functions, according to a property of the bargaining solution.
 By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$.
-If $g_{j}<g_{i}$, for example, then $\pi_{j}/\pi_{i}<g_{j}/g_{i}$ is more likely to occur than $\pi_{j}/\pi_{i}\geq g_{j}/g_{i}$, according to the effect.
-Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation fills the gap between these two ratios without changing the likelihood.
+Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation bridges the gap between these two ratios.
 Theorem 5
-The Nash bargaining solution to Eq. (16), (17), or (18) is
-$\left\{\begin{array}[]{rcl}\lambda_{ji}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},\hskip 14.22636pt\mathrm{and}\\ \lambda_{ij}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}\end{array}\right.$
+When countries $i$ and $j$ have opposite preferences on $p$ and $q$, the Nash bargaining solution is:
+$\lambda_{ji}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}\hskip 14.22636pt\mathrm{and}\hskip 14.22636pt\lambda_{ij}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}$
 for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
-The solution of $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$, in general, relies on the choice of status quo points.
-In Eq.
-(16), for example, both players could intensify or weaken their status quo positions.
-If the intensification or weakening is symmetrical with both players, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$.
+Figure III illustrates the bargaining solution.
+In the left plot (first scenario), $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$ while $i$ moves from bottom to top along the same curve.
+When they meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, the area of the rectangle cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and the status quo point $(p,1/q)$ is maximized.
+The right plot (second scenario) illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve.
+They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
+Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences.
+The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ rely on the choice of status quo points.
+The effective bargaining set is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$ because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second, respectively) scenario.
+If both players intensify or weaken their status quo positions symmetrically, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$.
 Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$.
@@
 Corollary 4
-For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$ in Eq.
-(16), (17), or (18), the solution is still $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
-In summary, when country $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
-$\lambda_{ji}=\left\{\begin{array}[]{cl}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}&\mathrm{otherwise.}\end{array}\right.$ (19)
+For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$, the bargaining solution remains $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
+Additionally, as Corollary 5 states, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$, as implied in Corollary 2 and Theorem 3.
+In summary, when countries $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
+$\lambda_{ji}=\left\{\begin{array}[]{ll}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},&\mathrm{otherwise.}\end{array}\right.$ (16)
 Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb.
-Country $j$ should avoid competing for a higher $\pi_{j}$ with a significantly larger-sized economy $i$, resulting in $\lambda_{ji}=\pi_{j}/\pi_{i}$.
-If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect.
-Consequently, by Eq.
-(10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, no matter trade surplus or deficit.
-This was probably China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy.
-Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq.
-(10).
-Thus, $j$ should also not compete with $i$.
-As a result, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, (1968)’s power transition theory).
-Finally, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ in Eq. (19) are also Nash bargaining solutions for $\lambda_{ji}$ because
-$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (20)
-and
-$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (21)
-In Eq.
-(20) and (21), both countries battle for better competitiveness and net trade balance, respectively.
-At the solution points, the objective functions are zero and no country gets extra welfare, already implied in Corollary 1 and Theorem 3.
-4.3 Posterior Bargaining Power
-In a real bargaining context, non-economic factors (such as military power, population, culture, tradition, legislation, and location) and alternative optimal criteria could considerably compromise the symmetry axiom and hence deviate the bargaining solutions substantially different from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
-Thus, the countries likely have unequal prior bargaining power in Theorem 5, and we introduce a posterior bargaining power that updates the prior one.
-To extend the bilateral Nash solution in Theorem 5 to multilateral, we define
+Corollary 5
+${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
+Finally, country $j$ should avoid direct confrontation with a significantly larger-sized economy $i$, according to Corollary 6.
+This was likely China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy.
+Consequently, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, 1968).
+Corollary 6
+Country $j$ should avoid homogeneous preferences with a significantly larger-sized country $i$.
+4.3 National Bargaining Power
+In deriving the solution $\lambda_{ji}^{*}$, both countries have a symmetric role and identical prior bargaining power, as per the Nash bargaining axiom.
+However, countries $i$ and $j$ typically possess asymmetric prior bargaining power.
+For instance, we could maximize $\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}$ or $\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ given some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013).
+Without prior knowledge about $\zeta_{ij}$, there is no justification to favor one country over another, resulting in $\lambda^{*}_{ji}$.
+In a real bargaining context, however, the parameter $\zeta_{ij}$ depends on numerous characteristics between these two, as well as other, countries.
+Non-economic factors such as military power, population, culture, tradition, legislation, and location, along with alternative optimal criteria, can significantly compromise the symmetry axiom and lead to bargaining solutions that differ substantially from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
+We introduce a posterior bargaining power that integrates both the symmetric prior and the data, capitalizing on the results in Theorem 5.
+To extend the bilateral Nash solution in Theorem 5 to multilateral negotiations, we define
 $\beta_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{\sqrt{\pi_{i}g_{i}}}{\sum\limits_{j\in\mathcal{N}}\sqrt{\pi_{j}g_{j}}}$
-as country $i$’s posterior bargaining power or national capability in the trade system.
-It is normalized such that all bargaining power sums to one and $\beta_{i}$ can be compared over time.
-Also, it measures bargaining power in multinational negotiations.
-When allocating votes in forming or reforming a multinational alliance, for example, using the power could avoid any bilateral dialogs and discussions.
-Clearly, $\lambda^{*}_{ij}=\beta_{i}/\beta_{j}$ for any $i,j\in\mathcal{N}$.
-The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ construes a linear ordering for all countries and holds a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$.
+as country $i$’s national bargaining power in the trade system.
+It is normalized so that all bargaining power sums to one, allowing $\beta_{i}$ to be compared over time.
+This measure aids multinational negotiations.
+For instance, when allocating votes in forming or reforming a multinational alliance, using this power can avoid bilateral dialogues and discussions.
+Despite factors other than production and competitiveness, bilateral trade agreements remain more prevalent than multilateral ones; Yilmazkuday and Yilmazkuday (2014) explored some reasons.
+The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ constructs a linear ordering for all countries and maintains a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$.
 If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$.
-Specifically, $i$’s bargaining power over $k$ equals the ratio of that of $i$ over $j$ and that of $j$ over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
+Specifically, $i$’s bargaining power over $k$ equals the ratio of $i$’s power over $j$ and $j$’s power over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
+Also, we can recover it directly from the relation $\lambda^{*}_{ik}=\beta_{i}/\beta_{k}$.
 Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$.
-So, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$.
-From the production aspect, $\beta_{i}$ declines when the companies in country $i$ rely on inputs from its partner countries.
-The more the companies depend on the partners, the less the country’s power to coerce concessions from the partners.
-From the consumption aspect, the bargaining power increases when the country consumes more imported goods and services by influencing the production of its counterparts.
-Between the determinants, $g_{i}$ is the country’s breadth of economy, and $\pi_{i}$ is predominantly the depth of the economy since the matrix $P$ in Eq.
-(3) already normalizes the production volumes.
-However, the positive correlation between them implies a risk of domino effect in which a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, which incurs further reduction of $g_{i}$.
-Consequently, $\beta_{i}$ would shrink more than the initial drop of $g_{i}$.
+Thus, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$.
+From the production aspect, $\beta_{i}$ declines when companies in country $i$ rely on inputs from partner countries.
+The more companies depend on partners, the less power the country has to coerce concessions from them.
+From the consumption aspect, bargaining power increases when the country consumes more imported goods and services, thereby influencing the production of its counterparts.
+Between the determinants, $g_{i}$ represents the country’s breadth of economy, while $\pi_{i}$ predominantly reflects the depth of the economy, as the matrix $P$ in Eq.
+(3) already normalizes production volumes.
+However, the positive correlation between them implies a risk of a domino effect, where a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, leading to further reduction in $g_{i}$.
+Consequently, $\beta_{i}$ would shrink more than the initial drop in $g_{i}$.
 5 Economic Globalization
-Globalization refers to removing the barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures.
+Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures.
 In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services.
-A recent example is China’s entry into the WTO in 2001, while the United Kingdom’s Brexit in 2020 was a clear rejection of globalization.
-On the contrary, a protectionism measure could raise tariffs on all imported goods, set import quotas for all other countries, and enact stricter government regulations on all imports.
-In 2018, for example, the Trump administration imposed punishing tariffs on all steel and aluminum imports from all other countries (e.g., Washington Post, 2018).
-Globalization creates winners and losers within a nation but it does not automatically compensate the losers.
+A recent example is China’s entry into the WTO in 2001, while the UK’s Brexit in 2020 was a clear rejection of globalization.
+Conversely, protectionist measures could raise tariffs on imported goods, set import quotas for other countries, and enact stricter government regulations on imports.
+For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports from other countries (Washington Post, 2018).
+Globalization creates winners and losers within a nation but does not automatically compensate the losers, thereby fueling income inequality and political divisions.
 Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers.
 This section questions whether the country should retreat from global economic integration without directly attacking any specific country.
-Again, the objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms.
+The objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms.
 We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
-Country $i$’s globalization action triggers adjustments on the whole matrix $P$.
-Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, say, with a small change $\Delta P_{ii}$ on $P_{ii}$.
+Country $i$’s globalization action triggers adjustments in the entire matrix $P$.
+Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, with a small change $\Delta P_{ii}$.
 If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism.
-As the action has no specific target countries, we make the following assumption: whenever $P$ has a change in the $i$th column, the other columns change proportionally to ensure each row still has a unit sum.
-Thus, the change $\Delta P_{ii}$ on $P_{ii}$ provokes the adjustments $\Delta P$ on $P$ in three levels of fallout.
-First, for any $j\not=i$, $P_{ij}$’s proportional response to $\Delta P_{ii}$ is
-$\Delta P_{ij}=-\frac{P_{ij}}{1-P_{ii}}\Delta P_{ii},$
-and the $i$th row of $\Delta P$ is
-$\Delta P_{ii}\left(-\frac{P_{i1}}{1-P_{ii}},\cdots,-\frac{P_{i,i-1}}{1-P_{ii}},1,-\frac{P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{P_{in}}{1-P_{ii}}\right),$
-which has a zero sum and the adjustments are proportional to the $i$th row of $P$. Secondly, the change $\Delta P_{ij}$ causes
-$\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}=-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}$
-due to country $j$’s reprisal reaction. Therefore, the $i$th column of $\Delta P$ is
-$\Delta P_{ii}\left(-\frac{\lambda_{i1}P_{i1}}{1-P_{ii}},\cdots,-\frac{\lambda_{i,i-1}P_{i,i-1}}{1-P_{ii}},1,-\frac{\lambda_{i,i+1}P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{\lambda_{in}P_{in}}{1-P_{ii}}\right)^{\top}.$
-Lastly, for any $t\not=i$, $P_{jt}$’s proportional response to $\Delta P_{ji}$ is
-$\Delta P_{jt}=-\frac{P_{jt}}{1-P_{ji}}\Delta P_{ji}=\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}.$
-The sum of the $j$th row in $\Delta P$ is
-$-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}+\sum\limits_{t\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}=\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\left[-1+\sum\limits_{t\not=i}\frac{P_{jt}}{1-P_{ji}}\right]\Delta P_{ii}=0.$
-We define the matrix $M$ by
-$\left[\begin{array}[]{ccccccc}\frac{\lambda_{i1}P_{i1}P_{11}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1,i-1}}{1-P_{1i}}&-\lambda_{i1}P_{i1}&\frac{\lambda_{i1}P_{i1}P_{1,i+1}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1n}}{1-P_{1i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i-1}}{1-P_{i-1,i}}&-\lambda_{i,i-1}P_{i,i-1}&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i+1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,n}}{1-P_{i-1,i}}\\ -P_{i1}&\cdots&-P_{i,i-1}&1-P_{ii}&-P_{i,i+1}&\cdots&-P_{in}\\ \frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i-1}}{1-P_{i+1,i}}&-\lambda_{i,i+1}P_{i,i+1}&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i+1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,n}}{1-P_{i+1,i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{in}P_{in}P_{n1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{n,i-1}}{1-P_{ni}}&-\lambda_{in}P_{in}&\frac{\lambda_{in}P_{in}P_{n,i+1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{nn}}{1-P_{ni}}\end{array}\right].$
-Then $\Delta P=M\Delta P_{ii}/(1-P_{ii})$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M/(1-P_{ii})$ as $\Delta P_{ii}\to 0$.
-In $M$, there are $22n-$ 2 negative numbers, all in the $i$th row or column; the others are positive.
-We also let $M_{i}$ be $M$ with its $i$th column dropped.
+As the action has no specific target countries, we assume that whenever $P$ changes the $i$th column, the other columns change proportionally to ensure each row of $P$ still has a unit sum.
+For example, the change $\Delta P_{ji}$ causes proportional changes $\Delta P_{ji}\sigma_{ji}$ to the $j$th row where:
+$\sigma_{ji}\overset{\mathrm{def}}{=\joinrel=}\frac{1}{1-P_{ji}}\left(-P_{j1},\cdots,-P_{j,i-1},1-P_{ji},-P_{j,i+1},\cdots,-P_{jn}\right).$
+Consequently, the change $\Delta P_{ii}$ provokes adjustments $\Delta P$ in three levels of fallout.
+First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$.
+Secondly, the $i$th column in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}$, due to reprisal reactions, where $\odot$ denotes the element-wise product or Hadamard product and $\Lambda_{i}$ is the $i$th row of $\Lambda$.
+Lastly, the $i$th column of $\Delta P$ triggers changes in all other columns.
+We define:
+$M\overset{\mathrm{def}}{=\joinrel=}\mathrm{diag}\left(\sigma_{ii}\odot\Lambda_{i}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right]$
+where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with the diagonal vector $\sigma_{ii}\odot\Lambda_{i}$.
+Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$, and thereby $\Delta P=\Delta P_{ii}M$.
+So, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M$ as $\Delta P_{ii}\to 0$.
+We let $M_{i}$ be $M$ with its $i$th column dropped.
 Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
 Theorem 6
-With the above setting and notations, we have
-$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (22)
+With the above setting and notations, we have:
+$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (17)
 and
-$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (23)
-Thus, the decision rule on globalization could be like this: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is significantly negative or positive, respectively.
-Otherwise, when $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4.
+$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\left(I_{n-1}-Z_{i}\right)^{-1}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (18)
+The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$.
+The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimal level because $P_{ii}$ cannot be $1$ or 00.
+However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may be positive or negative at this optimal level, allowing country $j$ to capitalize on this by making a corresponding negative or positive change in $P_{ij}$, thereby affecting $P_{ii}$.
+To disincentive other countries from making simultaneous movements, we set $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}$ to be a zero vector.
+This can be tested by checking if $\pi M_{i}$ is a zero vector, according to Corollary 7.
+Corollary 7
+$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$.
+Given $\Lambda$, Eqs.
+(17) and (18) for all $i\in\mathcal{N}$ define a dynamical system describing the instantaneous change of $\pi$ with respect to the diagonal of $P$.
+There are $n$ independent variables in the system since the proportions among off-diagonal elements remain unchanged.
+Due to the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on the diagonal.
+For example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% simultaneously; this succession continues until the matrix $P$ converges.
+Alternatively, we study the properties of the solution to the first-order conditions.
+If the solution exists, it satisfies a new counterbalance equilibrium in Eq.
+(19), as stated in Corollary 8.
+Eq.
+(19) places $n-1$ identities to the solution, requiring only one more condition to uniquely identify it.
+Also, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits some desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, resulting in bilateral authority or trade balance, according to Eq.
+(6) or Theorem 3, respectively.
+At the stable solution, any tiny perturbation on the diagonal of $P$ would have a nonsignificant effect on $\pi$, discouraging countries from making small movement.
+Yet, this solution is optimal for some countries only, and not every country maximizes its competitiveness with this solution.
+Additionally, the derivatives in Eqs.
+(17) and (18) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$.
+Therefore, seeking a far-reaching solution to the first-order conditions might be of limited practical use.
+Nevertheless, Theorem 6 provides each country with the right direction to steer.
+Corollary 8
+If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
+$\pi\left[\Lambda\odot P\right]^{\top}=\pi.$ (19)
+Eq. (19) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
+The decision rule on globalization could be as follows: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively.
+Otherwise, when $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4.
 In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies.
-Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and to prevent major internal trade disputes.
-Also, location and cultural proximity often foster the formation of a PTA or RTA.
+Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and prevent major internal trade disputes.
+Location and cultural proximity often foster the formation of a PTA or RTA.
 Besides, Eq.
-(23) measures the side effects of $i$’s protectionism on the rest of the world.
+(18) measures the side effects of $i$’s protectionism on the rest of the world, helping select countries for a PTA or RTA.
 In general, some would benefit from it, whereas others would not.
-Country $i$’s optimal level of globalization maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$; the first order condition is necessary at the optimal level.
-This can simply test if $\pi M_{i}$ is a zero vector or not, according to the first part of Corollary 4.
-Given $\lambda_{ji}$ for all $i$ and $j$ in $\mathcal{N}$, Eq.
-(22) and (23) for all $i\in\mathcal{N}$ define a dynamical system that describes the instant change of $\pi$ with respect to the diagonal of $P$.
-Given the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on $P$.
-In such an algorithm, for example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% at the same time; and this succession continues until the matrix $P$ converges.
-At the potential convergence, the final $P$ and $\pi$ reach a new counterbalance equilibrium in Eq.
-(24).
-However, $P$ needs no more adjustments if $\Lambda$ satisfies certain desirable properties, such as balanced bilateral trade according to Theorem 3 and the second part of Corollary 4.
-Also, $P$ needs no adjustments if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all countries fight for competitiveness in every bilateral conflict.
-At the convergence, any tiny perturbation on $P_{ii}$ would have a nonsignificant effect on $\pi$, discouraging $i$ from making small changes.
-Thus, the diagonal of $P$ is relatively stable levels of globalization in these two special cases.
-Yet, even if a point satisfies all the first-order conditions, it is optimal for some countries only and not every country maximizes its $\pi_{i}$ at this point.
-Besides, the derivatives in Eq.
-(22) and (23) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$.
-For these reasons, it might be of limited practical use to seek a far-reached solution to the first-order conditions, even if it exists.
-Nevertheless, Theorem 6 provides each country with a right direction to steer.
-Corollary 4
-1.
-$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$; 2.
-If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
-$\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$ (24)
-where $\bigotimes$ is the element-wise matrix product; as two special cases, Eq.
-(24) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ for all $i$ and $j$ in $\mathcal{N}$, or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
-Finally, Theorem 6 implies a mixed globalization and protectionism strategy.
-Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively.
+Finally, Theorem 6 implies a mixed globalization and protectionism strategy, helping $i$ to decide which PTA or RTA to join.
+Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively.
 Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq.
 (17) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$.
-Thus, when $j\in\mathcal{N}^{+}$ or $j\in\mathcal{N}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$; and the positive and negative contributions are partially offset.
-Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as
-$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (25)
+Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$, and the positive and negative contributions are partially offset.
+Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as:
+$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}_{i}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}_{i}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (20)
 In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting.
-This strategy is likely better than a pure globalization or protectionism strategy because the right-hand side of Eq.
-(25) is less than or equal to
-$-\sum\limits_{j\in\mathcal{N}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
-Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$, respectively.
+This strategy is better than a pure globalization or protectionism strategy because the right-hand side of Eq.
+(20) is less than or equal to:
+$-\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
+Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$, respectively.
+If most country members of a PTA or RTA are in $\mathcal{N}_{i}^{+}$, then country $i$ should consider participating in the trade agreement.
+However, country $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply that country $i\in\mathcal{N}_{j}^{+}$, possibly causing additional negotiations if $j$ is already in the PTA or RTA.
 6 Empirical Study
-This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$ for the years 2000 to 2019.
-The bilateral trade data come from the ComTrade database of the United Nations (2021).
-We use the import and export data not for consumer goods from the World Bank’s WITS Database (2021).
-Due to the rationality of importers and exporters, the data have largely reflected political, territorial, ideological, cultural, war, national security and other geopolitical considerations, as well as transaction and shipping costs, resource endowment, industrial distribution, and location advantages.
+In this empirical study, we apply our theoretical results to real-world trade data spanning from 2000 to 2019.
+This analysis provides quantitative evidence to support our previous arguments, making the theoretical insights more tangible and relevant for policymakers.
+The empirical analysis reveals trends in national competitiveness and bargaining power over time, highlighting how different countries have gained or lost.
+This provides a historical perspective on the impacts of globalization and trade wars.
+Between 2000 and 2019, the world economy underwent significant transformations and challenges, albeit with varying degrees of success across different regions.
+This period witnessed a substantial increase in global trade volumes, driven by rapid technological advancements, the expansion of international trade agreements, and the rise of emerging markets.
+The integration of China into the WTO in 2001 was a pivotal moment, leading to a surge in global trade activities.
+The rise of emerging markets, particularly China and India, played a crucial role in global economic growth but also led to a trade war between China and the USA.
+The 2008 financial crisis and the subsequent European debt crisis were major turning points, causing widespread economic downturns and leading to substantial policy interventions by governments and central banks worldwide.
+As trade volumes increased with globalization, nations demanded more of the world’s reserve currency, primarily through trade surpluses.
+Consequently, the reserve currency appreciates, decreasing the currency issuer’s exports and increasing trade deficits.
+The issuer may have three options: encouraging local currency settlement in international transactions or reducing its currency’s usage as the world’s reserve currency; ignoring the trade deficits and focusing on national competitiveness while still increasing exports; or rolling back globalization worldwide.
+For the USA, reshoring its manufacturing sectors would enhance its competitiveness and grow exports; however, trade deficit may still be an issue according to the impossibility trilemma.
+For China, significant growth is necessary to shrink the competitiveness gap with the USA, according to the Matthew effect in $\pi$.
+This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$.
+The bilateral trade data are sourced from the United Nations ComTrade database (2021).
+We use import and export data for non-consumer goods from the World Bank’s WITS Database (2021).
 For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018.
-During this period, there were a subprime mortgage crisis between 2007 and 2010 and a European debt crisis between 2009 and 2011.
-All these events heavily impacted relevant countries.
-Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the United Kingdom, and the USA).
+Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA).
 Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively.
@@
 Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
-First, Figure IV plots $\pi_{i}$ from 2000 to 2019.
-China and Russia gained, and Canada remained almost unchanged while six others lost.
-Specifically, the USA’s $\pi_{i}$ steadily grew before 2003; it then sharply dropped till 2011 and slowly recovered.
-Overall, it dropped by 12.7%.
-China’s $\pi_{i}$ sharply increased during the two crises but almost leveled off elsewhere.
-Canada and Russia showed their resilience in $\pi$ because of rising prices of commodities in these turmoiling years.
-The four Western European countries suffered significantly from the debt crisis, and their $\pi_{i}$ continued falling till 2015; manufacturing-focused Germany started falling in 2000 while three others from 2008.
-Japan’s $\pi_{i}$ plummeted nearly half in the first eight years but finally stopped this momentum between 2007 and 2014.
-It suffered a domino effect from 2000 to 2007 when China and South Korea challenged its automobile and electric industries.
-Beyond these nine countries, the others overall became more competitive with their total competitiveness improved by 22.7%, as seen in Figure VI.
-Many southern asian-, southeast asian-, and middle east countries had tripled their competitiveness including, but not limited to, India, Indonesia, Iraq, United Arab Emirates, and Vietnam.
-Also, the Matthew effect was clearly demonstrated.
-For example, the USA and China’s production ratio was $1.344$ in 2018, but their $\pi_{i}$ ratio was $3.070$.
-The effect can also be evidenced by Figure VI, in which $\pi_{i}$ has a strong positive correlation with $g_{i}^{2}$ after controlling $g_{i}$.
+Figure IV illustrates the competitiveness trends from 2000 to 2019.
+During this period, China and Russia experienced significant gains, while Canada remained relatively stable, and the other six countries saw declines.
+Notably, the USA’s competitiveness grew steadily until 2003, followed by a sharp decline until 2011, and a gradual recovery thereafter, resulting in an overall decrease of 12.7%.
+China’s competitiveness surged during the two crises but plateaued otherwise.
+Canada and Russia demonstrated resilience due to rising commodity prices during transformative years.
+The four Western European countries were significantly impacted by the debt crisis, with their competitiveness continuing to fall until 2015.
+Germany, focused on manufacturing, began its decline in 2000, while the other three countries started in 2008.
+Japan’s competitiveness nearly halved in the first eight years but stabilized between 2007 and 2014, suffering a domino effect from 2000 to 2007 as China and South Korea challenged its automobile and electric industries.
+Beyond these nine countries, others collectively improved their competitiveness by 22.7%, as depicted in Figure VI.
+Many Southern Asian, Southeast Asian, and Middle East countries, including India, Indonesia, Iraq, United Arab Emirates, and Vietnam, tripled their competitiveness.
+The Matthew effect is clearly demonstrated.
+For instance, the USA and China’s production ratio was $1.344$ in 2018, but their competitiveness ratio was $3.070$.
+This effect is further evidenced by Figure VI, which shows a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$ after controlling $g_{i}$.
 The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
@@
 * The numerators are for 2019 and the denominators for 2017.
-[…]
+A straightforward implication of Eq.
+(4) is country $j$’s direct contribution to country $i$’s competitiveness in percentage, i.e., $\pi_{j}P_{ji}/\pi_{i}\times 100\%$.
+Table I lists these contributions for 2017 and 2019.
+According to the diagonal numbers, the USA was the most self-sufficient in feeding its competitiveness, with a slight increase from 2017 to 2019.
+The ninth column also shows that the USA was the most significant contributor to Canada, China, Germany, Japan, and the UK in 2019.
+For the same year, China was the most significant contributor to Russia’s competitiveness and the second largest to those of Canada, Japan, and the USA.
+France, Germany, and Italy were highly reliant on each other’s contributions, and their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia.
+In value, the USA contributed .14492.025965=.003763.14492.025965.003763.14492*.025965=.003763.14492 ∗ .025965 = .003763 to Canada while Canada contributed .013878.293880=.004078.013878.293880.004078.013878*.293880=.004078.013878 ∗ .293880 = .004078 to the USA in 2019.
+Comparing the years before and after the beginning of the China-USA trade war, the USA’s contributions to China dropped by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)11.5849.0939052.1950.0865961-(1.5849*.093905)/(2.1950*.086596)1 - ( 1.5849 ∗ .093905 ) / ( 2.1950 ∗ .086596 ), while China’s contribution to the USA reduced by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1.8107.293881.0322.2721421-(.8107*.29388)/(1.0322*.272142)1 - ( .8107 ∗ .29388 ) / ( 1.0322 ∗ .272142 ).
 Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
-Figure VII plots the posterior bargaining power from 2000 to 2019.
-It has similar but smoother patterns as Figure IV.
-Due to rapid production growth, China’s bargaining power steadily rose, including in 2005 and 2015, when it had troughs in $\pi_{i}$.
-Also, its $\beta_{i}>\pi_{i}$, whereas the USA’s $\beta_{i}<\pi_{i}$ in each year, showing $\beta_{i}$’s mitigation on the Matthew effect in $\pi_{i}$.
-Additionally, Russia’s bargaining power sharply dropped after the Russo-Georgian War in 2008 and its annexation of Crimea in 2014.
-Figure VIII compares the bargaining power versus the USA.
-The four western European countries’ bargaining power against the USA slightly increased after introducing the Eurozone in 1999 and modestly decreased after 2008.
-On average, they lost 15.7% of their bargaining power to the USA.
-Besides, China and Russia improved their relative power multiple times in the first 15 years and remained constant after that.
-Meanwhile, Japan lost about 50%.
-Modestly immune to the crises, Canada’s relative bargaining strength sharply rose during the crises and quickly returned to its original level due to its strong alignment with the USA.
+Figure VII illustrates the national bargaining power from 2000 to 2019, displaying patterns similar to those in Figure IV, albeit smoother.
+China’s bargaining power steadily increased due to rapid production growth, even during its $\pi_{i}$’s troughs in 2005 and 2015.
+Notably, China’s bargaining power consistently surpassed its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness, indicating $\beta$’s mitigation of the Matthew effect in $\pi$.
+Russia’s bargaining power sharply declined following the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
+Figure VIII illustrates the comparative bargaining power of the USA against other countries.
+The bargaining power of four Western European countries experienced a slight increase following the establishment of the Eurozone in 1999, but saw a modest decline post-2008, culminating in an average loss of 15.7% of their bargaining power to the USA.
+Conversely, China and Russia significantly enhanced their relative power multiple times within the first 15 years, maintaining a steady level thereafter.
+Japan, however, suffered a substantial reduction, losing approximately 50% of its relative bargaining power.
+Canada’s relative bargaining strength surged during crises but swiftly reverted to its original level, attributable to its strong alignment with the USA.
 Figure VIII: Bargaining power versus the USA (2000-2019).
@@
 * The numerators are for 2017 and the denominators for 2000.
-Next, Table II lists the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for 2000 and 2017, calculated from Eq.
+Table II presents the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated from Eqs.
 (9) and (12).
-If the USA had cut 1% of its imports from China in 2017, for example, its $\pi_{i}$ would have increased by .04889%.
-While the USA may have had reasons to launch trade wars against all eight others, its top target was China in 2017 and Japan in 2000, according to the numerators and denominators of the last row.
-As seen from the last column, China and Japan would have been the two largest beneficiaries if they had further collaborated with the USA in 2017 and 2000, respectively.
-However, comparing the numerators and denominators in the column, the benefits were considerably reduced from 2000 to 2017.
-Also, starting trade friction from any of these eight countries against the USA would be a loss to the country.
-Additionally, any other country’s trade war against Russia would be nearly fruitless to increase its $\pi_{i}$, since both numerators and denominators are small in the eighth column.
-Also, the numbers symmetric with the diagonal have opposite signs, verifying Corollary 2.
+For instance, a 1% reduction in imports from China by the USA in 2017 would have resulted in an increase of .04889% in the USA’s competitiveness.
+Although the USA may have had reasons to initiate trade wars against all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row.
+As evidenced by the last column, China and Japan would have been the two largest beneficiaries had they further collaborated with the USA in 2017 and 2000, respectively.
+However, a comparison of the numerators and denominators in the column reveals a considerable reduction in benefits from 2000 to 2017.
+Additionally, initiating trade friction from any of these eight countries against the USA would result in a loss for the initiating country.
+A trade war against Russia would be nearly fruitless due to the small numerators and denominators in the eighth column.
+Furthermore, the numbers symmetric with the diagonal exhibit opposite signs, corroborating Corollary 3.
 Table III: $1,000\times\frac{\mathrm{d}\log\pi_{j}}{\mathrm{d}\log P_{ii}}$ for years 2000 and 2017
@@
 * The numerators are for 2017 and the denominators are for 2000.
-Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eq.
+Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eqs.
 (17), (18), and (12).
-From the diagonal, globalization significantly helped China, Germany, Italy, Japan, and Russia.
-It was a good choice for Canada in 2000 but a bad one in 2017.
-The effects on France were insignificant because of its small numerator and denominator.
-The USA and the United Kingdom would be anti-globalization advocates; this explains recent events such as the Trump administration’s withdrawal from international organizations and the United Kingdom’s Brexit.
-Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans and immigrants demanded more inexpensive goods from China, uplifting trade deficits and national debts.
+From the diagonal, globalization significantly benefited China, Germany, Italy, Japan, and Russia.
+It was advantageous for Canada in 2000 but detrimental in 2017.
+The effects on France were insignificant due to small numerator and denominator.
+The USA and the UK would be anti-globalization advocates, explaining recent events such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit.
+Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans from manufacturing offshoring demanded more inexpensive goods from China, uplifting trade deficits and national debts but benefiting from low inflation and affordable merchandises.
 Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$.
 Russia was the largest beneficiary of globalization in 2000.
-The side effects on third parties are off the diagonal.
-For example, the USA’s further globalization benefited all eight other countries, and so its protectionism would meet serious resistance from its partners.
-China was the largest beneficiary from the USA’s globalization in both years.
-Canada suffered the least with the USA’s protectionism in 2017; its further globalization would have slightly profited the USA, according to the first row.
-The last column implies that further globalization from seven other countries would hurt the USA’s $\pi_{i}$; Japan and China topped the list in 2000 and 2017, respectively.
-Their hurt declined by 70.2% and ascended by 380%, respectively, in the 17 years.
-In addition, the second row helps to explain the USA and its western allies’ objection to China’s BRI-alike initiatives, intensified by $4.74$ times from 2000 to 2017 according to the ratios of their numerators and denominators.
-Finally, we evaluate the side effects on the third parties due to the USA’s trade war against China in 2018 and 2019.
-Table IV is calculated from Eq.
+The side effects on third parties are positioned off the diagonal.
+For instance, the USA’s further globalization benefited all eight other countries, thereby making its protectionist policies likely to face significant resistance from its partners.
+China emerged as the largest beneficiary of the USA’s globalization in both years.
+Conversely, Canada experienced the least impact from the USA’s protectionism in 2017; its further globalization would have marginally benefited the USA, as indicated by the first row.
+The last column suggests that further globalization from seven other countries would adversely affect the USA; Japan and China topped the list in 2000 and 2017, respectively.
+Their adverse impact decreased by 70.2% and increased by 380%, respectively, over the 17 years.
+Additionally, the second row elucidates the USA and its Western allies’ objection to China’s BRI-like projects, which intensified by $4.74$ times from 2000 to 2017, according to the ratios of their numerators and denominators.
+We assess the externalities on the third parties due to the USA’s trade war against China in 2018 and 2019.
+Table IV is derived from Eqs.
 (9), (11), and (12).
 When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively.
-Canada was a free rider, and the six other countries were slightly negatively affected in 2018 and 2019.
-Therefore, these countries would likely take no formal action to prevent the war.
-Also, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict.
-The second best would be the United Kingdom for the USA and Japan for China, respectively.
+Canada was a free rider, and the other six countries were slightly negatively affected in 2018 and 2019.
+Consequently, these countries would likely refrain from taking formal action to prevent the trade war.
+Furthermore, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict.
+The second-best allies would be the UK for the USA and Japan for China, respectively.
 From 2018 to 2019, the effects were slightly mitigated for all these countries.
-Probably realizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term battle.
+Likely recognizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term conflict.
+However, the cumulative effects would be significant as the conflict is expected to continue for years.
 Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
@@
 * Country $i$ is for the USA, $j$ for China, and $k$ for others.
-The cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
+Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
 $\frac{\pi_{j}}{\pi_{i}}=\frac{.092618}{.284327}<\frac{g_{j}}{g_{i}}=\frac{17494.79}{23511.17}.$
-For any positive $\mathrm{d}P_{ji}$, therefore, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V).
-Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was
+For a positive $\mathrm{d}P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V).
+Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was:
 […]
@@
 According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164.
-So if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China.
-Also, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$.
+Therefore, if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China.
+Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$.
 To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
 7 Discussions
-Economic globalization and trade wars have emerged as broad public interests because of global competition among countries, large-scale restructuring of labor markets, and international distribution of added value.
-We argue that import and export data have already reflected these issues.
-Therefore, we attempt to extract a country’s competitive strength and bargaining power from the data, identify its best strategies for dealing with trade conflicts and cooperation, and determine if it should move more or less toward globalization.
-The employed network- and game-theoretical methodologies focus on contemporaneously interactive relationships among economies and are free of parameters and pre-set utility functions.
-The network game approaches could be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
-The forgoing account, however, passes over several vital factors that deserve further investigation.
-First, counterbalance disequilibrium may exist when trades are notably imbalanced.
-This could happen in many ways.
-For example, when country $i$ uses borrowed money to consume $j$’s exports, its consumption still has power over $j$’s production.
-However, the power comes not only from $i$, but also from the money lender, which has some control over $i$’s consumption.
-If the lender happens to be $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ underestimated.
-Secondly, unlike an economic superpower, an emerging economy may focus on its absolute growth of exports instead of competitive advantages.
-During China’s primitive capital accumulation in the 1980s and 1990s, for example, earning extra US dollars was necessary to finance it industrial modernization.
-Thus, maximizing $\pi_{i}$ may be a low priority for an emerging economy.
-The time dimension, absent from the current research, is required to formulate growth.
-Thirdly, we have yet to include economic globalization’s capital, information, and technology sides.
-They interplay with goods and services, but their data are not as easily accessible as goods and services.
-Also, several data issues may cast doubt on the results of the empirical study.
-There are many missing values in the service data, and country $i$’s reported imports from $j$ are often inconsistent with $j$’s reported exports to $i$.
-Next, the integration of regional economies merits further attention.
-Theorem 2 could be used to find the suitable candidates for the integration.
-If country $i$ belongs to several RTAs, however, $\Delta P_{ii}$ may not be proportionally offset by other countries in Theorem 6.
-Lastly, one could add other variables to the demand-sided counterbalance Eq.
-(3), especially those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$.
-Any factors directly impacting $i$’s production could be counted in the $i$th row of $P$.
-Though consumption may still dominate the fractions, not all productions are driven by consumption.
-One could extend the framework in several directions.
-First, between a bilateral trade war and anti-globalization, a country could target a select group of partners without directly affecting others.
-For example, as Table II indicates, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan.
-Secondly, though the dynamical system of Eq.
-(9)–(11) or Eq.
-(22)–(23) assumes a fixed $P$, we could include a time dimension to $P$.
-With the time horizon, we could discuss a trade war’s effects on inflation, $\pi_{i}$’s dependence on its historical values, and the improvement of trade deficits.
-The time dimension could also improve the forecast capacity of the model.
-Intertemporal delay often exists when production happens in one year but consumption in the following years.
-If the production takes two years, then the previous year’s imports of intermediate goods could be included in the current year’s completed production.
-Next, we could let $\lambda_{ji}$ depend on other determinants besides the competitiveness and production ratios.
-An industrial or geopolitical analysis could shed light on the choice.
-For example, as Japan and South Korea rivaled for automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was proper for $\lambda_{ji}$.
-In the same year, no industry could be a battlefield for China and the USA, so $g_{j}/g_{i}$ seemed a natural choice for $\lambda_{ji}$.
-Also, in the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954), the distance between trade partners is a vital determinant of their trade flows.
-Thus, one may discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
-Econometric analyses could also bring additional insights.
+Economic globalization and trade wars have emerged as significant public interests due to global competition among nations, large-scale restructuring of labor markets, and the international distribution of added value.
+This study posits that import and export data already reflect these issues, and that competitive advantages, rather than comparative advantages, are the primary drivers of trade wars and de-globalization.
+Consequently, we aim to extract a nation’s competitive strength and bargaining power from the data, identify optimal strategies for addressing trade conflicts and cooperation, and determine the extent to which a nation should pursue globalization.
+The methodologies employed, grounded in network and game theories, focus on contemporaneously interactive relationships among economies and are devoid of parameters and pre-set utility functions.
+These network game approaches can be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
+Several critical factors warrant further investigation.
+First, counterbalance disequilibrium may arise when trades are notably imbalanced.
+This imbalance can manifest in various ways.
+For instance, when country $i$ uses borrowed money to consume country $j$’s exports, country $i$’s consumption still exerts influence over country $j$’s production.
+However, this influence originates not only from country $i$, but also from the money lender, who possesses some control over $i$’s consumption.
+If the lender happens to be country $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ is underestimated, potentially leading to a debt trap for country $i$.
+Secondly, unlike an economic superpower, an emerging economy may prioritize the growth of exports over competitive advantages.
+Indeed, its economic growth must significantly surpass that of advanced economies to bridge the competitiveness gap, in accordance with $\pi$’s Matthew effect.
+Thus, maximizing competitive advantages may be a low priority for an emerging economy.
+For example, during China’s primitive capital accumulation in the 1980s and 1990s, earning additional US dollars was essential to finance its industrial modernization.
+The time dimension, absent from the current research, is necessary to formulate growth strategies.
+Thirdly, the facets of economic globalization encompassing capital, information, and technology have yet to be fully integrated.
+These elements interact with goods and services, but their data are not as easily accessible.
+Additionally, several data issues may undermine the empirical study’s results.
+Notably, there are numerous missing values in the service sectors, and discrepancies often arise between country $i$’s reported imports from country $j$ and country $j$’s reported exports to country $i$.
+Furthermore, the integration of regional economies warrants further examination.
+Theorem 2 could be employed to identify suitable candidates for integration.
+However, the underlying assumption may need modification.
+For instance, if country $i$ is a member of multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries as used in Theorem 6.
+Moreover, additional variables could be incorporated into the demand-sided counterbalance Eq.
+(3), particularly those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$.
+For example, incorporating labor loss into $g_{i}$.
+Any factors directly influencing country $i$’s production could be included in the $i$th row of $P$.
+Although consumption may still predominantly drive the fractions, not all production is driven by consumption.
+The framework can be extended in several directions.
+Firstly, amid a bilateral trade war and anti-globalization sentiments, a country may target specific groups of partners without directly impacting others.
+For instance, as illustrated in Table II, the USA could simultaneously increase tariffs on imports from China, Germany, and Japan.
+Secondly, while the dynamical system described by Eqs.
+(9)–(11) or Eqs.
+(17)–(18) assumes a fixed $P$, introducing a temporal dimension to $P$ could provide valuable insights.
+Incorporating a time horizon would facilitate the examination of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and the improvement of trade deficits.
+Additionally, the temporal dimension could enhance the model’s forecast capabilities.
+Intertemporal delays often occur when production takes place in one year but consumption in another.
+If production spans two years, the previous year’s imports of intermediate goods could be included in the current year’s completed production.
+Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios.
+An industrial or geopolitical analysis could provide valuable insights into this choice.
+For instance, as Japan and South Korea competed in automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was appropriate for $\lambda_{ji}$.
+In the same year, no industry served as a battleground for China and the USA, making $g_{j}/g_{i}$ a natural choice for $\lambda_{ji}$.
+Moreover, the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the importance of the distance between trade partners and the type of merchandise as crucial determinants of trade flows.
+Consequently, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
+Econometric analyses could also provide additional insights.
 In the linear regression of Eq.
-(8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is larger or less than 00, we may say $\pi_{i}$ is over- or under-valued, respectively.
-There are other regression forms and $(\pi_{i},g_{i})$ could be panel data.
-Inclusion of additional regressors, such as debt to GDP ratio, to the regression helps better understand $\pi$ by explanatory data.
-Finally, anticipating the side effects of $i$’s trade war against $j$, country $j$ could seek cooperation with third parties, which would substantially suffer from the conflict.
-However, adding more countries to both sides may result in a world war of trade, just like the Cold War between 1947 and 1991.
-Reference
+(8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is greater or less than 00, we may infer that $\pi_{i}$ is over- or under-valued, respectively.
+Including additional regressors, such as debt to GDP ratio, in the regression helps to better understand $\pi$ through explanatory data.
+Finally, anticipating the side effects of country $i$’s trade war against country $j$, country $j$ might seek cooperation with third parties that would substantially suffer from the conflict.
+However, adding more countries to both sides could result in a global trade war, akin to the Cold War between 1947 and 1991.
 Appendix
 A1. Proof of Theorem 1
-We multiply $\vec{1}_{n}$ on both right sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we then obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$.
+We multiply $\vec{1}_{n}$ from the right on both sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$.
 Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
-A2. Proof of Theorem 2
-We apply matrix calculus with the restriction of Eq.
-(3) to prove the theorem.
+A2. Proof of Corollary 1
+Assume that $g_{j}>g_{k}$.
+For example, country $i$ considers certain imports from either country $j$ or $k$, all else being equal.
+If $i$ selects $j$, by Eq.
+(4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
+$\pi_{j}\Delta P_{ji}=\pi_{j}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j}{g_{j}}=\frac{\pi_{j}}{g_{j}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j)$
+where $\Delta P_{ji}$ is the change on $P_{ji}$ due to the imports.
+The approximation ignores the ripple effect from the shock.
+If $i$ selects $k$, then the competitiveness country $i$ derives from $k$ in the imports is approximately
+$\pi_{k}\Delta P_{ki}=\pi_{k}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k}{g_{k}}=\frac{\pi_{k}}{g_{k}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k).$
+Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$, and so, country $i$ would likely choose $j$, all else being equal.
+A3. Proof of Theorem 2
+We apply matrix calculus with the restrictions of Eq.
+(3) and $\pi\vec{1}_{n}=1$ to prove the theorem.
 When making a small shock or perturbation $\Delta P$ to $P$ in Eq.
-(3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation of
+(3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation:
 $\pi+\Delta\pi=(\pi+\Delta\pi)[P+\Delta P]$ (A.1)
@@
 After subtracting $\pi=\pi P$ from Eq.
-(A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is
+(A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is:
 $\Delta\pi[I_{n}-P]\approx\pi\Delta P.$ (A.2)
-We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{j}$, and $\pi_{-i}$.
-Accordingly, $P$ have three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the same row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row.
+We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$.
+Accordingly, $P$ has three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row.
 Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$.
@@
 $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\left[I_{n}-P\right]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\left(\vec{0}^{\top}_{i-1},-\lambda_{ji}\pi_{i}+\pi_{j},\vec{0}^{\top}_{j-i-1},\lambda_{ji}\pi_{i}-\pi_{j},\vec{0}^{\top}_{n-j}\right)$ (A.3)
-where the right-hand side has all zeros except the $i$th and the $j$th elements.
+where the right-hand vector has all zeros except for the $i$th and the $j$th elements.
 Also, we partition the transpose of $P$ as
@@
 where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$.
-To solve $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ from Eq.
-(A.3), we write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq.
+We write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq.
 (A.3) as
 $\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&\vec{0}_{i-1}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&-\lambda_{ji}\pi_{i}+\pi_{j}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&\vec{0}_{j-i-1}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\lambda_{ji}\pi_{i}-\pi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&\vec{0}_{n-j}\end{array}\right].$
-Since all rows of Eq (A.5), except for the first one, add up to a zero vector, we multiply the following non-singular matrix
-$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{0}&\vec{0}_{j-i-1}^{\top}&\boxed{1}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$ (A.6)
-to the left side of Eq.
-(A.5).
-Because Eq.
-(A.6) is the $(n+1)\times(n+1)$ identity matrix except for $n$ ones in the $(i+1)$st row, the multiplication adds the rows of Eq.
-(A.5) (except for the first row) to its $(i+1)$st row.
-Consequently, the $(i+1)$st row of Eq.
-(A.5) becomes a zero vector and all others remain unchanged.
+Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the $(i+1)$st row, making the $(i+1)$st row of Eq.
+(A.5) a zero vector.
 After dropping the $(i+1)$st row in Eq.
-(A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
+(A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
 $\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$ (A.6)
@@
 $\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.7)
-to the left side of Eq. (A.7) to get
+to the left side of Eq. (A.6) to get:
 $\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$
-Therefore, by the first row of Eq. (A.9),
+Therefore, by the first row of Eq. (A.8):
 $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
-Also, by the second row of Eq. (A.9),
+Also, by the second row of Eq. (A.8):
 $-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
-and thus
+and thus:
 $\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}\right]\\ &=&(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
-Similarly, after we multiply the following non-singular matrix
-$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{1}&\vec{0}_{j-i-1}^{\top}&\boxed{0}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$
-to the left side of Eq.
-(A.5), the $(j+1)$st row of Eq.
-(A.5) becomes a zero vector and all others remain unchanged.
+Similarly, after we add all rows (except for the first one) in Eq.
+(A.5) to the $(j+1)$st row, the $(j+1)$st row becomes a zero vector.
 After dropping the $(j+1)$st row in Eq.
-(A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
+(A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
 $\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$ (A.9)
-We next multiply the following non-singular matrix
+We multiply the following non-singular matrix:
 $\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
-to the left side of Eq. (A.10) to get
+to the left side of Eq. (A.9) to get:
 $\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}&\vec{0}_{n-1}^{\top}&(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$
@@
 $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}}.$
-A3. Proof of Corollary 2
+A4. Proof of Corollary 3
 First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices.
 Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative.
-Therefore, by Eq.
+Therefore, by Eqs.
 (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$.
@@
 $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\lambda_{ji}\mathrm{d}P_{ji}}=\frac{1}{\lambda_{ji}^{2}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0.$
-A4. Proof of Theorem 4
+A5. Proof of Theorem 3
+In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$, while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$.
+If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these countries.
+This condition is for instant balance change.
+Over a specific time period, there is no trade surplus or deficit if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$, because $g_{i}P_{ij}$ represents exports from $i$ to $j$ and $g_{j}P_{ji}$ represents exports from $j$ to $i$.
+A6. Proof of Theorem 4
 When $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, there are three possible intervals and two values for $\lambda_{ji}$: $0<\lambda_{ji}<\pi_{j}/\pi_{i}$, $\lambda_{ji}=\pi_{j}/\pi_{i}$, $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$, $\lambda_{ji}=g_{j}/g_{i}$, or $g_{j}/g_{i}<\lambda_{ji}<\infty$.
@@
 When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$.
-In these ten scenarios, on the one hand, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes.
+In these ten scenarios, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes.
 On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
-A5. Proof of Theorem 5
-Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (16), the Nash bargaining solution is
+A7. Proof of Theorem 5
+In the first scenario, country $j$ prefers $q$ to $p$, whereas country $i$ prefers $1/p$ to $1/q$.
+Therefore, $u_{j}(x)=x$ is a utility function for $j$ because $u_{j}(q)>u_{j}(p)$.
+Similarly, $u_{i}(x)=x$ is a utility function for $i$.
+These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation.
+Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem:
+$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (A.11)
+In Eq.
+(A.11), $\lambda_{ji}-p$ and $\lambda_{ij}-\frac{1}{q}$ are the excessive payoffs or welfare for countries $j$ and $i$ in terms of their respective utility functions.
+The bargaining solution maximizes the product of the excessive utilities.
+In the second scenario of opposite preferences, country $j$ prefers $p$ to $q$ while country $i$ prefers $1/q$ to $1/p$.
+Their utility functions could be $u_{j}(x)=u_{i}(x)=-x$, and their status quo points are $u_{j}(q)=-q$ and $u_{i}(1/p)=-1/p$, respectively.
+The maximization problem becomes:
+$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(-\lambda_{ji}+q\right)\left(-\lambda_{ij}+\frac{1}{p}\right).$ (A.12)
+In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (A.12) by:
+$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (A.13)
+in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively.
+To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss.
+The final outcome is Eq.
+(A.13).
+Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq.
+(A.13) and, so, in Eq.
+(A.12).
+Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (A.11), the Nash bargaining solution is:
 $\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-p\right)\left(\frac{1}{x}-\frac{1}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{q}+\frac{p}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
-Also by Eq. (17) or (18),
+Also, by Eq. (A.12) or (A.13):
 $\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-q\right)\left(\frac{1}{x}-\frac{1}{p}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{p}+\frac{q}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
-A6. Proof of Corollary 3
+A8. Proof of Corollary 4
 When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$.
 Letting $x=\lambda_{ji}$, the bargaining solution to Eq.
-(16) becomes
+(A.11) becomes:
 $\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-cp\right)\left(\frac{1}{x}-\frac{c}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{cx}{q}+\frac{cp}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
 In the above, we need $c>0$ and $cp\leq q/c$, i.e., $0<c\leq\sqrt{q/p}$.
-After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (17) or (18) becomes
+After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (A.12) or (A.13) becomes:
 $\mathop{\mathrm{argmax}}\limits_{x>0}\left(\frac{q}{c}-x\right)\left(\frac{1}{cp}-\frac{1}{x}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{q}{cx}+\frac{x}{cp}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
 Finally, it is easy to check that $\sqrt{pq}\in[cp,q/c]$ for any $0<c\leq\sqrt{q/p}$.
-A7. Proof of Theorem 6
-With the change $\Delta P_{ii}$ on $P_{ii}$, we have $\Delta P=\frac{\Delta P_{ii}}{1-P_{ii}}M$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\frac{M}{1-P_{ii}}$ as $\Delta P_{ii}\to 0$.
-We write $\frac{M}{1-P_{ii}}$ as $\frac{M}{1-P_{ii}}=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $\frac{M}{1-P_{ii}}$.
-Then $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$.
+A9. Proof of Corollary 5
+This is because:
+$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (A.14)
+and
+$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (A.15)
+In Eqs.
+(A.14) and (A.15), both countries battle for better competitiveness and net trade balance, respectively.
+At the solution points, the objective functions are zero and no country gets extra welfare,
+A10. Proof of Corollary 6
+Country $j$’s direct competition with $i$ for a higher $\pi_{j}$ results in $\lambda_{ji}=\pi_{j}/\pi_{i}$.
+If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect.
+Consequently, by Eq.
+(10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, regardless of trade surplus or deficit.
+Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq.
+(10).
+Thus, $j$ should also not compete with $i$.
+A11. Proof of Theorem 6
+We write $M=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $M$.
+Since $\sigma_{ji}$ has a zero sum for all $j$, $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$.
 We divide Eq.
-(A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get
-$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi\frac{M}{1-P_{ii}}=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.12)
-With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.12) is
+(A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get:
+$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi M=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.16)
+With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.16) is:
 $\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&(\pi\xi_{1},\cdots,\pi\xi_{i-1})^{\top}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&\pi\xi_{i}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&(\pi\xi_{i+1},\cdots,\pi\xi_{j-1})^{\top}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\pi\xi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&(\pi\xi_{j+1},\cdots,\pi\xi_{n})^{\top}\end{array}\right].$
-As the sum of the last column of (A.13) $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we multiply the nonsingular matrix of Eq.
-(A.6) to the left side of Eq.
-(A.13) to make the $(i+1)$st row of Eq.
-(A.13) to zeros.
-After dropping the $(i+1)$st row in Eq.
-(A.13) and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$ as
-$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$ (A.14)
-We next multiply the matrix of Eq. (A.8) to the left side of Eq. (A.14) to get
-$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\frac{1}{1-P_{ii}}\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$
-Therefore, by the first row of Eq. (A.15),
-$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
-Also, by the second row of Eq. (A.15),
-$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}$
-and thus
-$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
-A8. Proof of Corollary 4
-Part 1: If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eq.
-(22) and Eq.
-(23).
+As the last column of Eq.
+(A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the $(i+1)$st row to make the $(i+1)$st row a zero vector.
+After dropping the $(i+1)$st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
+$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$ (A.18)
+We multiply the matrix of Eq. (A.7) to the left side of Eq. (A.18) to get:
+$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$
+Therefore, by the first row of Eq. (A.19):
+$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
+Also, by the second row of Eq. (A.19):
+$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
+and thus:
+$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
+A12. Proof of Corollary 7
+If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eqs.
+(17) and (18).
 Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
@@
 Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq.
-(22), we have
+(17), we have:
 $\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}=0.$ (A.20)
-We plug Eq. (A.16) into Eq. (23) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get
-$\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
-Therefore, $\left(\pi M_{i}\right)^{\top}=(1-P_{ii})\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
-Part 2: Using the definition of $M_{i}$, $\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that
-$\sum\limits_{j\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}P_{it}=0$
-for all $t\not=i$. We sum the above equation over all $t\not=i$ to get
+We plug Eq. (A.20) into Eq. (18) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get:
+$\left(I_{n-1}-Z_{i}\right)^{-1}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
+Therefore, $\left(\pi M_{i}\right)^{\top}=\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
+A13. Proof of Corollary 8
+Since
 […]
-which leads to
-$\sum\limits_{j\not=i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.17)
-and
-$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\bigotimes P\right]_{ji}^{\top}=\pi_{i}.$
-When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$.
-Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.17) is clearly valid and so is Eq. (24). If $\lambda_{ij}=P_{ji}/P_{ij}$, then Eq. (4) implies Eq. (A.17) and, therefore, Eq. (24) also holds.
+$\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that $-(1-P_{ii})\pi M$, i.e.,
+$\left(\pi_{1}\lambda_{i1}P_{i1},\cdots,\pi_{i-1}\lambda_{i,i-1}P_{i,i-1},-\pi_{i}(1-P_{ii}),\pi_{i+1}\lambda_{i,i+1}P_{i,i+1},\cdots,\pi_{n}\lambda_{in}P_{in}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right],$
+is a zero vector except for its $i$th element. Thus, for any $t\neq i$, the zero value in the $t$th element implies that:
+$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}P_{it}.$
+We sum the above equations over all $t\neq i$ to get:
+$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}\sum_{t\neq i}P_{it},$
+which leads to:
+$\sum\limits_{j\neq i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.21)
+and:
+$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\odot P\right]_{ji}^{\top}=\pi_{i}.$
+When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\odot P\right]^{\top}=\pi$.
+Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq.
+(A.21) is clearly valid and so is Eq.
+(19).
+If $\lambda_{ij}=P_{ji}/P_{ij}$, then $\left[\Lambda\odot P\right]^{\top}=P$ and, therefore, Eq.
+(19) also holds.

### v6

@@
 Abstract
-This paper explores the exciting dynamics of international trade, focusing on the strategic balance between competition and cooperation in the global trade network.
-It highlights how competitive advantages, rather than comparative advantages, drive trade conflicts and deglobalization.
-By using the balance of power, the paper introduces a quantitative measure of competitiveness that serves as a common goal for all countries, alongside trade balance.
-It then examines how nations can boost their competitive strength and trade balance through globalization, protectionism, trade collaboration, or conflict.
-The study provides practical insights for policymakers on managing trade relationships, resolving conflicts, and determining the right level of globalization by analyzing trade data.
-The analysis is supported by a comparison between theory-driven quantitative evidence and historical events, using real-world trade data from 2000 to 2019.
+This paper investigates the evolving dynamics of international trade, emphasizing the strategic interplay between competition and cooperation within the global trade network.
+It argues that competitive advantages — rather than traditional comparative advantages — are the primary drivers of trade conflicts and deglobalization.
+Drawing on the concept of the balance of power, the paper introduces a quantitative measure of competitiveness, which complements the trade balance as a long-term policy objective.
+It further explores how countries can enhance competitiveness and trade balance through globalization, protectionism, collaboration, or trade frictions.
+Using real-world trade data from 2000 to 2019, our empirical study finds parallels between historical developments and quantitative evidence based on this new theory, offering actionable insights to policymakers for managing trade relations, mitigating conflicts, and calibrating the optimal level of globalization.
 keywords:
-globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
-Declarations of interests: I have nothing to declare.
-Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
-Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author never used any Generative AI and AI-assisted technologies.
+globalization, trade war, counterbalance equilibrium, competitive advantage, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
+journal: Economic Analysis and Policy
 1 Introduction
-Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production would be maximized if all production factors could move freely.
-However, not all countries are equally prosperous; some succeed while others struggle.
-Even powerful economies have failed in the past, highlighting the need for fair profit sharing in international trade.
-Therefore, countries must compete for their share while cooperating to maximize global production.
-Both competition and cooperation are essential in international trade.
-This paper offers fair and strategic solutions for each nation within the trade network.
-Analyzing trade wars and globalization presents several challenges.
-First, international trade involves both conflict and cooperation.
-While voluntary trade of goods and services results from bilateral collaboration, a country’s growing strength can weaken its trade partner’s position.
-Thus, a purely cooperative or non-cooperative study may not capture the complexity of trade relationships.
-The second challenge is creating a high-level description of the interactive trade system that focuses on competitive advantage, as exports and imports already reflect the comparative advantages of specific goods and services.
-The system is a directional network from which we aim to quantitatively formulate the economies’ utility and social welfare functions in this conflicting and cooperative context.
-Policy variables, such as imports and exports as percentages of production, should be linked to these economic objectives.
-Countries can then apply optimization methods to choose appropriate policies to achieve their economic goals.
-Finally, incorporating data into the analytics can explain past events and find optimal solutions for the near future.
-The solutions can identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
-We view the international trade system as a game on a network where countries act as nodes and trades as directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016).
-Trade volumes flow along the edges, with imports and exports moving in opposite directions.
-In the multilateral network, a unilateral action could achieve unpredictable or even opposite results.
-Trade wars or anti-globalization efforts are often framed to protect domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping measures.
-These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005).
-Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of trade counterparts.
-For each country, the strengths and weaknesses lie in its domestic products and foreign imports and exports, forming a coalitional game.
-The Shapley value (1953) of the game provides a revealed cardinal preference or utility function for the individual country (Roth, 1977) and a power index in completing domestic production.
-The common battlefield for competition and cooperation is a social welfare function, which weights the individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987).
-These weights are chosen to balance the powers within the network (Hu and Shapley, 2003).
-The social welfare function measures the relative strengths of each country and serves as an objective in our analysis.
-It integrates both cooperation and competition within the global economic system, assessing the competitive advantages of countries and their long-term impacts on the entire system.
-This competitiveness generates additional welfare beyond economies of scale.
-The extra advantage, known as the Matthew effect, is another driving force for collaboration beyond comparative advantages.
-De-globalization and trade wars represent two levels of economic conflict.
-The former involves a country’s withdrawal from the global economy, while the latter involves direct conflict between two trade partners.
-Both can negatively affect global economic activity.
-We apply a unified game-theoretic approach to formulate simple yet decisive rules to mitigate either type of conflict.
-Given that competitiveness has a constant sum in the game, the pursuit of competitiveness becomes the primary reason for initiating a trade war or anti-globalization action, while comparative advantages promote globalization and trade collaboration.
-This study focuses on import and export activities, considering that geopolitical factors may already be incorporated into trade behaviors and volumes.
-Given the rationality of importers and exporters, trade data largely reflect political, territorial, ideological, cultural, war, national security, and other geopolitical considerations, alongside transaction and shipping costs, resource endowment, industrial distribution, and location advantages.
-From this perspective, a country’s primary objective in engaging in trade wars or de-globalization is to enhance its trade balance and competitiveness, as indicated by the data-driven social welfare function.
-The objective of this paper is to offer strategic solutions for policymakers to enhance national competitive strength, rather than focusing solely on the competitiveness of individual firms.
-Government policymakers, rather than trade companies, are responsible for improving national competitiveness, as trade companies often exploit comparative advantages.
-A nation is not merely a collection of firms, and the interests of trade companies do not always align with those of policymakers, who must also consider the negative impacts of trade, such as the contraction of domestic labor markets and the relocation of industries overseas.
-Therefore, our study disregards production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead maximizes the social welfare function, emphasizing domestic production decision-making by global consumers.
-However, the measure of competitiveness is also a weighted average of comparative advantages, weighted by the preferences of consumers, whether affluent or impoverished.
-When a country initiates a trade war against a trade partner, it raises tariffs or imposes other barriers to enhance competitiveness.
-However, the outcome depends on the partner’s response.
-We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in conflict or cooperation.
-Additionally, the reaction may benchmark their production ratio to avoid trade deficits.
-Utilizing the threshold and benchmark ratios, we derive Nash bargaining solutions (1950), generating consistent national bargaining power and providing methods to resolve bilateral trade frictions.
-A country with no specific targets to confront could choose between further globalization or protectionism.
-By making minor adjustments to its overall exports, we calibrate the effect on its competitiveness.
-The sign of the effect indicates whether the country should pursue further globalization.
-The analyses of trade wars and globalization also shed light on forming multilateral trade agreements.
-The solutions are data-driven, relying solely on trade and production data.
-For a specific context with a set of fair assumptions, many researchers have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005).
-There are also numerous debates over economic globalization in the literature, with various detailed viewpoints documented in Stiglitz (2002, 2017).
-Different perspectives, assumptions, and considerations can lead to opposing conclusions about globalization for a country.
-Our data-driven approach disregards arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security.
-Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them.
-Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time.
-We test our new theories by analyzing ComTrade data from the United Nations (2021) for the years 2000 through 2019, comparing theory-derived results with major economic events during the same period.
-Our findings suggest that protectionism could have been a more advantageous strategy for the USA and the United Kingdom (UK), but a less favorable option for China, Germany, Japan, and Russia, particularly if their objectives were centered on competitiveness.
-Data from 2017 indicate that trade wars might be a viable strategy for the USA to maintain its competitiveness and reduce its trade deficits, and we assess the impacts of the China–USA trade war in 2018 and 2019.
-However, we also observe significant adverse side effects on third parties.
-According to our theorem of impossibility trilemma, the USA must reduce imports if it aims to lower its national debt by generating trade surpluses while preserving its competitiveness.
-Our contribution to the trade literature lies in introducing a novel strategic and comprehensive framework for analyzing cooperation and competition within the global economic system.
-Our analyses eschew unnecessary assumptions and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015).
-Unlike traditional studies, our approach does not rely on unknown parameters and formulates a quantitative measure of competitiveness for each country using the balance of power.
-This measure encompasses overall strength, including both direct and indirect influences.
-The study also derives national bargaining powers and offers strategic solutions for policymakers to manage trade relationships effectively.
-In the following exposition, we first introduce the competitiveness objective in Section 2.
-Section 3 formulates a necessary condition for a country to initiate a trade war or form a trade partnership to optimize its objectives.
-Section 4 derives Nash bargaining solutions to resolve bilateral conflicts or share profits.
-Section 5 discusses a similar condition for globalization or protectionism.
-In an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019.
-Finally, Section 7 discusses several extensions and limitations of this framework.
-Our exposition is self-contained, and the proofs are in the Appendix.
+Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production could be maximized if all factors of production moved freely.
+Yet, prosperity remains uneven: some nations flourish while others falter.
+Even historically dominant economies have faced setbacks, underscoring the importance of fair competition.
+To navigate this landscape, nations must strike a balance between competition and cooperation — securing their fair share while contributing to global output.
+Both dynamics are essential.
+This paper proposes strategic and fair solutions for countries engaged in international trade. We identify three key challenges:
+  1. 1.
+    The Dual Nature of Trade: International trade embodies both conflict and cooperation.
+While voluntary exchange reflects collaboration, a country’s growing strength can undermine its partners.
+Models that are purely cooperative or non-cooperative fail to capture this complexity.
+  2. 2.
+    System-Level Representation: A framework that emphasizes competitiveness is crucial, as trade flows already reflect comparative advantages.
+We conceptualize international trade as a directional network and aim to formulate utility functions that integrate both conflict and cooperation.
+Policy variables — such as imports and exports as shares of production — should align with these objectives, enabling countries to optimize their strategies.
+  3. 3.
+    Data-Driven Insights: Real-world data can illuminate past events and guide optimal strategies for the near future — such as selecting trade partners, resolving disputes, and determining the appropriate degree of globalization.
+We model international trade as a network game, where countries are nodes and trade flows are directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016).
+In this multilateral setting, unilateral actions often lead to unpredictable or counterproductive outcomes.
+Trade wars and anti-globalization measures — though intended to protect domestic industries or reduce deficits — frequently cause more harm than good (e.g., Chung et al., 2016; Read, 2005).
+A realistic approach requires mutual understanding of each partner’s strengths and weaknesses, as reflected in domestic production and trade flows.
+These flows, along with domestic consumption, form a coalitional game for each nation.
+The Shapley value (1953) of the game provides a revealed preference or individual utility function (Roth, 1977) and serves as a power index in completing domestic production.
+The commonly shared utility function — referred to as the social welfare function — aggregates individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987) and balances power within the network (Hu and Shapley, 2003).
+It evaluates national competitiveness and captures the dual dynamics of cooperation and competition, thereby providing a battleground for both across nations.
+Competitiveness generates additional welfare beyond economies of scale — a phenomenon often described as the Matthew effect — and offers further incentive for collaboration, complementing traditional comparative advantages.
+While comparative advantage refers to producing goods or services at a lower opportunity cost, thereby promoting specialization and inclusive growth, competitive advantage emphasizes strategic management aimed at achieving national dominance, fostering exclusive growth.
+We adopt a unified game-theoretic framework to analyze two distinct forms of economic conflict: deglobalization and trade wars.
+Deglobalization refers to a country’s withdrawal from the global economy, whereas trade wars involve direct confrontation between trading partners.
+Both phenomena can significantly disrupt global economic activity.
+This paper proposes clear and practical strategies to mitigate these conflicts.
+In our framework, national competitiveness is modeled as a constant-sum game, suggesting that its pursuit often fuels trade wars and anti-globalization measures.
+In contract, comparative advantage promotes globalization and trade cooperation.
+Our approach centers on trade flows, recognizing that geopolitical factors — such as political, territorial, ideological, cultural, and security considerations — are inherently embedded in trade data.
+These data also reflect transaction costs, resource endowments, industrial structures, and locational advantages.
+From this perspective, a country’s primary objective in trade wars or deglobalization is to improve its trade balance and competitiveness.
+The goal of this paper is to offer strategic guidance for policymakers to enhance national competitiveness, rather than focusing on firm-level performance.
+Unlike trade firms, which primarily exploit comparative advantages, nations are not mere aggregations of firms.
+Policymakers must also address the adverse effects of trade, such as labor market disruptions and manufacturing offshoring.
+Accordingly, this study moves beyond production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead seeks to maximize a social welfare function that emphasizes domestic production decisions shaped by global demand.
+This function incorporates comparative advantages, weighted by consumer preferences across nations, products, and income levels.
+When a country initiates a trade war, it typically raises tariffs or imposes barriers to boost competitiveness.
+The outcome depends on the partner’s response.
+We find that a simple threshold — the ratio of competitiveness — determines whether conflict or cooperation is optimal.
+Additionally, responses may benchmark production ratios to avoid trade deficits.
+Using these threshold and benchmark ratios, we derive Nash bargaining solutions (1950) for fair responses.
+These solutions establish consistent measures of national bargaining power and offer practical methods for resolving bilateral trade frictions.
+For countries without explicit confrontation targets, the strategic choice lies between globalization and protectionism.
+By adjusting overall exports, we can assess the impact on competitiveness; the sign of this effect indicates whether deeper globalization is beneficial.
+Additionally, our analyses of trade wars, globalization, and bargaining power all help form and reform multilateral trade agreements.
+The proposed solutions are data-driven, relying exclusively on trade and production data.
+While prior research has examined bilateral or multilateral negotiations under specific assumptions (e.g., Bagwell et al., 2020; Schneider, 2005), and globalization has been extensively debated (e.g., Stiglitz, 2002 and 2017), these discussions often yield opposing conclusions depending on assumptions and priorities.
+In contrast, our data-driven approach sidesteps normative debates on labor markets, supply chains, environmental concerns, and national security — factors that trade data may already implicitly capture.
+Ultimately, whether a country embraces further globalization or protectionism depends on its unique circumstances at a given point of time.
+Using UN ComTrade data (2021), we test our theories against major economic events from 2000 to 2019.
+Results suggest that protectionism could have benefited the United States and the United Kingdom, but not China, Germany, Japan, or Russia, when competitiveness was the primary goal.
+Data from 2017 indicate that trade wars might have helped the USA maintain competitiveness and reduce deficits.
+We also assess the China–USA trade war in 2018 and 2019, noting significant spillover effects on third-party economies.
+Our impossibility trilemma shows that the USA should reduce imports if it aims to lower national debt by generating trade surpluses while preserving competitiveness.
+This study contributes to the trade literature by introducing a strategic, comprehensive framework for analyzing cooperation and competition in global trade.
+Unlike traditional studies, it avoids unknown parameters, unnecessary assumptions, and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015).
+The remainder of this paper is organized as follows.
+We introduce the revealed social welfare function in Section 2.
+Section 3 formulates conditions for a country to initiate a trade war or form a trade partnership to optimize its competitiveness.
+Section 4 derives Nash bargaining solutions for resolving bilateral conflicts and sharing profits.
+Section 5 discusses conditions for choosing globalization or protectionism.
+Using trade data from 2000 to 2019, Section 6 presents an empirical analysis of historical events.
+Finally, Section 7 concludes with policy implications, extensions, and limitations of the framework.
+All proofs are provided in the Appendix, and the exposition is self-contained.
 2 A Network-based Social Welfare Function
-This section introduces a novel type of competitiveness and social welfare function for countries within the trade network.
-Competitiveness is a relative measure derived from a constant-sum game.
-Absolute production and total trade volume may provide limited information regarding an economy’s competitive position.
-The game simulates the inflows and outflows within the network; without change, these steady flows reach a static equilibrium due to long-term accumulation.
-The equilibrium represents the competitiveness vector, which describes the utilities of the countries’ positions in the network game.
-We detail the steps toward the data-based social welfare function and its unique properties, which are unfamiliar in the trade literature.
-Domestic production is paramount for policymakers, whereas profit maximization is crucial for firms or producers.
-Production with zero or slightly negative profit may be insignificant to firm owners but can still be an option for the government to boost employment.
-For example, production maximization results in less unemployment, while labor may be resourced overseas to maximize shareholders’ profits.
-Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors and the distribution of profit to the firms.
-Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
-For countries in the network, we describe their aggregate trade activities with a square matrix $P$.
-Assume there are $n$ countries, labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$.
-For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ be the fraction of country $i$’s production that exports to country $j$.
-The exports include all services and goods, including intermediate goods; they may contain intermediate goods made by third parties or country $j$.
-Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as all domestically produced final products.
-Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are directly attributable to the production, such as raw materials, intermediate goods, and crude oil.
-It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$.
-Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP.
-Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$.
-We use $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ for the non-exporting fraction.
-Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $i\in\mathcal{N}$.
-We place all the fractions $P_{ij}$ into an $n\times n$ stochastic matrix $P=[P_{ij}]$.
-Each row of $P$ defines an individual utility function that determines a specific country’s production from a demand- or consumption-side perspective.
-For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
+This section introduces a novel concept of competitiveness, which is a social welfare function for countries within the global trade network.
+Competitiveness is defined as a relative measure derived from a constant-sum game.
+Unlike absolute production or total trade volume, which provide limited insight into an economy’s competitive position, this measure captures the strategic dynamics of trade.
+The game employs inflows and outflows within the network.
+In the absence of changes, these flows converge to a static equilibrium over time, reflecting long-term accumulation.
+This equilibrium yields a competitiveness vector, which represents the utilities of countries’ positions in the network game.
+We outline the steps for constructing this data-driven social welfare function and highlight its unique properties, which are largely unfamiliar in the trade literature.
+For policymakers, domestic production is a primary concern, whereas firms prioritize profit maximization.
+Production that yields zero or slightly negative profit may be unattractive to firms but can still be a viable policy option for governments seeking to boost employment.
+For example, maximizing production can reduce unemployment, while firms might prefer outsourcing labor overseas to maximize shareholder returns.
+Accordingly, our objective function and strategic responses focus on production decision-making, leaving the allocation of production factors and profit distribution to firms.
+Governments, however, play a dual role: regulating profit-maximization behavior and enhancing consumer welfare, while redistributing firm profits through taxation.
+To model trade relationships, we represent aggregate trade activities using a square matrix $P$.
+Assume there are $n$ countries, labeled $1,2,\cdots,n$, collectively denoted by the set $\mathcal{N}=\{1,2,\cdots,n\}$.
+For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ denote the fraction of country $i$’s production exported to country $j$.
+Exports include all goods and services, including intermediate goods, which may incorporate components from third countries or from country $j$ itself.
+The production value of country $i$, denoted $g_{i}$, consists of all types of exports, all domestically produced final products, and imports directly attributable to production (e.g., raw materials, intermediate goods, crude oil).
+However, $g_{i}$ excludes imported final goods not used in domestic production.
+We define $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ as the non-exporting fraction.
+Clearly, $P_{ij}\geq 0$ for all $i,j\in\mathcal{N}$, and for each $i\in\mathcal{N}$, $\sum\limits_{j=1}^{n}P_{ij}=1$.
+All fractions $P_{ij}$ form an $n\times n$ stochastic matrix $P=[P_{ij}]$.
+Each row of the matrix $P$ defines an individual utility function that breaks down a specific country’s production from either a demand-sided or consumption-side perspective.
+For any subset $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
 $v_{i}(S)\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{j\in S}P_{ij},$ (1)
 with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$.
-Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$.
-Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$.
-Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern utility function (1953), i.e., $P_{ij}$ measures an appraisal agent’s satisfaction when they play country $j$’s role in the game of $(\mathcal{N},v_{i})$.
-The agent is neutral with respect to ordinary and strategic risks detailed in Roth (1977).
-Thus, countries have cardinal preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$.
-Yet, their preferences change as $i$ varies in $\mathcal{N}$.
-In the matrix $P$, therefore, each row is an individual utility function, whereas column countries are the arguments or social states of the function.
-A social welfare function generally combines these individual utility functions into one, accounting for the diverse value judgments of countries in $\mathcal{N}$.
-It rates the set of countries in $\mathcal{N}$ as a whole and serves as a common platform for the countries to compete and collaborate.
+Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ represents the value of production driven by the countries in $S$.
+Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is given by the $i$th row of $P$, i.e., $(P_{i1},P_{i2},\cdots,P_{in})$.
+Roth (1977) demonstrates that the Shapley value is a von Neumann-Morgenstern utility function (1953).
+In this context, $P_{ij}$ measures an appraisal agent’s satisfaction when playing the role of country $j$ in the game of $(\mathcal{N},v_{i})$, assuming neutrality toward both ordinary and strategic risks (Roth, 1977).
+In the matrix $P$, therefore, rows represent individual utility functions of countries and columns represent the arguments or social states of these functions.
+Consequently, countries exhibit cardinal preferences in the game: $P_{ij}>P_{ik}$ implies that country $j$ is preferred to $k$.
+However, these preferences vary with $i\in\mathcal{N}$.
+A social welfare function aggregates these individual utility functions into a single function, incorporating the diverse value judgments of countries in $\mathcal{N}$.
+It evaluates the set of countries as a whole and serves as a common platform for competition and cooperation.
 2.1 Long-Run Influence in the Trade System
-With the normalization (i.e., $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$) and monotonicity of $v_{i}$, $P_{ij}$ bears much similarity to the Shapley-Shubik power index (1954), measuring country $j$’s chance of pivoting in forming the result $v_{i}(\mathcal{N})$.
-Indeed, $P_{ij}$ is a type of pro rata power when country $i$ makes its full production.
-The multilinear extension of $v_{i}$ is a function defined by:
-$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
-where $x_{j}\in[0,1]$ is the probability with which country $j$ participates in the consumption-driven production.
-Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$.
-When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production.
-Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also represents $j$’s progress in completing its imports from $i$.
-Additionally, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$.
-Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$.
-For example, if country $j$ completes 25% of its consumption, all other countries also finish their 25% at the same time.
-Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s overall influence or power that propels country $i$ to complete the production.
-According to Owen (1972), the integral equals the Shapley value for the coalitional game, i.e.,
+With the normalization conditions $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$, and the monotonicity of $v_{i}$, the measure $P_{ij}$ closely resembles the Shapley-Shubik power index (1954), which quantifies country $j$’s pivotal role in determining the outcome $v_{i}(\mathcal{N})$.
+In essence, $P_{ij}$ represents a form of pro rata power when country $i$ realizes its full production.
+The multilinear extension of $v_{i}$ is defined as:
+$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S).$
+Here, $x_{j}\in[0,1]$ denotes the probability that country $j$ participates in the consumption-driven production.
+When restricted to the corners of the unit cube $0,1]^{n}$, $\tilde{v}_{i}$ simplifies to the original function $v_{i}$.
+Under the assumption of independent actions, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ represents the expected production of country $i$.
+As production evolves from zero at the origin $(0,0,\cdots,0)$ to full completion at $(1,1,\cdots,1)$, each $x_{j}$ also reflects country $j$’s progress in completing its imports from country $i$.
+The partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ captures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at a given point.
+In the absence of prior knowledge about participation probabilities or progress, we assume uniform advancement along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$.
+For instance, if country $j$ completes 25% of its consumption, all other countries are assumed to have completed 25% as well.
+Consequently, the integral of the gradient of $\tilde{v}_{i}$, denoted $\bigtriangledown\tilde{v}_{i}$, along this diagonal quantifies each country’s overall influence in driving country $i$’s production to completion.
+According to Owen (1972), this integral equals the Shapley value for the coalitional game:
 $\int_{0}^{1}\bigtriangledown\tilde{v}_{i}(x,\cdots,x)\mathrm{d}x=(P_{i1},P_{i2},\cdots,P_{in})^{\top}$
-where $\top$” denotes vector or matrix transpose. Therefore, $P$ contains $n\times n$ quantitative powers, and the balance of power in $P$ refers to a state of stability.
-From a supply-side perspective, inputs to the production function in Eq.
-(1) can include any variables that affect the production and services of country $i$.
-However, it is impossible to compile a complete list of these variables.
-Their effects are also time-varying and country-specific; each product has a unique way of combining them.
-For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products.
-These assets are primarily payments for past, current, or future exports.
-Trade involves exchanging goods and services; and monetary flow as the medium of exchange makes transactions possible in modern international trade.
-Nevertheless, payments alone cannot fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only.
-Zero payments may occur for intrafirm transactions of exports within a multinational company.
-Additionally, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively.
-A country can accumulate trade deficits if it is a highly desirable destination for foreign investment.
-For example, U.S.
-Treasury Bonds are lucrative and risk-free investment instruments after the Cold War, so foreign governments or companies must earn extra U.S.
-dollars from their trades to add to their foreign investment portfolios.
-The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system.
-In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere.
-When China exports a smartphone to the USA, for example, the chip and design software of the phone could come from the USA, and the camera from Korea.
-Japan could manufacture the camera screen using US patents.
-To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$.
-In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world.
-For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
-$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
-where $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$.
-It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs through all countries in $\mathcal{N}$.
-Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely.
-These conditions are obviously satisfied in the closed and well-connected trade system.
-Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system.
-The direct influences in $P$ pass through interlinked global value chains and there are millions of them.
-Consequently, $P^{t}$ converges only if $t$ is large enough.
-The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012).
-The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.
+where $\top$” denotes vector or matrix transpose.
+Therefore, this matrix $P$ contains $n\times n$ quantitative measures of power, and the balance of power in $P$ indicates a stable state in the trade system.
+From a supply-side perspective, the inputs to the production function in Eq.
+(1) may include any variables that influence the production and services of country $i$.
+However, compiling a comprehensive list of these variables is impractical.
+Their effects vary over time and across countries, and each product integrates them in unique ways.
+For example, inflows of financial assets — such as remittances and foreign direct investments — play a significant role in domestic production.
+These assets often represent payments for past, current, or future exports.
+Trade involves the exchange of goods and services; and monetary flows facilitate these transactions in modern international trade.
+Yet, payments alone do not fully capture trade activities.
+Transactions between two firms may be partially offset, resulting in net payments only.
+Intra-firm trade within multinational corporations may involve zero payments.
+Moreover, trade surpluses and deficits can be influenced by currency manipulation — depreciation and appreciation, respectively.
+A country may accumulate trade deficits if it is a highly attractive destination for foreign investment.
+For instance, U.S.
+Treasury Bonds became lucrative and low-risk investment instruments after the Cold War.
+As a result, foreign governments or companies often seek to earn additional U.S.
+dollars through trade to expand their investment portfolios.
+In a closed global trade system, direct influences captured in matrix $P$ naturally give rise to indirect ones .
+Within a production chain, even if the final product is labeled as originating from a single country, its components often come from multiple sources.
+For instance, when China exports a smartphone to the USA, the chip and design software may be sourced from the USA, the camera from Korea, and the camera screen manufactured in Japan using US patents.
+To account for these spillovers, we consider both indirect and long-term impacts through successive powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$.
+In general, $P^{t}$ represents the aggregate $t$-step impacts across all supply chains in the globalized economy.
+Specifically, the entry at the $i$th row and $j$th column of $P^{t}$, denoted $\left(P^{t}\right)_{ij}$, captures all paths of length $t$ from country $i$ to country $j$:
+$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}.$ (2)
+Here, each term $P_{k_{s},k_{s+1}}$ represents the spillover from country $k_{s}$ to $k_{s+1}$, or equivalently, the direct influence of $k_{s+1}$ on $k_{s}$, for $s=1,2,\cdots,t-1$.
+These interactions span all countries in the set $\mathcal{N}$.
+Under general conditions — specifically, aperiodicity and irreducibility — the rows of $P^{\infty}$ converge to a unique constant row vector $\pi$.
+These conditions are naturally satisfied in our trade system.
+The element $\pi_{j}$ is defined as: $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ and is independent of the choice of $i$, thereby measuring country $j$’s long-run influence on the entire trade system.
+The existence and uniqueness of $\pi$ are well-established in the theory of Markov chains (e.g., Karlin and Taylor, 2012).
+Given the vast number of interlinked global value chains, convergence of $P^{t}$ requires a sufficiently large $t$.
+Moreover, the limit of $P^{t}$ offers a practical computational approach to determine $\pi$.
 2.2 Mixed Cooperation and Noncooperation
-The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also solves the following counterbalance equilibrium:
+The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also satisfies the following counterbalance equilibrium:
 $\pi=\pi P$ (3)
 subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$.
-Numerous counterbalanced systems exist in both the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
-The counterbalance in Eq.
-(3) implies $\pi$’s mixed cooperative and noncooperative properties.
-This is a distinctive feature between comparative and competitive advantages; while both promote collaborations, competitive advantage may bring trade wars and deglobalization.
-For instance, between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years.
-Although unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147).
-Econometric analysis often struggles to capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
-From a non-cooperative perspective, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$.
-Thus, theoretically, there are $n(n-1)/2$ potential trade conflicts, either small or large, within the trade system.
-Most of these are minor trade disagreements or disputes, making the “war” metaphor hyperbolic in these cases.
-However, it is worth mitigating even minor disputes before they become problematic or catastrophic.
-Additionally, the increase in $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initiated by changes in the matrix $P$.
-In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively, for bilateral trade wars and globalization.
-From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, as $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in:
+Counterbalanced systems are prevalent in both the physical and human domains — examples include ecological systems, the U.S.
+government structure, and China’s five-element theory.
+The counterbalance expressed in Eq.
+(3) reflects the mixed cooperative and noncooperative nature of $\pi$.
+This duality distinguishes comparative advantage from competitive advantage: while both encourage collaboration, competitive advantage can also lead to trade wars and deglobalization.
+For example, in the trade relationship between China and the United States from 1980 to 2020, the first thirty years were marked by greater cooperation than the last ten.
+Although this mixture of cooperation and noncooperation is less familiar to empirical macroeconomics, it may offer valuable insights moving forward (e.g., Allen, 2000, page 147).
+Also, econometric models often struggles to capture the unity of opposites, as estimated coefficients or effects cannot simultaneously be significantly positive and significantly negative.
+From a noncooperative standpoint, given that $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$.
+Theoretically, this leads to $\dfrac{n(n-1)}{2}$ potential trade conflicts — ranging from minor disagreements to major disputes — within the trade system.
+Nonetheless, even minor disputes are worth addressing before they escalate into serious or catastrophic issues.
+Moreover, the increase in $\pi_{j}$ is not uniformly offset by reductions in other economies.
+These shifts often follow shared causal patterns triggered by changes in the matrix $P$.
+In Sections 3 and 5, we differentiate between the first move off the diagonal and the move along the diagonal of $P$, corresponding to bilateral trade wars and globalization, respectively.
+From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, because $\pi_{j}P_{ji}$ contributes to $\pi_{i}$, as shown in:
 $\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}P_{ji}$ (4)
-from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t},$ where
+from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ where:
 $\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}\left(P^{t}\right)_{ji}$ (5)
-for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$.
+for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, it follows that $\left(P^{t}\right)_{ji}>0$ for all sufficiently large $t$ and for all $j\in\mathcal{N}$.
 Thus, Eq.
-(5) further implies that country $i$ should assist all other countries in the trade system, including the poorest and least competitive ones, to enhance their $\pi_{j}$.
-Hence, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$.
-The trade-off is determining how much it should assist or contest $j$ without sacrificing its national interests and potentially increasing $\pi_{i}$.
+(5) further implies that country $i$ should assist all other countries in the trade system — including the poorest and least competitive ones — to enhance their $\pi_{j}$.
+Therefore, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$.
+The key trade-off lies in determining how much assistance or contestation to provide to $j$ without sacrificing national interests while potentially increasing $\pi_{i}$.
 […]
-Figure I: Dynamics of authority flow for country $i$ and $j$.
-We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary driving forces for the flows.
+Figure I: Dynamics of authority flow for country $i$ and $j$
+We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary forces driving the flows.
 In the inflow equation Eq.
 (4), country $i$ accumulates or absorbs authority from other countries through its direct influences.
-It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, all else being equal.
-Additionally, it derives more authority from countries on which it has a significant direct influence (i.e., $P_{ji}$ is large), all else being equal.
-In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$, and thus, authority flows from $j$ to $i$ along the same column.
-Conversely, country $i$ also contributes to other countries, as illustrated in the following outflow equation:
+It gains more authority from influential trade partners (i.e., when $\pi_{j}$ is large) than from less influential ones, all else being equal.
+Additionally, it derives more authority from countries on which it exerts significant direct influence (i.e., when $P_{ji}$ is large), again holding other factors constant.
+As shown in Figure I, country $i$ imports goods from $j$ through the right column, and thus authority flows from $j$ to $i$ along the same column.
+Conversely, country $i$ also contributes to other countries, as illustrated in the outflow equation:
 $\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\neq i}\pi_{k}P_{kj}.$
 The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal.
-In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
-In the accounting of $\pi$, the two countries reach a break-even point when:
-$\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
-Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either takes more authority from or gives more to country $j$, respectively.
-For country $i$, a straightforward implication of Eq.
-(6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively.
+In Figure I, country $i$ redistributes $\pi_{i}$ through the bottom row, according to its export shares.
+These two countries reach a break-even point when:
+$\pi_{i}P_{ij}=\pi_{j}P_{ji}$ (6)
+where $\pi_{i}P_{ij}$ and $\pi_{j}P_{ji}$ represent the authority flows from $i$ to $j$ and from $j$ to $i$, respectively.
+Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either gains more authority from or gives more to country $j$, respectively.
+For country $i$, a straightforward implication is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively.
 It can also adjust its exports to $j$ accordingly.
-However, the outflow from $i$ to $j$ does not necessarily match the inflow.
-With the involvements of third parties, the inflows and outflows meet at the equilibrium Eq.
-(3) for $i$, i.e.
+However, the outflow from $i$ to $j$ does not necessarily match the inflow. With the involvement of third parties, inflows and outflows meet at the equilibrium described by Eq. (3) for $i$:
 $\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}.$
-For the same reason, it could be overly restrictive for all countries to balance their bilateral trades.
-For example, country $i$ can still maintain an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
+For the same reason, requiring all countries to balance their bilateral trades would be overly restrictive.
+For example, country $i$ can still maintain an overall zero net balance if it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
 Finally, Eq.
-(3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$.
-It represents a type of eigenvector centrality, a subclass of network centrality.
+(3) implies that $\pi$ is the unit-sum row eigenvector corresponding to the largest row eigenvalue of $P$.
+This represents a form of eigenvector centrality, a subclass of network centrality.
 In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987).
-Our $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market, and the eigenvector $\pi$ measures competitiveness (see Section 2.3).
-The same domestic products are used both domestically or overseas.
-Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal.
-Furthermore, $P$ normalizes the scales of economies such that each row has a unit sum.
+Our matrix $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market and the same domestic products are used both domestically and overseas.
+Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with others that also accept the student and are off the diagonal.
+Furthermore, $P$ normalizes the scales of economies such that each row sums to one.
 Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
-2.3 $\pi$ as Competitive Advantages
-The competitive advantage of country $i$ refers to its ability to outperform its competitors within the trade system.
-For any product exported from country $i$ to country $j$, the competitors of country $i$ include all other countries.
-Thus, its competitive advantage extends from bilateral country relations to a relationship with the entire system — it is the ability to surpass all other countries in the $j$th market for any $j\in\mathcal{N}$.
-In contrast, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over the trade partner’s opportunity cost; it is a bilateral relationship.
-Therefore, globalization may have made competitive advantages more prevalent than comparative advantages.
-According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, land, market affordability, and low cost of capital.
-These factors, imported from outside or manufactured domestically, are assembled in the final products.
-Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker.
-In the smartphone example, the competitive factors include the chips and patents from the USA, the labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea.
-Various studies have also described qualitative strategies to enhance firm competitiveness.
-For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
-We find that $\pi$ quantifies national competitiveness.
-To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus.
+2.3 $\pi$ as a Measure of Competitive Advantage
+The competitive advantage of country $i$ refers to its ability to outperform rivals within the global trade system.
+For any product exported from country $i$ to country $j$, the competitors of $i$ include all other countries.
+Thus, competitive advantage extends from bilateral relations to encompass the entire system — it reflects the ability to surpass all other countries in market $j$ for any $j\in\mathcal{N}$.
+In contrast, comparative advantage is based on an economy’s specialization, resource endowment, and technology relative to its trade partner’s opportunity cost, making it inherently bilateral.
+Globalization, however, may have amplified the importance of competitive advantages over comparative advantages.
+According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantage rests on direct control of key production factors, such as access to rare resources, advanced technology, low-cost but highly skilled labor, affordable markets, and inexpensive capital.
+These factors — whether imported or domestically produced — are integrated into final products.
+Consequently, imported goods carry the original producer’s competitiveness and should be attributed to their source.
+For example, in smartphone production mentioned before, competitive factors include U.S.
+chips and patents, Chinese labor, Japanese precision manufacturing, Korean camera technology.
+Various studies also outline qualitative strategies to enhance firm competitiveness.
+Porter (1985), for instance, identifies three generic strategies: cost leadership, differentiation, and focus.
+We propose that $\pi$ quantifies national competitiveness.
+To introduce this quantitative measure, we temporarily ignore home bias in trade (cf., McCallum, 1995) and assume no bilateral trade surplus or deficit.
 Denote $\tau_{j}$ as the $j$th row of $P$.
-Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production.
-So, it measures other countries’ aggregate comparative advantages over country $j$.
-Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means solid competitiveness to occupy the $j$th market and win a lion’s share of $j$’s imports.
-If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum.
-Yet, the $j$th value is not excluded from $\tau_{j}$, because domestic products also compete with country $j$’s imports when consumers make their consumption choices, and they have no home bias.
-We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
-These individual cardinal utilities are not comparable because it is more challenging to occupy a competitive market than a less competitive one.
-We make them comparable by weighting, and the weights are not exogenously assigned but endogenously determined.
-More weight should be placed on a more competitive reference market than on a less competitive one.
-Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$.
-Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$.
-Both the weighted score and the weight vector quantify the competitiveness of the countries.
-To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$.
-By Theorem 1, $\rho$ is just $\pi$.
+Under the assumption of no bilateral surplus or deficit, $\tau_{j}$ — excluding its $j$th element — represents imports from other countries into $j$, expressed as fractions of $j$’s production.
+This measures other countries’ aggregate comparative advantages over $j$.
+Moreover, all countries compete for comparative advantages in market $j$.
+A large value in $\tau_{j}$ indicates strong competitiveness in capturing a significant share of $j$’s imports.
+If we use market $j$ as the reference point for ranking all countries’ competitiveness, the resulting score is the vector $\tau_{j}$, which sums to one.
+The $j$th element is included because domestic products also compete with imports when consumers make purchasing decisions, assuming no home bias.
+We apply endogenous weighting to the reference scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
+These individual cardinal utilities are not directly comparable because operating in a highly competitive market is inherently more challenging than in a less competitive one.
+To make them comparable, we introduce weights that are endogenously determined rather than exogenously assigned.
+Greater weight should be given to more competitive reference markets.
+Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$, where $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. The weighted competitiveness score is:
+$\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P.$
+Both the weighted score and the weight vector quantify national competitiveness.
+For consistency, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$.
+By Theorem 1, $\rho$ equals $\pi$.
 Theorem 1
-If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
-In the endogenous weighting, worldwide consumers unconsciously participate in evaluating national competitiveness, encompassing all products and services.
-Essentially, countries as a whole unknowingly assess themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats.
-The structure of the global trade network is largely determined endogenously by the distribution of firm productivity within individual countries.
-Nevertheless, government policies can influence this distribution without significantly affecting the welfare of international trade by shifting imports or exports between countries.
-Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing heavy tariffs on foreign products.
-However, national competitiveness $\pi_{i}$ is not merely the sum of the individual competitiveness of domestic firms; firm competitiveness are weighted in $\pi$ by consumers, wealthy or poor, domestic or overseas.
-Affordability thus plays a vital role in $\pi$.
-The positive weights imply a positive correlation between firm and national competitiveness.
-Additionally, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
-Thus, the weighted utility function $\pi$ is determined not by the consumption of production factors or the added value, but by consumer preferences, judged independently of national interests or preferences.
-Pareto optimality is achieved when weighting these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$.
-The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant, such as scaling $\tau_{i}$ by $g_{i}$.
-Harsanyi (1955) and Hammond (1987) provide additional ethical, risk, and utilitarianism connotations for this type of social welfare function.
-With the weights being $\pi$ itself, $\pi$ exhibits some special properties, discussed in Section 2.4.
-However, the social welfare function $\pi$ significantly departs from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014).
-It is not explicitly expressed in a specific functional form of production factors and total factor productivity but implied from the counterbalance equilibrium, as indirect imports also impact the final products or services.
-Exported goods and services have different utilities for exporting and importing countries.
-Therefore, each country has a unique utility function, and a single utility function cannot capture the diversity of all countries.
-We use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages production factors into final products and how differently it trades the products with others.
-It is precisely derived from the data and uses no parametric models, thus no unknown parameters need to be estimated, and no residuals exist to differentiate the data from a parametric model.
-Additionally, according to Barney and Felin (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors.
-Due to broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges.
-With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network.
-The index $\pi_{i}$ contains a unique component, $\pi_{i}P_{ii}$, which does not originate from external source, as indicated by Eq.
+If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and therefore $\rho=\pi$.
+In this endogenous weighting framework, global consumers unconsciously participate in evaluating national competitiveness across all products and services.
+Essentially, countries collectively assess themselves without external influence from business interests, media, advertising, governments, or military pressure.
+The structure of the global trade network is largely shaped by the distribution of firm productivity within individual countries.
+However, government policies can influence this distribution — without significantly affecting international trade welfare — by shifting imports or exports between countries (e.g., Gopinath et al., 2024).
+Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing tariffs on foreign products.
+Importantly, national competitiveness $\pi_{i}$ is not simply the sum of domestic firms’ competitiveness.
+Instead, firm competitiveness is weighted in $\pi$ by consumer preferences — wealthy or poor, domestic or overseas.
+Affordability thus plays a critical role in determining $\pi$.
+Positive weights imply a positive correlation between firm-level and national competitiveness.
+Moreover, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as it is a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$.
+This weighted utility function $\pi$ is determined not by factor consumption or added value, but by consumer preferences, independent of national interests.
+Pareto optimality is achieved because:
+$\pi_{j}\geq\pi_{k}\ \mathrm{if}\ P_{ij}\geq P_{ik}\ \mathrm{for}\ \mathrm{all}\ i\in\mathcal{N}.$
+This optimality remains invariant if any individual utility function is scaled or shifted by a positive constant (e.g., scaling $\tau_{i}$ by $g_{i}$).
+Harsanyi (1955) and Hammond (1987) provide further ethical, risk, and utilitarian interpretations of this type of social welfare function.
+This social welfare function $\pi$ differs significantly from widely used utility or welfare functions in the literature on trade wars and globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014).
+Unlike traditional approaches that specify a functional form based on production factors and total factor productivity, our formulation is implicitly derived from the counterbalance equilibrium, accounting for the indirect effects of imports on final goods and services.
+Because exported goods and services yield different utilities for exporting and importing countries, each country possesses a distinct utility function.
+Consequently, a single utility function cannot capture the heterogeneity across nations.
+Our approach minimizes assumptions by disregarding how efficiently countries transform production factors into final products or how they engage in trade.
+The index $\pi$ is constructed directly from observed data without relying on parametric models.
+As a result, there are no unknown parameters to estimate and no residuals to reconcile with a model specification.
+Furthermore, as Barney and Felin (2013) note, the debate over micro-foundations remain unresolved, with scholars in management, strategy, and organization offering divergent views on the relationship between macroeconomic phenomena and micro-level behaviors.
+Given the vast heterogeneity among firms and nations, modeling micro-founded strategic interactions within a network comprising millions of product flows across tens of thousands of directional edges is particularly challenging.
+A certain degree of abstraction is therefore necessary to capture the essential features of such a complex system.
+The index $\pi_{i}$ includes a unique component, $\pi_{i}P_{ii}$, which originates entirely from domestic sources, as indicated in Eq.
 (4).
-This encompasses, but is not limited to, raw oil fields, natural gas reserves, rare minerals, and patented technologies, all of which remain integral to the nation’s core competitive factors.
-A global value chain typically involves finite steps but also includes indefinite domestic loops when tracing all the added values of the final products.
-For instance, in the smartphone example, the education, skills, and experience of manufacturing labor are ultimately developed domestically over many years, contributing to the final product.
-Additionally, numerous U.S.
-workers develop patents and design software domestically for extended periods.
-These inner loops affect other countries but not vice versa; this aspect of competitiveness never flows in from external sources.
-Consequently, social infrastructure, political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that facilitate business all contribute to national competitiveness (e.g., Delgado et al., 2012).
-For a specific global value chain, at the other end of the unique component is the design and assembly of the final product.
-Product leadership bridges the market, consumers, strategic direction, technology innovation, sales, and production factors together; this leadership represents another type of competitiveness.
-Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods.
-Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing.
-Technological advancements reduce labor costs, and public safety and social stability mitigate the risk of costs.
+This component encompasses critical resources such as oil fields, natural gas reserves, rare minerals, and patented technologies — factors that remain central to a nation’s competitive advantage.
+A global value chain typically consists of a finite sequence of steps but also includes indefinite domestic loops when tracing the value added to final products.
+For example, in the case of smartphones, the education, skills, and experience of manufacturing labor are developed domestically over many years, contributing to the final product.
+Similarly, U.S.
+workers invest significant time in developing patents and design software domestically.
+These internal loops influence other countries but are not reciprocally affected, underscoring the role of domestic competitiveness $\pi_{i}P_{ii}$.
+Elements such as social infrastructure, political institutions, monetary and fiscal policies, and the quality of the microeconomic environment all contribute to national competitiveness (e.g., Delgado et al., 2012).
+Within a specific global value chain, the opposite end of this unique component involves the design and assembly of the final product.
+Product leadership integrates market orientation, consumer preferences, strategic direction, technology innovation, sales, and production factors — representing another dimension of competitiveness.
+Between raw resources and the final product lies in a series of value-added processes for intermediate goods, where competitiveness is driven by cost efficiency, including financing costs.
+Technological progress reduces labor costs, while public safety and social stability mitigate risk-related expenses.
 Eq.
-(2) indicates a robust positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy.
-After adjusting for the sizes of $g_{i}$ in $P$, the interaction effects between competitiveness and economies of scale remain, thereby conferring some cost advantages.
-Nations export goods that are abundant and import goods that are scarce domestically.
-With significant abundance and shortage, country $i$ could have substantial nonzero elements in the $i$th column and row of the matrix $P$, resulting in considerable nonzero items summed in Eq.
+(2) reveals a strong positive correlation between $\pi_{i}$ and $g_{i}$, which reflect the depth and breadth, respectively, of a country’s economy.
+After adjusting for the magnitude of $g_{i}$ in matrix $P$, the interaction between competitiveness and economies of scale persists, offering notable cost advantages.
+Countries tend to export goods they possess in abundance and import those they lack.
+When these are significant, country $i$ will exhibit substantial nonzero elements in both the $i$th column and row of matrix $P$, contributing meaningfully to the summation in Eq.
 (2).
-A large economy also generates a high differentiation of goods and services, which offer location advantages and minimal transaction costs for domestic consumers.
-Free from tariffs and overseas shipping, these goods also possess a price advantage.
-Consequently, $P_{ii}$ is often the largest in the $i$th row of $P$, irrespective of home bias.
-Therefore, a vast economy with a comprehensively integrated industrial system (such as China and the USA) is likely, though not necessarily, to have a large value in $\pi$.
-A country with a single industrial structure is susceptible to intense competition, and its $\pi_{i}$ is fragile and volatile when competing with the USA for competitiveness.
-Hence, we use $\pi_{i}$ as an objective for country $i$ when formulating trade policies.
-It quantifies the economy’s depth and its interplay with breadth, encompassing both cooperation and non-cooperation with other nations.
-This measure is endogenously derived from the comparative advantages of all products and services.
-Given that each country has unique locations, natural and human resources, and economic systems, a single utility function that assumes one stereotype for all economies may be insufficient to capture this diversity.
-Our study acknowledges that each country, along with its trade stance with partners, is distinct.
-Each contributes inclusiveness value to the global economy, where international trade is a cornerstone.
+Large economies also foster a high degree of product and service differentiation, which provides location-specific advantages and reduces transaction costs for domestic consumers.
+These goods, being exempt from tariffs and international shipping fees, enjoy a price advantage.
+As a result, $P_{ii}$ is often the largest in the $i$th row of $P$, regardless of any home bias.
+Therefore, a large and comprehensively integrated industrial system — such as those found in China or the United States — is likely, though not guaranteed, to exhibit a high value in $\pi$.
+In contrast, countries with narrow industrial structures are more vulnerable to competitive pressures, and their $\pi_{i}$ tend to be fragile and volatile, especially when competing with the U.S.
+in terms of competitiveness.
+We thus adopt $\pi_{i}$ as a strategic objective for country $i$ in the formulation of trade policies.
+It encapsulates both the depth and breadth of the economy, reflecting cooperative and non-cooperative dynamics with other nations.
+This metric is endogenously derived from the comparative advantages across all goods and services.
+Recognizing that each country possesses unique geographic, natural, human, and institutional characteristics, we argue that a single utility function cannot adequately represent the diversity of global economies.
+Our study embraces the distinctiveness of each nation and its trade relationships, emphasizing the inclusive value each contributes to the global economy — where international trade remains a foundational pillar.
 2.4 The Matthew Effect in $\pi$
-In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness:
+In Eq. (4), country $i$’s competitiveness $\pi_{i}$ accumulates in proportion to its initial level of competitiveness:
 $\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\neq i}\pi_{j}P_{ji},$ (7)
-where the proportion $P_{ii}$ is generally larger than .7.7.7.7.
-This illustrates $\pi$’s Matthew effect of accumulated advantage, often summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies exhibit substantial competitiveness compared to less powerful ones, revealing a quadratic relationship between $\pi_{i}$ and $g_{i}$.
-For instance, the USA’s $\pi_{i}$ could be significantly higher than Japan’s, relative to their production ratio.
-The amplification also stems from the USA’s relative advantages over other countries compared to Japan’s.
-In Eq.
-(7), with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$, respectively,
-$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
-the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$.
-Alternatively, the Matthew effect is largely explained by preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among countries according to their existing competitiveness.
-The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$.
-If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$.
-Thus, it exhibits quadratic growth when $\tilde{g}_{i}$ increases linearly, and in a linear regression:
+where the self-reinforcing proportion $P_{ii}$ is typically greater than .7.7.
+This formulation illustrates the Matthew effect — the phenomenon of accumulated advantage — often summarized by the aphorism: “the rich get richer and the poor get poorer.”
+For example, the United States may exhibit a substantially higher $\pi_{i}$ than Japan, compared to their production ratio.
+This amplification arises not only from the U.S.’s internal advantages but also from its relative superiority over other countries compared to Japan.
+Specifically, in Eq.
+(7), for $i=\mathrm{USA}$ or $i=\mathrm{Japan}$:
+$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}}.\end{array}\right$.
+The Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $1>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>0$; and $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$.
+Alternatively, the Matthew effect can be interpreted through preferential attachment in endogenous weighting, where weights are distributed proportionally to existing competitiveness.
+This implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$.
+Normalizing $g_{i}$ as $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{g_{i}}{\sum_{j=1}^{n}g_{j}}$, the effect suggests that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$, after controlling for $\tilde{g}_{i}$.
+Thus, $\pi_{i}$ exhibits quadratic growth as $\tilde{g}_{i}$ increases linearly.
+In a linear regression framework:
 $\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i},$ (8)
-for some $c_{1}>0,c_{2}>0$, where $\epsilon_{i}$ is the residual term.
-Therefore, when dealing with a rapidly thriving country $i$, country $j$ should mitigate its conflicts with $i$ before the counterpart becomes too strong to compromise.
-Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say:
+for some $c_{1}>0,c_{2}>0$, and $\epsilon_{i}$ is the residual term.
+Therefore, when country $i$ is rapidly ascending, country $j$ should seek to resolve conflicts with $i$ before $i$ becomes too powerful to compromise.
+If country $i$ leverages its accelerating $\pi_{i}$ to boost production $g_{i}$, such as:
 $\tilde{g}_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i}$
-for some $c_{4}>0$, then $\tilde{g}_{i}$ would be accelerated quadratically, leading to $\pi_{i}$’s quartic growth.
-Continuing this iteration, eventually, both $\pi_{i}$ and $\tilde{g}_{i}$ would grow exponentially to their maximum capacities.
-Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
-Additionally, the welfare associated with competitive advantage leads to an agglomeration effect, clustering all countries around a select few of economic superpowers.
-According to Corollary 1, imports from a large economy are preferred over imports from a smaller economy, all else being equal.
-Consequently, the smaller economy should counter this preference by lowering its export prices.
-Meanwhile, a third party could impose different tariffs on large and small countries to ensure that the same imported products have uniform prices in the third country.
-Furthermore, the large market, characterized by numerous producers and consumers where no single entity can dominate, may be favored by the third party for its market competitiveness.
-In either scenario, a small country tends to fear the emergence of a neighboring superpower that shares similar location advantages and resource endowments.
+with $c_{4}>0$, then $\tilde{g}_{i}$ accelerates quadratically, leading to quartic growth in $\pi_{i}$.
+Iterating this process results in exponential growth of both $\pi_{i}$ and $\tilde{g}_{i}$ toward their maximum capacities.
+Historically, many empires have emerged rapidly, gaining prominence with astonishing speed.
+Moreover, the welfare associated with competitive advantage induces an agglomeration effect, clustering countries around a few economic superpowers.
+According to Corollary 1, imports from a large economy are preferred over those from a smaller one, all else being equal.
+Consequently, smaller economies may counteract this preference by lowering export prices.
+A third party might impose differentiated tariffs on large and small countries to equalize import prices.
+Additionally, large markets — characterized by numerous producers and consumers without a dominant entity — may be favored for their market competitiveness.
+In either case, smaller countries often fear the rise of neighboring superpowers with similar geographic and resource advantages.
 Corollary 1
-Assume the Matthew effect. If $g_{j}>g_{k}$, then the third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
-Thus, providing additional welfare to competitive economies is another significant force driving trade flows across the network, alongside comparative advantage.
-However, trade wars and protectionism are more likely due to competitive advantages rather than comparative advantages, which benefit all trade partners.
-Accordingly, an economic superpower would exert all efforts to maintain its status when facing challenges from emerging competitors; preserving its advantage in $\pi$ prevents the domino effect, which is contrary to the Matthew effect.
-As Copeland (2014) states, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$.
-An emerging superpower’s aspiration to assume its rightful place would encounter significant challenges.
-World War I and II were intense conflicts between existing and emerging superpowers, leading to the establishment of new world orders.
+Assume the Matthew effect. If $g_{j}$ is significantly greater than $g_{k}$, then third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
+Thus, the additional welfare granted to competitive economies becomes a major force driving trade flows, alongside comparative advantage.
+However, trade wars and protectionism are more likely to stem from competitive advantages, which benefit only select partners, rather than comparative advantages, which benefit all.
+Accordingly, an economic superpower will strive to preserve its status when challenged by emerging competitors.
+Maintaining its advantage in $\pi$ helps prevent a reverse Matthew effect.
+As Copeland (2014) notes: “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, the incentive is more proportional to $\pi_{i}$ than to $g_{i}$.
+Emerging superpowers seeking their rightful place face formidable challenges.
+World Wars I and II were intense confrontations between established and rising powers, ultimately reshaping the global order.
 3 Bilateral Trade War for Competitiveness
-Anticipating a significant positive change to $\pi_{i}$, country $i$ considers initiating a trade war against its trade partner $j$.
-Gaming on the matrix $P$, country $i$’s first move could alter the element $P_{ji}$ — the portion of $j$’s production exported to $i$.
-For instance, raising tariffs or decreasing import quotas on $j$ reduces imports from $j$ and decreases $P_{ji}$.
-Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$.
-For simplicity, we assume the bilateral conflict involves only the two counterparties in $P$ without directly affecting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$.
-However, changes in $P_{ji}$ or $P_{ij}$ eventually impact global value chains in multiple steps, indirectly involving all third parties.
-This simplification may not reflect the complexity of the 2018 China-USA trade war, where the Biden administration urged allies to ban Chinese companies from purchasing advanced chips and chip-making equipment (New York Times, 2022).
-In retaliation, country $j$ might decrease $P_{ij}$ or $P_{ji}$ sooner or later.
-Without retaliation, the war ends at inception, and country $i$ wins.
-No reaction or unconditional surrender could avoid further damages but might also invite additional sanctions from $i$ if $j$ shows weakness.
-Overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ may also severely harm $\pi_{j}$.
-For example, after the Trump administration imposed tariffs and trade barriers on China in January 2018, China responded in April 2018 by imposing tariffs on 128 products imported from the USA (Washington Post, 2018).
-We introduce a matrix $\Lambda$ for retaliation actions.
-In retaliation for a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$.
-Similarly, in retaliation for $i$’s change $\Delta P_{ij}$ in $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$.
-Thus, $\lambda_{ij}=1/\lambda_{ji}$, in the veil of any first-mover advantage.
-We assume $\lambda_{ji}>0$ to align actions from both countries.
-Ignoring temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact.
-We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
-Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; thus, the reduced exports will be consumed domestically, and the modified $P$ remains a stochastic matrix.
-The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services.
-Generally, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks.
-In the following exposition, we first assume the coefficient is an exogenous or predetermined constant in the potential trade war.
-Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts.
-The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries.
-In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
-We introduce a few notations for the next five theorems and their corollaries.
-Let $I_{n}$ be the $n\times n$ identity matrix and let $\pi_{{}_{-i}}$ be the transpose of $\pi$ with $\pi_{i}$ removed.
-The square matrix $Z_{i}$ is the transpose of $P$ with its $i$th row and $i$th column removed.
-Also, the column vector $\alpha_{i}$ extracts the $i$th row from $P$ and then drops its $i$th element.
-As usual, $\vec{1}_{n}$ and $\vec{0}_{n}$ are the $n\times 1$ column vectors with all ones and zeros, respectively.
-For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has one for its $j$th element and zeros elsewhere.
-Finally, the vector $\gamma_{ji}$ takes the values of $e_{j}$ and then removes its $i$th element.
-So, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
+Anticipating a significant positive shift in its competitiveness $\pi_{i}$, country $i$ may consider initiating a trade war against its trade partner $j$.
+By strategically manipulating the trade matrix $P$, country $i$’s initial move could alter the element $P_{ji}$ — the share of country $j$’s production exported to country $i$.
+For example, increasing tariffs or reducing import quotas on goods from country $j$ would lower imports from $j$, thereby decreasing $P_{ji}$.
+Country $i$ might also restrict certain exports to $j$, directly reducing $P_{ij}$.
+For simplicity, we assume the bilateral conflict affects only the two counterparties in matrix $P$, without directly impacting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$.
+However, changes in $P_{ji}$ or $P_{ij}$ can propagate through global value chains, indirectly involving third parties over multiple steps.
+This simplification does not fully capture the complexity of events such as the 2018 China–USA trade war, during which the Biden administration encouraged allies to prohibit Chinese firms from acquiring advanced chips and chip-making equipment (New York Times, 2022).
+3.1 Reciprocal Retaliations
+In response, country $j$ may retaliate by reducing $P_{ij}$ or $P_{ji}$.
+For instance, following the Trump administration’s imposition of tariffs and trade barriers on China in January 2018, China retaliated in April 2018 by levying tariffs on 128 U.S.
+products (Washington Post, 2018a).
+If country $j$ does not retaliate, the conflict ends at its inception, and country $i$ emerges victorious.
+A lack of response or unconditional surrender may prevent further damage but could also invite additional sanctions from country $i$ if country $j$ appears weak.
+Conversely, an overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ could also severely harm $\pi_{j}$.
+To formalize retaliation, we introduce a matrix $\Lambda$ representing retaliation actions.
+In response to a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ adjusts $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$.
+Similarly, in retaliation for a change $\Delta P_{ij}$, country $j$ modifies $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$.
+Thus, $\lambda_{ij}=1/\lambda_{ji}$, assuming no first-mover advantage.
+We assume $\lambda_{ji}>0$ to ensure aligned actions from both countries.
+Ignoring temporally lagged reactions on the long-run effects $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ yields equivalent impacts.
+Letting $\lambda_{ii}=1$, we compile all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, for all $i,j\in\mathcal{N}$.
+To preserve the unit sum in each row of $P$, we deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$.
+The reduced exports are then consumed domestically, ensuring that the modified $P$ remains a stochastic matrix.
+The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both countries, including political considerations, anti-dumping measures, trade deficits, and the nature of traded commodities or services.
+In practice, there is no universal formula for every contingency, and negotiations may involve multiple rounds of bilateral discussions.
+In the subsequent section, we assume $\lambda_{ji}$ is an exogenous or predetermined constant in a potential trade war.
+Section 4 explores a generic yet simple bargaining solution for $\lambda_{ji}$ across most of the n(n1)/2n(n-1)/2 bilateral conflicts.
+The solution aims not to eliminate conflicts but to mitigate them and fairly distribute the associated costs, serving the diverse interests of the involved countries.
+In a mixed cooperative and noncooperative setting, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ may be of limited interest to either party.
+We introduce several notations for the next five theorems and their corollaries.
+Let $I_{n}$ denote the $n\times n$ identity matrix and $\pi_{{}_{-i}}$ represent the transpose of $\pi$ with its $i$th element removed.
+Define $Z_{i}$ as the transpose of $P$ with its $i$th row and column removed.
+The column vector $\alpha_{i}$ is obtained by extracting the $i$th row from $P$ and removing its $i$th element.
+Vectors $\vec{1}_{n}$ and $\vec{0}_{n}$ are $n\times 1$ column vectors of all ones and all zeros, respectively.
+For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has a $1$ in its $j$th position and zeros elsewhere.
+Finally, the vector $\gamma_{ji}$ is derived from $e_{j}$ by removing its $i$th element.
+Thus, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
 Theorem 2
-With the above setting and notations, for any $j\neq i$,
+Under the above setting and notations, for any $j\neq i$:
 $\hskip 8.5359pt\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}},$ (9)
@@
 $\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (11)
-Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain:
-$\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
-Using Theorem 2, we have the following Corollary 2, already implied in Eq.
-(6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$.
+Using the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also have:
+$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\qquad\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\qquad\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}=\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.$
+Theorem 2 implies Corollary 2, a variant of Eq.
+(6) — $\pi_{i}$ remains unchanged when the authority flow from $i$ to $j$ equals that from $j$ to $i$.
 However, Eq.
-(6) applies no assumptions in Theorem 2.
+(6) does not rely on the assumptions in Theorem 2.
 Corollary 2
-When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
-Additionally, since
-$(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
-has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3.
-In particular, when $\lambda_{ji}=0$ in Eqs.
-(9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$.
-Eq.
-(10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
+When $\lambda_{ji}=\pi_{j}/\pi_{i}$ (or equivalently, $\lambda_{ij}=\pi_{i}/\pi_{j}$), $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}=\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}=\vec{0}_{n}^{\top}$.
+Additionally, since $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ has all non-negative elements, the derivatives $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3.
+In a trade friction (i.e., $\mathrm{d}P_{ji}<0$), $j$’s overreaction with $\lambda_{ji}>\pi_{j}/\pi_{i}$ benefits $\pi_{i}$ but deteriorates $\pi_{j}$, according to Eqs.
+(9)–(10).
 Corollary 3
 For any $i\neq j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
-By varying $j$ in Eq.
+3.2 Identification of Competitors and Collaborators
+By varying country $j$ in Eq.
 (9), we can identify country $i$’s trade collaborators and competitors.
 In Eq.
-(9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$.
-If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively.
-Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration.
-Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks.
-When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$.
-If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change in $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the sign of $\mathrm{d}P_{ji}$.
-In search of the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$.
-The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
+(9), the derivative $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has the opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$.
+Accordingly, if $\lambda_{ji}>\pi_{j}/\pi_{i}$, then country $j$ becomes a potential target for conflict with country $i$ because a small negative change in $P_{ji}$ would result in a positive change in $\pi_{i}$.
+Conversely, if $\lambda_{ji}<\pi_{j}/\pi_{i}$, then country $j$ is a potential candidate for deeper collaboration.
+In this case, when $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, a larger $\lambda_{ji}$ slows the growth of $\pi_{i}$ and accelerates the decline of $\pi_{j}$.
+Finally, if country $j$ maintains $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any small change in $P_{ji}$ will have negligible effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the direction of $\mathrm{d}P_{ji}$.
+To facilitate cross-country comparisons, we express changes in terms of percentage variation in both $\pi_{i}$ and $P_{ji}$.
+The limit of the percentage change of $\pi_{i}$ with respect to that in $P_{ji}$ is given by:
 $\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}=\frac{\mathrm{d}\pi_{i}/\pi_{i}}{\mathrm{d}P_{ji}/P_{ji}}=\frac{P_{ji}}{\pi_{i}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}.$ (12)
-[…]
-Furthermore, country $i$ would choose:
+When Eq.
+(12) exceeds a positive threshold $\theta$ (e.g., $\theta=.1$) or falls below $\theta$, the derivative in Eq.
+(9) is considered significantly positive or negative, respectively.
+For instance, if Eq.
+(12) equals .2.2, then a $1\%$ increase in $P_{ji}$ would approximately lead to a $2\%$ increase in $\pi_{i}$.
+Therefore, a significantly negative (or positive) derivative in Eq.
+(9) serves as a necessary — but not sufficient — condition for country $i$ to initiate a trade war (or form an economic alliance) with country $j$, even if Eq.
+(10) remains insignificant.
+Furthermore, country $i$ would select
 $\mathop{\mathrm{argmax}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
-as $i$’s best trade partner and:
+as its best trade partner and
 $\mathop{\mathrm{argmin}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
-as $i$’s worst competitor.
+as its worst competitor.
 In Eqs.
-(13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$.
-However, $i$ is not his best partner’s best partner nor his worst competitor’s worst competitor, according to Corollary 3.
-Thus, countries $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched.
-For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
+(13) and (14), $\pi_{i}$ experiences the largest increase or decrease, respectively, in percentage terms for each percentage rise in $P_{ji}$.
+However, according to Corollary 3, country $i$ is not necessarily its best partner’s best partner, nor its worst competitor’s worst competitor.
+This implies that international trade dynamics are not akin to sports-like competitions or zero-sum games — unless the countries involved are evenly matched.
+For example, during the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, while China largely adopted a defensive posture.
 Eq.
-(11) describes the side effects on all other countries due to country $i$’s change in $P_{ji}$.
-With the perturbation in $P_{ji}$, some countries gain, while others lose their competitiveness; the gain or loss may be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$.
-Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not of the same magnitude.
-One movement could be significant, while the other remains insignificant; it is also possible that both movements are insignificant.
-Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
-The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$.
-The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$.
-However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero.
-In practice, however, $i$ is unlikely to exercise these extreme actions because its gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq.
-(4), which could be negligible to $i$.
-Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq.
+(11) describes the spillover effects on other countries resulting from country $i$’s change in $P_{ji}$.
+When $P_{ji}$ is perturbed, some countries gain competitiveness while others lose it; these gains or losses can be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$.
+Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, their magnitude differ: one change may be significant while the other remains negligible, or both may be minor.
+Third countries share the difference between these movements unevenly, and their individual or collective impacts can be substantial.
+These spillover effects may influence the determination of $\lambda_{ji}$ when country $i$ is significantly more powerful than country $j$.
+The conflict appears resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$.
+However, if this condition instead maximizes $\pi_{j}$, then, in theory, $i$ could aggressively counteract.
+In practice, such extreme actions are unlikely because $i$’s gain from $j$ is capped by $\pi_{j}P_{ji}$ (as shown in Eq.
+(4)), which may be negligible for $i$.
+Moreover, $\pi_{j}P_{ji}$ could be smaller than the losses incurred by third parties, as measured by Eq.
 (11).
-So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$.
-It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$.
-For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages.
-For centuries, many norms and standards have been established to tame unilateral trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
+Excessively non-cooperative behavior could provoke backlash from these countries, potentially leading $j$ to form alliances and pursue collective bargaining against $i$.
+Alternatively, $j$ might adopt objectives beyond maximizing $\pi_{j}$, such as avoiding direct confrontation with $i$.
+For instance, it could deepen economic engagement with $i$ by developing complementary industries and leveraging its comparative advantages.
+Historically, norms and standards have evolved to restrain unilateral trade dominance; today, the World Trade Organization (WTO) plays a central role in regulating trade rules among nations.
 4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
 When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$.
-Among these factors, trade deficits are paramount.
+Among these factors, trade deficits are particularly significant.
 This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.
 4.1 Trade Balance
-For a given period, a trade surplus for country $i$ with country $j$ implies a trade deficit of the same magnitude for $j$ with $i$, and vice versa.
-A trade surplus generates job opportunities, reduces unemployment, and expands the economy of scale.
-It also enhances creditworthiness by enabling the country to pay down foreign debt.
-A surplus can result from competitive advantages in tradable goods and services.
-Opinions on trade deficits vary, though large deficits can undermine economic sustainability.
-In the short term, trade deficits help avoid shortages and mitigate issues like inflation and poverty.
+For any given period, a trade surplus for country $i$ with country $j$ implies an equivalent trade deficit for $j$ with $i$, and vice versa.
+A trade surplus creates job opportunities, reduces unemployment, and expands economies of scale.
+It also enhances creditworthiness by enabling debt repayment.
+Surpluses often result from comparative and competitive advantages in tradable goods and services.
+Opinions on trade deficits vary.
+While large deficits can undermine economic sustainability, in the short term, they may help avoid shortages and mitigate issues like inflation and poverty.
 The implications of trade deficits also depend on their impact on national security and how they are financed.
-No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis will not endure.
-There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances, $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3.
-For country $i$ to have a trade surplus with country $j$ due to the change $\Delta P_{ij}$, $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$.
+No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis is unsustainable.
+There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances: $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3.
+For country $i$ to have a trade surplus with country $j$ due to a change $\Delta P_{ij}$, $i$ must export more to $j$ than it imports, i.e., $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$.
 Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$.
-To sustain trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$.
-Excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$.
+To sustain trade peace under $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$ — excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$.
 Therefore, $\lambda_{ji}$ should be neither too large nor too small.
-However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$.
-In this case, both $i$ and $j$ lose exports, though $i$ loses fewer.
-This could adversely affect future production, so most countries avoid mutually damaging stalemates.
-Generally, $\lambda_{ji}$ is determined by both players, but the sign and size of $\Delta P_{ji}$ are often decided by one, especially if it has overwhelming power.
+Conversely, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$.
+In this case, both $i$ and $j$ lose exports, though $i$ loses less.
+Such outcomes can harm future production, so most countries avoid mutually damaging stalemates.
+At a cumulative level, $g_{i}P_{ij}$ and $g_{j}P_{ji}$ represent exports from $i$ to $j$ and from $j$ to $i$, respectively.
+Thus, $g_{j}/g_{i}=P_{ij}/P_{ji}$ results in zero trade surplus and deficit over the accumulation period.
+During this period, $g_{i}$, $g_{j}$, $P_{ij}$, and $P_{ji}$ all grow from zero, so $\Delta P_{ji}=P_{ji}-0$, $\Delta P_{ij}=P_{ij}-0$, and $\lambda_{ji}=\dfrac{\Delta P_{ij}}{\Delta P_{ji}}=\dfrac{P_{ij}}{P_{ji}}=\dfrac{g_{j}}{g_{i}}$.
 Theorem 3
-$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
-To ease trade tension, we have studied two ways to determine $\lambda_{ji}$.
-On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $\Delta P_{ji}$.
-On the other hand, if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates any future trade deficits.
-Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for long-term rivalry due to $\pi$’s involvement in all countries and industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and alleviate related problems such as unemployment and national debt.
-According to Pozsar (2022), for example, the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis.
-Nevertheless, when China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort.
-In this example, the focal point is competitiveness rather than trade balance.
-Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is an appropriate choice for $\lambda_{ji}$.
-We examine a scenario where a superpower country $i$ faces imminent challenges from an emerging economy $j$.
-According to the Matthew effect, it is likely that $\pi_{j}/\pi_{i}<g_{j}/g_{i}$.
-Based on Theorem 4, country $i$ cannot simultaneously have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$.
-This situation occurs only when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages.
-Figure II illustrates this trilemma, where up to two policy objectives are achievable for country $i$.
-For instance, if country $i$ adopts a position on edge $b$, it can overturn comparative advantage and earn a trade surplus but sacrifices competitiveness unless it absorbs competitiveness from other countries.
-In the early 2000s, bilateral exports between China and the USA increased as a percentage of their respective production.
+$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains zero net trade deficit and surplus between countries $i$ and $j$ at the instantaneous or cumulative level, respectively.
+To ease trade tensions, we consider two approaches for determining $\lambda_{ji}$: if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $P_{ji}$; if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates future trade deficits.
+In general, $\lambda_{ji}=\pi_{j}/\pi_{i}$ applies to long-term rivalry because $\pi$ reflects competitiveness across all countries and industries, whereas $\lambda_{ji}=g_{j}/g_{i}$ can immediately resolve trade deficits and alleviate related issues such as unemployment and national debt.
+For example, according to Pozsar (2022), the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting inexpensive goods overseas, while the USA lost much of its manufacturing base and accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis.
+When China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block these moves and urged its allies to join the effort.
+In this case, the focal point was competitiveness rather than trade balance, suggesting a prolonged conflict where $\pi_{j}/\pi_{i}$ is the appropriate choice for $\lambda_{ji}$.
+We examine a scenario in which a dominant superpower, country $i$, faces imminent challenges from an emerging economy, country $j$.
+According to the Matthew effect, it is likely that $\frac{\pi_{j}}{\pi_{i}}<\frac{g_{j}}{g_{i}}$.
+Based on Theorem 4, country $i$ cannot simultaneously achieve three objectives in its bilateral trade with country $j$: increasing its competitiveness $\pi_{i}$; maintaining a trade surplus; expanding exports to country $j$.
+This constraint creates what we call the Impossibility Trilemma.
+The trilemma implies that if country $i$ attempts to grow exports to country $j$ while increasing competitiveness, it cannot also sustain a trade surplus.
+This situation occurs when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages.
+Theorem 4 (Impossibility Trilemma)
+Given $\frac{\pi_{j}}{\pi_{i}}<\frac{g_{j}}{g_{i}}$, country $i$ cannot simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; a trade surplus; positive export growth (i.e., $\Delta P_{ij}>0$).
+Figure II illustrates this trilemma: at most two of the three policy objectives can be achieved by country $i$.
+For example, if country $i$ adopts a position along edge $b$, it can overturn comparative advantage and secure a trade surplus, but only by sacrificing competitiveness — unless it offsets this by absorbing competitiveness from other countries.
+A historical example is the early 2000s, when bilateral exports between China and the USA grew as a share of their respective production.
 According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
-Theorem 4 (Impossibility Trilemma)
-Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
-Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
-Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned and do not completely contradict those of country $j$.
-In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas country $j$’s interests could lie at another point.
-A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries benefit from escalating economies of scale and diminishing marginal costs of production.
-According to the proof of Theorem 4, country $j$ could simultaneously achieve all three objectives listed in the theorem.
-This seems unfair to country $i$, but it is conditional on the Matthew effect, which is somewhat biased toward $\pi_{i}$.
-Therefore, Theorem 4 serves as a redress to the effect.
+Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$
+Country $i$’s interests in increasing competitiveness, maintaining a trade surplus, and achieving export growth are not perfectly aligned, and they do not completely contradict those of country $j$.
+In Figure II, country $i$’s optimal policy position lies somewhere within the triangle, while country $j$’s interests may lie elsewhere.
+A mutually beneficial outcome occurs when $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$, as both countries gain from economies of scale and declining marginal production costs.
+According to the proof of Theorem 4, country $j$ could, under certain conditions, achieve all three objectives simultaneously.
+While this may seem unfair to country $i$, it is conditional on the Matthew effect, which inherently favors $\pi_{i}$.
+Thus, Theorem 4 serves as a corrective measure to this bias.
 4.2 Nash Bargaining Solutions
-This subsection examines a compromise solution between trade balance and competitiveness.
-In the trade system, some countries prioritize competitiveness while others aim to improve trade balance, either by reducing deficits or increasing surpluses.
-Their preferences influence the choice of $\lambda_{ji}$, which maintains $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively.
-If countries are indifferent between competitiveness and trade balance, the midpoint appears to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$.
-However, this midpoint solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$.
-Alternatively, when countries $i$ and $j$ have opposing priorities between enhancing competitiveness and improving trade balance, the Nash bargaining solution (1950) can be employed to find a cooperative solution that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while adhering to the condition $\lambda_{ji}\lambda_{ij}=1$.
-Let
+This subsection explores a compromise between trade balance and competitiveness.
+In international trade, some countries prioritize competitiveness, while others focus on improving trade balance — either by reducing deficits or increasing surpluses.
+These preferences influence the choice of $\lambda_{ji}$, which can maintain either competitiveness $\pi_{i}$ or trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ji}=g_{j}/g_{i}$, respectively.
+If countries are indifferent between competitiveness and trade balance, a natural midpoint might seem like a solution: $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$.
+However, this midpoint does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$.
+When countries $i$ and $j$ have opposing priorities — one favoring competitiveness and the other trade balance — the Nash bargaining solution (1950) provides a cooperative approach that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while ensuring $\lambda_{ji}\lambda_{ij}=1$.
+Define:
 $p\overset{\mathrm{def}}{=\joinrel=}\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}\quad\mathrm{and}\quad q\overset{\mathrm{def}}{=\joinrel=}\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}$ (15)
-which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. By Corollary 2 and Theorem 3, without bargaining, $\lambda_{ji}$ can secure $p$ and $q$, and $\lambda_{ij}$ can secure $1/p$ and $1/q$ because:
+which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. Clearly,
 $\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}\quad\mathrm{and}\quad\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
-We consider two scenarios when countries $i$ and $j$ have opposite preferences on $p$ and $q$.
-In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$.
+Without bargaining, according to Corollary 2 and Theorem 3, $\lambda_{ji}$ can secure $p$ or $q$, and $\lambda_{ij}$ can secure $1/p$ or $1/q$.
+These represent their respective status quo points if no agreement is reached.
+Two scenarios arise when countries $i$ and $j$ have conflicting preferences.
+In the first scenario, country $j$ prefers $q$ over $p$, while $i$ prefers $1/p$ over $1/q$.
 However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$.
@@
 (15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other.
-In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$.
-Theorem 5 addresses the bargaining problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design.
-The solution is consistent across both preference scenarios, thus it remains effective when either country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance.
-The solution remains unchanged under any positive affine transformation of their utility functions, according to a property of the bargaining solution.
-By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$.
-Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation bridges the gap between these two ratios.
+In the second scenario, country $j$ prefers $p$ over $q$, while $i$ prefers $1/q$ over $1/p$.
+A preference over these alternatives defines a von Neumann-Morgenstern utility function for the bargaining problem.
+Theorem 5 resolves these bargaining problems.
+The solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design.
+It remains consistent across both scenarios and under any positive affine transformation of utility functions (a property of Nash bargaining).
+By taking the square root of $\pi_{j}/\pi_{i}$, the solution also mitigates the Matthew effect in $\pi$, bridging the gap between these $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$.
 Theorem 5
-When countries $i$ and $j$ have opposite preferences on $p$ and $q$, the Nash bargaining solution is:
+When countries $i$ and $j$ have opposing preferences on $p$ and $q$, the Nash bargaining solution is:
 $\lambda_{ji}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}\hskip 14.22636pt\mathrm{and}\hskip 14.22636pt\lambda_{ij}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}$
@@
 Figure III illustrates the bargaining solution.
-In the left plot (first scenario), $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$ while $i$ moves from bottom to top along the same curve.
-When they meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, the area of the rectangle cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and the status quo point $(p,1/q)$ is maximized.
-The right plot (second scenario) illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve.
-They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
-Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences.
-The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ rely on the choice of status quo points.
-The effective bargaining set is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$ because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second, respectively) scenario.
-If both players intensify or weaken their status quo positions symmetrically, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$.
-Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$.
-However, the bargaining solution remains unchanged, according to Corollary 4.
-Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
+In the first scenario (left plot), country $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$, while country $i$ moves from bottom to top along the same curve.
+They meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, which maximizes the area of the rectangle formed by this point and the status quo point $(p,1/q)$.
+The second scenario (right plot) shows a similar process where where $j$ moves from $q$ to $p$ and $i$ moves from $1/p$ to $1/q$ along the curve.
+Again, they meet at $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, maximizing the area of the rectangle defined by this point and $(q,1/p)$.
+Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences
+The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ depend on the choice of status quo points.
+The effective bargaining set (or retaliation corridor) is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$, because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second) scenario.
+These corridors help prevent hostile overreactions and unintended consequences, such as third-party blowback.
+However, they can be adjusted if both players symmetrically strengthen or weaken their status quo positions.
+If we scale the two status quo points by a positive number $c\leq\sqrt{q/p}$, the effective bargaining set for $\lambda_{ji}$ expands from $[p,q]$ to $[cp,q/c]$, and for $\lambda_{ij}$ from $[1/q,1/p]$ to $[c/q,1/(cp)]$.
+Despite this scaling, the bargaining solution remains unchanged (Corollary 4), while the new effective bargaining sets approach $(0,\infty)$ as $c\to 0$.
 Corollary 4
-For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$, the bargaining solution remains $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
-Additionally, as Corollary 5 states, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$, as implied in Corollary 2 and Theorem 3.
-In summary, when countries $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
-$\lambda_{ji}=\left\{\begin{array}[]{ll}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},&\mathrm{otherwise.}\end{array}\right.$ (16)
-Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb.
+For any positive $c\leq\sqrt{q/p}$, replacing $p$ with $cp$ and $q$ with $q/c$ does not change the bargaining solution: $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
+Additionally, as stated in Corollary 5, the ratios ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$ when countries $i$ and $j$ directly compete for competitiveness and trade balance, respectively.
+The solution $\lambda_{ji}=\lambda_{ji}^{*}$ or $\pi_{j}/\pi_{i}$ — but not $g_{j}/g_{i}$ — appears endogenous to the derivative $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$, whose zero value provides a status quo point for the Nash solution.
+Nevertheless, the solution remains unchanged regardless of the derivative.
 Corollary 5
-${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
-Finally, country $j$ should avoid direct confrontation with a significantly larger-sized economy $i$, according to Corollary 6.
-This was likely China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy.
-Consequently, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, 1968).
+Both ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
+However, direct confrontation may be an inappropriate option when $g_{i}$ and $g_{j}$ differ significantly, according to Corollary 6.
+This aligns with China’s strategy with the USA from 1980 to 2010, when Deng Xiaoping advocated “keeping a low profile and taking no lead” (e.g., Yan, 2014).
+Historically, an even distribution of economic, political, and military capabilities between rival countries tends to increase the likelihood of war, whereas peace is best preserved when there is an imbalance of national capabilities (e.g., Organski, 1968).
+Therefore, Corollary 6 suggests using $\lambda_{ji}=\lambda_{ji}^{*}$ as the rule of thumb, unless their sizes of economies are comparable — in this case, $\frac{\pi_{j}}{\pi_{i}}$ and $\frac{g_{j}}{g_{i}}$ are also viable options for $\lambda_{ji}$.
 Corollary 6
-Country $j$ should avoid homogeneous preferences with a significantly larger-sized country $i$.
+During globalization era when $\Delta P_{ji}>0$, country $j$ should strategically avoid homogeneous preferences with a significantly larger country $i$.
 4.3 National Bargaining Power
-In deriving the solution $\lambda_{ji}^{*}$, both countries have a symmetric role and identical prior bargaining power, as per the Nash bargaining axiom.
-However, countries $i$ and $j$ typically possess asymmetric prior bargaining power.
-For instance, we could maximize $\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}$ or $\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ given some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013).
-Without prior knowledge about $\zeta_{ij}$, there is no justification to favor one country over another, resulting in $\lambda^{*}_{ji}$.
-In a real bargaining context, however, the parameter $\zeta_{ij}$ depends on numerous characteristics between these two, as well as other, countries.
-Non-economic factors such as military power, population, culture, tradition, legislation, and location, along with alternative optimal criteria, can significantly compromise the symmetry axiom and lead to bargaining solutions that differ substantially from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
-We introduce a posterior bargaining power that integrates both the symmetric prior and the data, capitalizing on the results in Theorem 5.
-To extend the bilateral Nash solution in Theorem 5 to multilateral negotiations, we define
+In deriving the solution $\lambda_{ji}^{*}$, both countries are assumed to have symmetric roles and identical prior bargaining power, consistent with the Nash bargaining axiom.
+In practice, however, countries $i$ and $j$ typically exhibit asymmetric prior bargaining power.
+For example, one could maximize
+$\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}\quad\mathrm{or}\quad\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ (16)
+for some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013).
+In real-world bargaining, however, $\zeta_{ij}$ depends on numerous factors, including non-economic elements such as military strength, population, culture, legal systems, and geographic location.
+These factors, along with alternative optimal criteria, can significantly violate the symmetry axiom, resulting in bargaining outcomes that deviate from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
+Without prior knowledge of $\zeta_{ij}$, there is no justification to favor one country over another, leading to the symmetric solution $\lambda^{*}_{ji}$, which may provide information about bargaining power.
+To address this, we introduce a posterior bargaining power measure that combines the symmetric prior with empirical data, building on Theorem 5.
+Extending the bilateral Nash solution to multilateral negotiations, we define:
 $\beta_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{\sqrt{\pi_{i}g_{i}}}{\sum\limits_{j\in\mathcal{N}}\sqrt{\pi_{j}g_{j}}}$
-as country $i$’s national bargaining power in the trade system.
-It is normalized so that all bargaining power sums to one, allowing $\beta_{i}$ to be compared over time.
-This measure aids multinational negotiations.
-For instance, when allocating votes in forming or reforming a multinational alliance, using this power can avoid bilateral dialogues and discussions.
-Despite factors other than production and competitiveness, bilateral trade agreements remain more prevalent than multilateral ones; Yilmazkuday and Yilmazkuday (2014) explored some reasons.
-The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ constructs a linear ordering for all countries and maintains a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$.
-If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$.
-Specifically, $i$’s bargaining power over $k$ equals the ratio of $i$’s power over $j$ and $j$’s power over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
-Also, we can recover it directly from the relation $\lambda^{*}_{ik}=\beta_{i}/\beta_{k}$.
-Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$.
-Thus, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$.
-From the production aspect, $\beta_{i}$ declines when companies in country $i$ rely on inputs from partner countries.
-The more companies depend on partners, the less power the country has to coerce concessions from them.
-From the consumption aspect, bargaining power increases when the country consumes more imported goods and services, thereby influencing the production of its counterparts.
-Between the determinants, $g_{i}$ represents the country’s breadth of economy, while $\pi_{i}$ predominantly reflects the depth of the economy, as the matrix $P$ in Eq.
-(3) already normalizes production volumes.
-However, the positive correlation between them implies a risk of a domino effect, where a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, leading to further reduction in $g_{i}$.
-Consequently, $\beta_{i}$ would shrink more than the initial drop in $g_{i}$.
+where $\beta_{i}$ represents country $i$’s national bargaining power in the trade system.
+This measure is normalized so that the sum of all bargaining powers equals one, enabling comparisons over time.
+It is particularly useful in multilateral context, such as allocating vote rights in forming or reforming international alliances, as it avoids the need for bilateral negotiations.
+Despite this, bilateral trade agreements remain more common than multilateral ones (e.g., Yilmazkuday and Yilmazkuday, 2014).
+The set $\{\beta_{i}\}_{i=1}^{n}$ establishes a linear ordering of countries and preserves the transitivity property of the Nash solution $\lambda_{ij}^{*}$.
+If $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $\beta_{i}>\beta_{k}$.
+Specifically, country $i$’s bargaining power relative to $k$ equals:
+$\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}=\beta_{i}/\beta_{k},\quad\forall j\in\mathcal{N}.$
+Two equally weighted and highly correlated factors determine $\beta_{i}$: production volume $g_{i}$ and competitiveness strength $\pi_{i}$.
+Consequently, $\beta_{i}$ inherits some of the “controversial” properties of $\pi_{i}$.
+From the production aspect, $\beta_{i}$ decreases when domestic firms rely heavily on imported inputs, reducing the country’s leverage in negotiations.
+From the consumption perspective, bargaining power rises when the country imports more goods and services, influencing foreign production.
+However, the positive correlation between the breadth $g_{i}$ and depth $\pi_{i}$ of the economy introduces a risk of a reverse Matthew effect: a decline in $g_{i}$ can trigger a reduction in $\pi_{i}$, which in turn further depresses $g_{i}$.
+As a result, $\beta_{i}$ may shrink more than the initial drop in $g_{i}$.
 5 Economic Globalization
 Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures.
-In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services.
-A recent example is China’s entry into the WTO in 2001, while the UK’s Brexit in 2020 was a clear rejection of globalization.
-Conversely, protectionist measures could raise tariffs on imported goods, set import quotas for other countries, and enact stricter government regulations on imports.
-For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports from other countries (Washington Post, 2018).
-Globalization creates winners and losers within a nation but does not automatically compensate the losers, thereby fueling income inequality and political divisions.
-Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers.
-This section questions whether the country should retreat from global economic integration without directly attacking any specific country.
-The objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms.
-We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
-Country $i$’s globalization action triggers adjustments in the entire matrix $P$.
-Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, with a small change $\Delta P_{ii}$.
-If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism.
-As the action has no specific target countries, we assume that whenever $P$ changes the $i$th column, the other columns change proportionally to ensure each row of $P$ still has a unit sum.
-For example, the change $\Delta P_{ji}$ causes proportional changes $\Delta P_{ji}\sigma_{ji}$ to the $j$th row where:
+In economic terms, it describes the growing interdependence among countries, driven by trade in goods and services.
+A notable example is China’s accession to the WTO in 2001, while the UK’s Brexit in 2020 represented a clear rejection of globalization.
+Conversely, protectionist policies can raise tariffs on imports, impose quotas, and enforce stricter regulations.
+For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports (Washington Post, 2018b).
+Globalization creates both winners and losers within nations, but it does not automatically compensate those who lose, often fueling income inequality and political divisions.
+Therefore, a more inclusive approach is needed — one that not only maximizes the gains of winners but also mitigates negative impacts on others.
+The benefits of globalization are also unevenly distributed across countries.
+This section explores whether a nation should retreat from global economic integration without directly targeting any specific country.
+The goal is to enhance national competitiveness rather than simply maximizing the aggregate utility of trade firms.
+Our analysis focuses on actions that influence the global economy as a whole, while disregarding bilateral trade balances.
+5.1 Optimal Level of Globalization
+When country $i$ takes an action to adjust its level of globalization, it triggers changes across the entire matrix $P$.
+The country’s first move affects the diagonal element $P_{ii}$ with a small change $\Delta P_{ii}$.
+If globalization increases, $P_{ii}$ decreases (i.e., $\Delta P_{ii}<0$); if protectionism rises, $P_{ii}$ increases (i.e., $\Delta P_{ii}>0$).
+Since the action does not target specific countries, we assume that whenever column $i$ of $P$ changes, the other columns adjust proportionally to maintain the unit sum of each row.
+For example, a change $\Delta P_{ji}$ induces proportional changes $\Delta P_{ji}\sigma_{ji}$ within row $j$ where:
 $\sigma_{ji}\overset{\mathrm{def}}{=\joinrel=}\frac{1}{1-P_{ji}}\left(-P_{j1},\cdots,-P_{j,i-1},1-P_{ji},-P_{j,i+1},\cdots,-P_{jn}\right).$
-Consequently, the change $\Delta P_{ii}$ provokes adjustments $\Delta P$ in three levels of fallout.
+Consequently, a change $\Delta P_{ii}$ in $P$ produces three levels of adjustments $\Delta P$ in $P$.
 First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$.
-Secondly, the $i$th column in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}$, due to reprisal reactions, where $\odot$ denotes the element-wise product or Hadamard product and $\Lambda_{i}$ is the $i$th row of $\Lambda$.
-Lastly, the $i$th column of $\Delta P$ triggers changes in all other columns.
+Secondly, the $i$th column in $\Delta P$ is $\left[\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}\right]^{\top}$, due to reprisal reactions, where $\odot$ denotes the Hadamard (element-wise) product and $\Lambda_{i}$ is the $i$th row of $\Lambda$.
+Lastly, the change in column $i$ propagates to all other columns.
 We define:
 $M\overset{\mathrm{def}}{=\joinrel=}\mathrm{diag}\left(\sigma_{ii}\odot\Lambda_{i}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right]$
-where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with the diagonal vector $\sigma_{ii}\odot\Lambda_{i}$.
-Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$, and thereby $\Delta P=\Delta P_{ii}M$.
-So, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M$ as $\Delta P_{ii}\to 0$.
-We let $M_{i}$ be $M$ with its $i$th column dropped.
-Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
+where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with $\sigma_{ii}\odot\Lambda_{i}$ on its diagonal.
+Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$.
+Therefore:
+$\Delta P=\Delta P_{ii}M,\quad\mathrm{and}\ \mathrm{as}\ \Delta P_{ii}\to 0,\quad\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M.$
+Finally, let $M_{i}$ denote $M$ with its $i$th column dropped. Theorem 6 provides the derivative of $\pi$ with respect to $P_{ii}$.
 Theorem 6
-With the above setting and notations, we have:
+Under the above setting and notations, we have:
 $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (17)
@@
 $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\left(I_{n-1}-Z_{i}\right)^{-1}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (18)
-The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$.
-The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimal level because $P_{ii}$ cannot be $1$ or 00.
-However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may be positive or negative at this optimal level, allowing country $j$ to capitalize on this by making a corresponding negative or positive change in $P_{ij}$, thereby affecting $P_{ii}$.
-To disincentive other countries from making simultaneous movements, we set $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}$ to be a zero vector.
-This can be tested by checking if $\pi M_{i}$ is a zero vector, according to Corollary 7.
+The decision rule on globalization could be as follows: country $i$ should pursue further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively.
+If this derivative is insignificantly different from zero, country $i$ may instead seek trade war targets or trade partnerships as outlined in Sections 3 and 4.
+In addition, a positive (negative) sign of $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ indicates the country is over-globalized (under-globalized, respectively).
+The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$, thereby rebalancing $P$ by $\Lambda$.
+The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimum level because $P_{ii}$ cannot be $1$ or 0.
+However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may remain positive or negative at this optimal level, allowing country $j$ to exploit this by adjusting $P_{ij}$, thereby influencing $P_{ii}$.
+To deter simultaneous strategic moves by other countries, we impose $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
+This condition holds if and only if $\pi M_{i}=\vec{0}_{n-1}$, as stated in Corollary 7.
 Corollary 7
@@
 Given $\Lambda$, Eqs.
-(17) and (18) for all $i\in\mathcal{N}$ define a dynamical system describing the instantaneous change of $\pi$ with respect to the diagonal of $P$.
-There are $n$ independent variables in the system since the proportions among off-diagonal elements remain unchanged.
-Due to the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on the diagonal.
-For example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% simultaneously; this succession continues until the matrix $P$ converges.
-Alternatively, we study the properties of the solution to the first-order conditions.
-If the solution exists, it satisfies a new counterbalance equilibrium in Eq.
+(17) and (18) for all $i\in\mathcal{N}$ define a dynamical system that describes the instantaneous change of $\pi$ with respect to the diagonal elements of $P$.
+The system has $n$ independent variables since the proportions among off-diagonal elements remain unchanged.
+Due to its complexity, numerical algorithms can approximate a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive small adjustments on the diagonal elements.
+For example, over-globalized countries may increase $P_{ii}$ by 1% at each step, while under-globalized countries reduce $P_{ii}$ by 1% simultaneously.
+This iterative process continues until both the matrix $P$ and the vector $\pi$ converge.
+Unlike the convergences discussed in Lau et al.
+(2022), the diagonal elements do not necessarily converges to the same value.
+Alternatively, we examine the properties of the solution to the first-order conditions.
+If such a solution exists, it satisfies a new counterbalance equilibrium given in Eq.
 (19), as stated in Corollary 8.
 Eq.
-(19) places $n-1$ identities to the solution, requiring only one more condition to uniquely identify it.
-Also, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits some desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, resulting in bilateral authority or trade balance, according to Eq.
-(6) or Theorem 3, respectively.
-At the stable solution, any tiny perturbation on the diagonal of $P$ would have a nonsignificant effect on $\pi$, discouraging countries from making small movement.
-Yet, this solution is optimal for some countries only, and not every country maximizes its competitiveness with this solution.
+(19) imposes n1n-1 identities to the solution, requiring only one additional condition to uniquely identify it.
+Moreover, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits certain desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, which correspond to balanced bilateral authority or trade volumes according to Eq.
+(6) and Theorem 3, respectively.
+Notably, Eq.
+(19) reduces to Eq.
+(3) when $\lambda_{ji}=P_{ij}/P_{ji}$, indicating no globalization or deglobalization is likely occur in the absence of trade imbalance.
+Corollary 8
+If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies:
+$\pi\left[\Lambda\odot P\right]^{\top}=\pi.$ (19)
+Eq. (19) holds if all $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all $\lambda_{ij}=P_{ji}/P_{ij}$.
+At the stable solution, any small perturbation to the diagonal of $P$ has a insignificant effect on $\pi$, discouraging countries from making minor adjustment.
+However, this solution is optimal for some countries only and does not necessarily maximize competitiveness for all.
 Additionally, the derivatives in Eqs.
-(17) and (18) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$.
-Therefore, seeking a far-reaching solution to the first-order conditions might be of limited practical use.
-Nevertheless, Theorem 6 provides each country with the right direction to steer.
-Corollary 8
-If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
-$\pi\left[\Lambda\odot P\right]^{\top}=\pi.$ (19)
-Eq. (19) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
-The decision rule on globalization could be as follows: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively.
-Otherwise, when $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4.
-In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies.
-Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and prevent major internal trade disputes.
-Location and cultural proximity often foster the formation of a PTA or RTA.
-Besides, Eq.
-(18) measures the side effects of $i$’s protectionism on the rest of the world, helping select countries for a PTA or RTA.
-In general, some would benefit from it, whereas others would not.
-Finally, Theorem 6 implies a mixed globalization and protectionism strategy, helping $i$ to decide which PTA or RTA to join.
-Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively.
-Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq.
-(17) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$.
-Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$, and the positive and negative contributions are partially offset.
-Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as:
+(17) and (18) rely on the exogeneity assumption of $\Lambda$ and proportional rebalancing in $M$.
+Therefore, pursuing a comprehensive solution to the first-order conditions may have limited practical use.
+Nevertheless, Theorem 6 provides each country with guidance on the appropriate direction to move.
+5.2 Mixed Globalization–Protectionism Strategy
+In addition to globalization and trade wars, country $i$ can also form or join preferential trade agreements (PTAs) or regional trade agreements (RTAs) with other economies.
+Ideally, within a PTA or RTA, the ratios $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should remain balanced to facilitate consensus and prevent significant internal trade disputes.
+Geographic and cultural proximity often encourage the formation of such agreements.
+Moreover, Eq.
+(18) measures the external impact of country $i$’s protectionist policies on the rest of the world, aiding in the selection of suitable partners for PTAs or RTAs.
+While some countries benefit from these, others may not.
+Theorem 6 suggests a mixed globalization–protectionism strategy, guiding country $i$ in deciding which PTAs or RTAs to join.
+To implement this, we divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on the sign of their components in the vector $\pi M_{i}$:
+  • 1.
+    $\mathcal{N}_{i}^{+}$: countries with positive components;
+  • 2.
+    $\mathcal{N}_{i}^{-}$: countries with negative components.
+Since $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ is a non-negative vector, we rewrite Eq.
+(17) as: $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$.
+Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, its contribution to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is negative or positive, respectively, partially offsetting each other.
+Given $\Delta P_{ii}$, the approximate change in $\pi_{i}$ is:
 $\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}_{i}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}_{i}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (20)
-In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting.
-This strategy is better than a pure globalization or protectionism strategy because the right-hand side of Eq.
+Under the mixed strategy, country $i$ increases protectionism toward countries in $\mathcal{N}_{i}^{-}$ and deepens globalization with countries in $\mathcal{N}_{i}^{+}$, thereby avoiding the partial offsetting effect.
+This approach outperforms a purely globalization or purely protectionism strategy because the right-hand side of Eq.
 (20) is less than or equal to:
-$-\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
-Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$, respectively.
-If most country members of a PTA or RTA are in $\mathcal{N}_{i}^{+}$, then country $i$ should consider participating in the trade agreement.
-However, country $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply that country $i\in\mathcal{N}_{j}^{+}$, possibly causing additional negotiations if $j$ is already in the PTA or RTA.
+$\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
+Therefore, if most members of a PTA or RTA belong to $\mathcal{N}_{i}^{+}$, then country $i$ should consider joining that agreement.
+However, note that $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply $i\in\mathcal{N}_{j}^{+}$, which may require additional negotiations if country $j$ is already part of the PTA or RTA.
 6 Empirical Study
-In this empirical study, we apply our theoretical results to real-world trade data spanning from 2000 to 2019.
-This analysis provides quantitative evidence to support our previous arguments, making the theoretical insights more tangible and relevant for policymakers.
-The empirical analysis reveals trends in national competitiveness and bargaining power over time, highlighting how different countries have gained or lost.
-This provides a historical perspective on the impacts of globalization and trade wars.
-Between 2000 and 2019, the world economy underwent significant transformations and challenges, albeit with varying degrees of success across different regions.
-This period witnessed a substantial increase in global trade volumes, driven by rapid technological advancements, the expansion of international trade agreements, and the rise of emerging markets.
-The integration of China into the WTO in 2001 was a pivotal moment, leading to a surge in global trade activities.
-The rise of emerging markets, particularly China and India, played a crucial role in global economic growth but also led to a trade war between China and the USA.
-The 2008 financial crisis and the subsequent European debt crisis were major turning points, causing widespread economic downturns and leading to substantial policy interventions by governments and central banks worldwide.
-As trade volumes increased with globalization, nations demanded more of the world’s reserve currency, primarily through trade surpluses.
-Consequently, the reserve currency appreciates, decreasing the currency issuer’s exports and increasing trade deficits.
-The issuer may have three options: encouraging local currency settlement in international transactions or reducing its currency’s usage as the world’s reserve currency; ignoring the trade deficits and focusing on national competitiveness while still increasing exports; or rolling back globalization worldwide.
-For the USA, reshoring its manufacturing sectors would enhance its competitiveness and grow exports; however, trade deficit may still be an issue according to the impossibility trilemma.
-For China, significant growth is necessary to shrink the competitiveness gap with the USA, according to the Matthew effect in $\pi$.
+In this empirical study, we apply our theoretical framework to real-world trade data spanning from 2000 to 2019.
+This analysis offers quantitative evidence that supports our earlier arguments, making the theoretical insights more concrete and relevant for policymakers.
+The empirical findings reveals evolving trends in national competitiveness and bargaining power, illustrating how various countries have gained or lost ground over time.
+These insights provide a historical lens through which to examine the impacts of globalization and trade conflicts.
+Between 2000 and 2019, the global economy experienced profound transformations and challenges, with varying outcomes across regions.
+This period saw a marked increase in global trade volumes, fueled by rapid technological advancements, the expansion of international trade agreements, and the emergence of new economic powers.
+A pivotal moment was China’s accession to the WTO in 2001, which triggered a surge in global trade activity.
+The rise of emerging markets — particularly China and India — was instrumental in driving global economic growth.
+However, it also contributed to escalating trade tensions, most notably between China and the USA.
+The 2008 financial crisis and the subsequent European debt crisis were major inflection points, causing widespread economic downturns and prompting significant policy interventions by governments and central banks worldwide.
+As globalization intensified and trade volumes expanded, nations increasingly demanded the world’s reserve currency, primarily through trade surpluses.
+This demand led to appreciation of the reserve currency, which in turn reduced the currency issuer’s export competitiveness and widened trade deficits.
+Thus, reversing globalization trends could be a viable option for the issuer.
+For the USA, reshoring manufacturing sectors could reduce imports and deficits.
+For China, sustained economic growth is essential to narrow the competitiveness gap with the USA, in line with the Matthew effect in $\pi$.
 This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$.
-The bilateral trade data are sourced from the United Nations ComTrade database (2021).
-We use import and export data for non-consumer goods from the World Bank’s WITS Database (2021).
-For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018.
-Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA).
-Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively.
-We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations.
-The results are in the following tables and figures.
-Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
-Figure IV illustrates the competitiveness trends from 2000 to 2019.
-During this period, China and Russia experienced significant gains, while Canada remained relatively stable, and the other six countries saw declines.
-Notably, the USA’s competitiveness grew steadily until 2003, followed by a sharp decline until 2011, and a gradual recovery thereafter, resulting in an overall decrease of 12.7%.
-China’s competitiveness surged during the two crises but plateaued otherwise.
-Canada and Russia demonstrated resilience due to rising commodity prices during transformative years.
-The four Western European countries were significantly impacted by the debt crisis, with their competitiveness continuing to fall until 2015.
-Germany, focused on manufacturing, began its decline in 2000, while the other three countries started in 2008.
-Japan’s competitiveness nearly halved in the first eight years but stabilized between 2007 and 2014, suffering a domino effect from 2000 to 2007 as China and South Korea challenged its automobile and electric industries.
-Beyond these nine countries, others collectively improved their competitiveness by 22.7%, as depicted in Figure VI.
-Many Southern Asian, Southeast Asian, and Middle East countries, including India, Indonesia, Iraq, United Arab Emirates, and Vietnam, tripled their competitiveness.
-The Matthew effect is clearly demonstrated.
-For instance, the USA and China’s production ratio was $1.344$ in 2018, but their competitiveness ratio was $3.070$.
-This effect is further evidenced by Figure VI, which shows a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$ after controlling $g_{i}$.
-The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
-Figure V: The sum of $\pi_{i}$ for all other countries (2000-2019).
-Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling $g_{i}$ (2000-2019).
+Bilateral trade data are sourced from the United Nations ComTrade database (2021), while import and export data for non-consumer goods come from the World Bank’s WITS Database (2021).
+For comparison, we focus on two benchmark years: 2000, prior to China’s accession to the WTO in 2001, and 2017, just before the onset of the China-U.S.
+trade war in 2018.
+Although the calculations include all economies in the dataset, we present results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA), identified by their ISO-3 codes: CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA.
+The computations use the default coefficients, $\lambda_{ji}=\lambda_{ji}^{*}$, and the results are summarized in the following tables and figures.
+Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019)
+Figure IV illustrates competitiveness trends from 2000 to 2019.
+During this period, China and Russia recorded significant gains, while Canada remained relatively stable, and the other six countries experienced declines.
+Notably, the USA’s competitiveness rose steadily until 2003, then dropped sharply through 2011, followed by a gradual recovery — resulting in an overall decline of 12.7%.
+China’s competitiveness surged during the two major crises but otherwise plateaued.
+Canada and Russia demonstrated resilience, supported by rising commodity prices.
+The four Western European countries were severely affected by the debt crisis, with competitiveness falling until 2015.
+Germany’s decline began in 2000, while the other three started in 2008.
+Japan’s competitiveness nearly halved in the first eight years, then stabilized between 2007 and 2014.
+From 2000 to 2007, Japan faced a reverse Matthew effect as China and South Korea challenged its automobile and electric industries.
+Beyond these nine countries, other economies collectively improved competitiveness by 22.7%, as shown in Figure VI.
+Many countries in Southern Asian, Southeast Asian, and the Middle East — including India, Indonesia, Iraq, the United Arab Emirates, and Vietnam — tripled their competitiveness.
+The Matthew effect is clearly evident in the data.
+For example, in 2018, the production ratio between the USA and China was $1.344$, but their competitiveness ratio reached $3.070$.
+This disparity underscores the cumulative advantage phenomenon.
+Further evidence in presented in Figure VI, which illustrates a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$, after controlling for $g_{i}$.
+However, the partial correlation declined from .757.757 to .527.527 over the study period, coinciding with the transition of the trade system from a less multipolar to a more multipolar structure.
+Figure V: Sum of $\pi_{i}$ for all other countries (2000-2019)
+Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling for $g_{i}$ (2000-2019)
 Table I: Country $j$’s direct contribution to $\pi_{i}$ for 2017 and 2019 (in percentage)
@@
 USA $\frac{1.3878}{1.4971}$ $\frac{0.8107}{1.0322}$ $\frac{0.4527}{0.4749}$ $\frac{0.2580}{0.2740}$ $\frac{0.5935}{0.5643}$ $\frac{0.1733}{0.1802}$ $\frac{0.3514}{0.3897}$ $\frac{0.0368}{0.0464}$ $\frac{92.142}{91.369}$ $\frac{.293880}{.272142}$
-* The numerators are for 2019 and the denominators for 2017.
-A straightforward implication of Eq.
-(4) is country $j$’s direct contribution to country $i$’s competitiveness in percentage, i.e., $\pi_{j}P_{ji}/\pi_{i}\times 100\%$.
-Table I lists these contributions for 2017 and 2019.
-According to the diagonal numbers, the USA was the most self-sufficient in feeding its competitiveness, with a slight increase from 2017 to 2019.
-The ninth column also shows that the USA was the most significant contributor to Canada, China, Germany, Japan, and the UK in 2019.
-For the same year, China was the most significant contributor to Russia’s competitiveness and the second largest to those of Canada, Japan, and the USA.
-France, Germany, and Italy were highly reliant on each other’s contributions, and their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia.
-In value, the USA contributed .14492.025965=.003763.14492.025965.003763.14492*.025965=.003763.14492 ∗ .025965 = .003763 to Canada while Canada contributed .013878.293880=.004078.013878.293880.004078.013878*.293880=.004078.013878 ∗ .293880 = .004078 to the USA in 2019.
-Comparing the years before and after the beginning of the China-USA trade war, the USA’s contributions to China dropped by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)11.5849.0939052.1950.0865961-(1.5849*.093905)/(2.1950*.086596)1 - ( 1.5849 ∗ .093905 ) / ( 2.1950 ∗ .086596 ), while China’s contribution to the USA reduced by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1.8107.293881.0322.2721421-(.8107*.29388)/(1.0322*.272142)1 - ( .8107 ∗ .29388 ) / ( 1.0322 ∗ .272142 ).
-Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
-Figure VII illustrates the national bargaining power from 2000 to 2019, displaying patterns similar to those in Figure IV, albeit smoother.
-China’s bargaining power steadily increased due to rapid production growth, even during its $\pi_{i}$’s troughs in 2005 and 2015.
-Notably, China’s bargaining power consistently surpassed its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness, indicating $\beta$’s mitigation of the Matthew effect in $\pi$.
-Russia’s bargaining power sharply declined following the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
-Figure VIII illustrates the comparative bargaining power of the USA against other countries.
-The bargaining power of four Western European countries experienced a slight increase following the establishment of the Eurozone in 1999, but saw a modest decline post-2008, culminating in an average loss of 15.7% of their bargaining power to the USA.
-Conversely, China and Russia significantly enhanced their relative power multiple times within the first 15 years, maintaining a steady level thereafter.
-Japan, however, suffered a substantial reduction, losing approximately 50% of its relative bargaining power.
-Canada’s relative bargaining strength surged during crises but swiftly reverted to its original level, attributable to its strong alignment with the USA.
-Figure VIII: Bargaining power versus the USA (2000-2019).
+* The numerators correspond to 2019 values, and the denominators to 2017.
+A direct implication of Eq.
+(4) is that country $j$’s contribution to country $i$’s competitiveness can be expressed as a percentage: $\pi_{j}P_{ji}/\pi_{i}\times 100\%$.
+Table I reports these contributions for 2017 and 2019.
+The diagonal values indicate that the USA was the most self-sufficient in sustaining its competitiveness, with a slight increase from 2017 to 2019.
+The ninth column further shows that, in 2019, the USA was the largest contributor to Canada, China, Germany, Japan, and the UK.
+In the same year, China was the primary contributor to Russia’s competitiveness and the second-largest contributor to Canada, Japan, and the USA.
+France, Germany, and Italy exhibited strong mutual reliance, while their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia.
+Bilateral break-even points in Eq.
+(6), in general, do not exist.
+For example, in 2019, the USA contributed .14492.025965=.003763.14492*.025965=.003763 to Canada, whereas Canada contributed .013878.293880=.004078.013878*.293880=.004078 to the USA.
+Comparing the periods before and after the onset of the China-USA trade war, the USA’s contributions to China decreased by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)1-(1.5849*.093905)/(2.1950*.086596), while China’s contribution to the USA fell by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1-(.8107*.29388)/(1.0322*.272142).
+Figure VII: Bargaining power $\beta_{i}$ (2000-2019)
+Figure VII illustrates national bargaining power from 2000 to 2019, showing patterns similar to Figure IV but smoother.
+China’s bargaining power rose steadily due to rapid production growth, even during troughs in its $\pi_{i}$ in 2005 and 2015.
+Notably, China’s bargaining power consistently exceeded its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness — evidence of $\beta$’s role in mitigating the Matthew effect in $\pi$.
+Russia’s bargaining power declined sharply after the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
+Figure VIII compares the USA’s bargaining power against the other countries.
+Western European countries saw a slight increase after the Eurozone’s formation in 1999, followed by a modest decline post-2008, resulting in an average 15.7% loss of bargaining power relative to the USA.
+Conversely, China and Russia significantly strengthened their relative power during the first 15 years, then stabilized.
+Japan, however, experienced a substantial decline, losing about 50% of its relative bargaining power.
+Canada’s bargaining strength spiked during crises but quickly reverted to its original level, reflecting its strong alignment with the USA.
+Figure VIII: Bargaining power versus the USA (2000-2019)
 Table II: $1,000\times\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}$ for years 2000 and 2017
@@
 * The numerators are for 2017 and the denominators for 2000.
-Table II presents the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated from Eqs.
+Table II reports the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated using Eqs.
 (9) and (12).
-For instance, a 1% reduction in imports from China by the USA in 2017 would have resulted in an increase of .04889% in the USA’s competitiveness.
-Although the USA may have had reasons to initiate trade wars against all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row.
-As evidenced by the last column, China and Japan would have been the two largest beneficiaries had they further collaborated with the USA in 2017 and 2000, respectively.
-However, a comparison of the numerators and denominators in the column reveals a considerable reduction in benefits from 2000 to 2017.
-Additionally, initiating trade friction from any of these eight countries against the USA would result in a loss for the initiating country.
-A trade war against Russia would be nearly fruitless due to the small numerators and denominators in the eighth column.
-Furthermore, the numbers symmetric with the diagonal exhibit opposite signs, corroborating Corollary 3.
+For example, a 1% reduction in U.S.
+imports from China in 2017 would have increased U.S.
+competitiveness by .04889%.
+Although the USA may have had reasons to initiate trade disputes with all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row.
+The last column shows that China and Japan would have been the two largest beneficiaries had they strengthened collaboration with the U.S.
+in 2017 and 2000, respectively.
+However, comparing the numerators and denominators across years reveals a significant decline in potential benefits from 2000 to 2017.
+Moreover, any of these eight countries initiating trade friction against the USA would incur losses.
+A trade war with Russia, in particular, would be almost fruitless due to the small values in the eighth column.
+Finally, the numbers symmetric to the diagonal exhibit opposite signs, confirming Corollary 3.
 Table III: $1,000\times\frac{\mathrm{d}\log\pi_{j}}{\mathrm{d}\log P_{ii}}$ for years 2000 and 2017
@@
 * The numerators are for 2017 and the denominators are for 2000.
-Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eqs.
+Table III presents the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated using Eqs.
 (17), (18), and (12).
-From the diagonal, globalization significantly benefited China, Germany, Italy, Japan, and Russia.
-It was advantageous for Canada in 2000 but detrimental in 2017.
-The effects on France were insignificant due to small numerator and denominator.
-The USA and the UK would be anti-globalization advocates, explaining recent events such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit.
-Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans from manufacturing offshoring demanded more inexpensive goods from China, uplifting trade deficits and national debts but benefiting from low inflation and affordable merchandises.
-Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$.
-Russia was the largest beneficiary of globalization in 2000.
-The side effects on third parties are positioned off the diagonal.
-For instance, the USA’s further globalization benefited all eight other countries, thereby making its protectionist policies likely to face significant resistance from its partners.
-China emerged as the largest beneficiary of the USA’s globalization in both years.
-Conversely, Canada experienced the least impact from the USA’s protectionism in 2017; its further globalization would have marginally benefited the USA, as indicated by the first row.
-The last column suggests that further globalization from seven other countries would adversely affect the USA; Japan and China topped the list in 2000 and 2017, respectively.
-Their adverse impact decreased by 70.2% and increased by 380%, respectively, over the 17 years.
-Additionally, the second row elucidates the USA and its Western allies’ objection to China’s BRI-like projects, which intensified by $4.74$ times from 2000 to 2017, according to the ratios of their numerators and denominators.
-We assess the externalities on the third parties due to the USA’s trade war against China in 2018 and 2019.
+The diagonal values indicates that globalization significantly benefited China, Germany, Italy, Japan, and Russia.
+Canada gained in 2000 but faced negative effects in 2017.
+For France, the impact was negligible due to small numerator and denominator values.
+The USA and the UK appeared as anti-globalization advocates, which aligned with recent developments such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit.
+While business interest groups profited from vast international markets and low-cost raw materials, many low-income Americans — affected by manufacturing offshoring — demanded cheaper goods from China.
+This dynamic increased trade deficits and national debts but also delivered low inflation and affordable consumer products.
+China emerged as the largest beneficiary of globalization in 2017, with a 2.324% increase in $\pi_{i}$ for every 1% decrease in $P_{ii}$.
+This advantage underscores the positive momentum behind its Belt and Road Initiative (BRI).
+In contrast, Russia was the largest winner in 2000.
+Off-diagonal values reveal spillover effects on third parties.
+For example, further globalization by the USA benefited all eight other countries, suggesting that U.S.
+protectionist policies would likely face resistance from its partners.
+China consistently gained the most from U.S.
+globalization in both years.
+Conversely, Canada experienced minimal impact from U.S.
+protectionism in 2017, and its own globalization would have only marginally benefited the USA, as indicated in the first row.
+The last column shows that additional globalization from seven other countries would negatively affect the USA, with Japan and China exerting the greatest adverse impact in 2000 and 2017, respectively.
+Over 17 years, Japan’s negative effect decreased by 70.2%, while China’s increased by 380%.
+Furthermore, the second row highlights the growing opposition of the USA and its Western allies to China’s BRI-like projects, which intensified by a factor of $4.74$ between 2000 and 2017.
+We assess the externalities imposed on third parties by the China-USA trade war during 2018 and 2019.
 Table IV is derived from Eqs.
 (9), (11), and (12).
-When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively.
-Canada was a free rider, and the other six countries were slightly negatively affected in 2018 and 2019.
-Consequently, these countries would likely refrain from taking formal action to prevent the trade war.
-Furthermore, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict.
-The second-best allies would be the UK for the USA and Japan for China, respectively.
-From 2018 to 2019, the effects were slightly mitigated for all these countries.
-Likely recognizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term conflict.
-However, the cumulative effects would be significant as the conflict is expected to continue for years.
-Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
+When the USA reduced its imports from China by 1% in 2018 and 2019, China’s competitiveness $\pi_{j}$ suffered significant losses of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ experienced non-significant gains of .0540% and .0460%, respectively.
+Canada benefited as a free rider, whereas the other six countries were slightly negatively affected in both years.
+From 2018 to 2019, the negative effects were slightly mitigated for all countries.
+Consequently, these countries were unlikely to take immediate action to prevent the dispute.
+However, cumulative spillover effects are expected to be significant as the conflict is likely to persist for years.
+Among potential allies, Canada and Russia would have been the best partners for the USA and China, respectively.
+The second-best allies would have been the UK for the USA and Japan for China.
+Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ — Side Effects of the China-USA Trade War on Third Parties
 Year CAN DEU FRA GBR ITA JPN RUS CHN USA
@@
 2019 -21.815 3.223 2.133 .039 5.324 12.392 14.894 93.744 -45.954
-* Country $i$ is for the USA, $j$ for China, and $k$ for others.
-Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
+* Country $i$ refers to the USA, $j$ to China, and $k$ to others.
+Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of the conflict in 2018:
 $\frac{\pi_{j}}{\pi_{i}}=\frac{.092618}{.284327}<\frac{g_{j}}{g_{i}}=\frac{17494.79}{23511.17}.$
-For a positive $\mathrm{d}P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V).
-Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was:
-[…]
-for 2018.
-For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$.
-According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164.
-Therefore, if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China.
-Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$.
-To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
-7 Discussions
-Economic globalization and trade wars have emerged as significant public interests due to global competition among nations, large-scale restructuring of labor markets, and the international distribution of added value.
-This study posits that import and export data already reflect these issues, and that competitive advantages, rather than comparative advantages, are the primary drivers of trade wars and de-globalization.
-Consequently, we aim to extract a nation’s competitive strength and bargaining power from the data, identify optimal strategies for addressing trade conflicts and cooperation, and determine the extent to which a nation should pursue globalization.
-The methodologies employed, grounded in network and game theories, focus on contemporaneously interactive relationships among economies and are devoid of parameters and pre-set utility functions.
-These network game approaches can be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
-Several critical factors warrant further investigation.
-First, counterbalance disequilibrium may arise when trades are notably imbalanced.
-This imbalance can manifest in various ways.
-For instance, when country $i$ uses borrowed money to consume country $j$’s exports, country $i$’s consumption still exerts influence over country $j$’s production.
-However, this influence originates not only from country $i$, but also from the money lender, who possesses some control over $i$’s consumption.
-If the lender happens to be country $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ is underestimated, potentially leading to a debt trap for country $i$.
-Secondly, unlike an economic superpower, an emerging economy may prioritize the growth of exports over competitive advantages.
-Indeed, its economic growth must significantly surpass that of advanced economies to bridge the competitiveness gap, in accordance with $\pi$’s Matthew effect.
-Thus, maximizing competitive advantages may be a low priority for an emerging economy.
-For example, during China’s primitive capital accumulation in the 1980s and 1990s, earning additional US dollars was essential to finance its industrial modernization.
-The time dimension, absent from the current research, is necessary to formulate growth strategies.
-Thirdly, the facets of economic globalization encompassing capital, information, and technology have yet to be fully integrated.
-These elements interact with goods and services, but their data are not as easily accessible.
-Additionally, several data issues may undermine the empirical study’s results.
-Notably, there are numerous missing values in the service sectors, and discrepancies often arise between country $i$’s reported imports from country $j$ and country $j$’s reported exports to country $i$.
-Furthermore, the integration of regional economies warrants further examination.
-Theorem 2 could be employed to identify suitable candidates for integration.
-However, the underlying assumption may need modification.
-For instance, if country $i$ is a member of multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries as used in Theorem 6.
-Moreover, additional variables could be incorporated into the demand-sided counterbalance Eq.
-(3), particularly those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$.
-For example, incorporating labor loss into $g_{i}$.
-Any factors directly influencing country $i$’s production could be included in the $i$th row of $P$.
-Although consumption may still predominantly drive the fractions, not all production is driven by consumption.
-The framework can be extended in several directions.
-Firstly, amid a bilateral trade war and anti-globalization sentiments, a country may target specific groups of partners without directly impacting others.
-For instance, as illustrated in Table II, the USA could simultaneously increase tariffs on imports from China, Germany, and Japan.
-Secondly, while the dynamical system described by Eqs.
-(9)–(11) or Eqs.
-(17)–(18) assumes a fixed $P$, introducing a temporal dimension to $P$ could provide valuable insights.
-Incorporating a time horizon would facilitate the examination of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and the improvement of trade deficits.
-Additionally, the temporal dimension could enhance the model’s forecast capabilities.
-Intertemporal delays often occur when production takes place in one year but consumption in another.
-If production spans two years, the previous year’s imports of intermediate goods could be included in the current year’s completed production.
+For a positive $\Delta P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V).
+Thus, an economic conflict with a negative $\Delta P_{ji}$ was possible, and in this lose-lose scenario, the Nash bargaining solution was:
+$\lambda_{ji}^{*}=\sqrt{\frac{.092618\times 17494.79}{.284327\times 23511.17}}=.49233.$
+For each dollar decrease in exports from China to the USA ($\Delta P_{ji}=-\mathdollar 1/g_{j}$), China would have adjusted its imports from the USA by: $\Delta P_{ij}=\lambda_{ji}^{*}\Delta P_{ji}$ of the USA’s production, which is $49233*(-1/g_{j})*g_{i}=-\mathdollar.66164.$ Therefore, if this resolution had been applied, the USA would have earned a surplus of .33836=(.66164)(1).33836=(-.66164)-(-1) for each dollar of reduced imports from China.
+Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\Delta P_{ji}<0$.
+To maintain competitive advantage and reduce national debt by earning trade surpluses, the USA, by the impossibility trilemma, had to reduce its imports and exports — i.e., $\Delta P_{ji}<0$ — through policies such as tariffs and manufacturing reshoring.
+7 Policy Implications and Discussions
+Economic globalization and trade wars have become pressing public concerns, driven by intensifying global competition, large-scale labor market restructuring, and the international distribution of added value.
+This study argues that import and export data already capture these dynamics and that competitive advantage — rather than comparative advantage — is the primary force behind trade wars and the trend toward deglobalization.
+Our objective is to extract insights into a nation’s competitive strength and bargaining power from trade data, identify optimal strategies for managing trade conflicts and fostering cooperation, and assess the appropriate level of engagement in globalization.
+When applying this policy-oriented framework in real-world trade scenarios, it is essential to explore alternative approaches that adapt and extend the framework to specific national and economic contexts.
+7.1 Practical and Policy Implications
+This paper conceptualizes national competitiveness as a systemic property, rather than a simple reflection of unilateral production metrics or bilateral trade balances.
+Specifically, the metric $\pi$ represents stationary standings among nations, derived from their economic interactions — namely, bilateral trade flows.
+This formulation positions $\pi$ as a network-wide measure of economic strength, accounting for both direct and indirect effects across global value chains.
+Policymakers should monitor $\pi$ and its fluctuations alongside trade balances and production metrics to gain a more comprehensive understanding of national competitiveness.
+7.1.1 Prioritizing Trade Objectives
+When navigating the trade-off between competitiveness and trade balance, two actionable ratios can guide bilateral trade decisions: long-run competitiveness-neutral point $\lambda_{ji}=\dfrac{\pi_{j}}{\pi_{i}}$, useful for managing strategic rivalries without escalation; short-run trade-balance-neutral point $\lambda_{ji}=\dfrac{g_{j}}{g_{i}}$, suitable for addressing immediate concerns such as employment, debt sustainability, or foreign exchange reserves.
+These benchmarks help trade negotiators align policy actions with either long-term strategic goals or short-term economic stability.
+The choice between strategies based on competitive versus comparative advantage should reflect national characteristics.
+Small economies are advised to adopt comparative advantage strategies to promote inclusive growth and mutual benefit, while superpowers may pursue competitive advantage strategies that favor dominant players.
+Most countries will adopt mixed, country-specific objectives.
+For example, in 2020, the USA faced strategic challenges from China alongside massive national debt.
+Yet, competitiveness is often overlooked by incumbent administrations due to its intangible nature and long-term impact.
+Few quantitative metrics like $\pi$ have been used to measure competitiveness growth.
+7.1.2 Recognizing Asymmetries Early
+The convex nature of competitiveness creates reinforcing dynamics — often referred to as the Matthew effect — which makes it difficult for latecomers to catch up unless they grow substantially faster.
+This has critical implications for strategic sequencing, such as determining when to confront or collaborate with emerging rivals.
+Convexity also increases the risk of a Thucydides trap, where instability arises from competition between the two leading superpowers.
+Stability is more likely when a single superpower coexists with several great powers.
+Our empirical studies suggest that such instability may persist due to ongoing competition between China and the USA.
+Consequently, incumbent superpowers should prioritize maintaining their leadership position.
+Meanwhile, small, open economies can leverage agglomeration effects by aligning with high-$\pi_{i}$ hubs and using pricing strategies to offset biases toward large markets.
+Fast-growing challengers aiming to close the competitiveness gap should avoid systematic confrontations with significantly larger economies.
+Another critical asymmetry is the impossibility trilemma, which clarifies trade-offs often obscured in policy debates.
+Large economies should use it to set realistic bilateral goals, recognizing that achieving both a surplus and higher competitiveness may require reducing exports to challengers or broadly curbing imports.
+7.1.3 Competition and Cooperation Corridors
+The decision to compete or collaborate with a specific country depends on the marginal impact of imports and exports on $\pi_{i}$.
+Given limited resources, country $i$ may prioritize its top competitors and collaborators while ignoring others.
+The counterbalance equilibrium provides a formal rationale for mixed trade policies that combine competitive pressure with cooperative incentives.
+This unity of opposites suggests avoiding extreme actions in both retaliations and cooperation.
+For example, controlling rare minerals and advanced technologies could escalate issues to national security levels.
+Trade relationships evolve over time, making timing critical.
+Countries should develop an early-warning dashboard that annually tracks changes in $\pi$, $\beta$, and derivatives in Theorems 2 and 6.
+Thresholds — such as elasticities exceeding $\pm.1\%$ — should trigger interagency reviews if signs flip or magnitudes shift significantly.
+7.1.4 Negotiation leverage and institutional design
+The square-root formulation in $\lambda_{ji}^{*}$ symmetrically balances competitiveness and trade balance.
+It is scale-free and can serve as a default “fair retaliation” coefficient in bilateral trade protocols.
+However, this solution can be strategically manipulated (Corollary 6), particularly if a weaker nation appears to avoid competition with a dominant one, which may overlook this trap.
+The asymmetric Nash solution in Eq.
+(16) also benefits smaller countries more than the non-cooperative solutions in Corollary 5.
+Moreover, when analyzing the relationship between a weaker country $j$ and a dominant country $i$, the Matthew effect indicates that the Nash bargaining solution yields $\lambda_{ji}^{*}>\pi_{j}/\pi_{i}$.
+Consequently, the United States may identify only a limited number of countries as suitable partners for deeper collaboration in its pursuit of enhanced national competitiveness (see Section 3.2).
+Nevertheless, all other economies can benefit from cooperating with the sole superpower, as outlined in Corollary 1 and supported by the empirical study.
+Both the bilateral solution $\lambda_{ji}^{*}$ and the multilateral solution $\beta_{i}$ faces drawbacks due to simplifying assumptions that often fail to capture the complexity of international relations.
+Incorporating unequal bargaining power and third-party influence can improve these solutions.
+For instance, the USA may impose stricter retaliation measures (e.g., tariffs and quotas) on nations running trade surpluses with it.
+This may exclude China, given their comparable production sizes, with or without purchasing power parity.
+7.1.5 Globalization and Supply-Chain strategy
+Before adjusting border measures, country $i$ should assess its sensitivity to changes in its domestic share $P_{ii}$, i.e., $\dfrac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$.
+This indicates whether further globalization or protectionism enhances competitiveness.
+This empirical approach offers a pragmatic alternative to ideological debates over liberation versus reshoring.
+A mixed strategy appears optimal: deepen ties with partners in $\mathcal{N}_{i}^{+}$ and selectively ring-fence against those in $\mathcal{N}_{i}^{-}$.
+In practice, however, both sets are too large.
+Accordingly, country $i$ should consider joining or upgrading PTAs/RTAs rich in $\mathcal{N}_{i}^{+}$, while negotiating carve-outs for areas with high $\mathcal{N}_{i}^{-}$ exposure — such as rules-of-origin thresholds or snap-back clauses.
+With these agreements, the spillover effects in Eqs.
+(11) and (18) require adjustments.
+Due to the 2018 China-USA trade war, fragmentation between two blocs of nations has increased tension in the international trade system.
+The first bloc includes the over-globalized USA and its traditional allies, while the second comprises nations heavily reliant on China’s infrastructure-building capacity and affordable goods.
+Trade between blocs may decrease, while intra-bloc trade may increase.
+Countries outside these blocs may benefit.
+7.1.6 Third-party spillovers and Trade Alliances
+Partners play a pivotal role in the trade network.
+While uniform trade policies rarely benefit all partners, heterogeneous policies can generate gains across countries.
+Several methods — such as Eqs.
+(6) and (12), or the sets $\mathcal{N}^{+}_{i}$ and $\mathcal{N}^{-}_{i}$ — help identify competitors and collaborators.
+Given any partner, country $i$ can enhance its competitiveness by adopting targeted strategies: weakening ties with competitors or strengthening ties with collaborators.
+These strategies adjust existing trade relationships, and before implementing them, a country should anticipate who stands to gain and who might resist — critical for coalition-building and mitigating opposition.
+Gains and losses extend beyond principal actors to third-party countries.
+To manage these dynamics, countries should use Eqs.
+(11) and (18) to measure spillover effects, forecast ally reactions, and design compensatory mechanisms — thereby reducing coalition leakage during disputes.
+Ignoring such side effects can turn collaborators into competitors.
+Partner-specific collaboration can yield substantial returns, but these benefits often diminish over time, underscoring the need to regularly update partner rankings and reassess traditional alliances using current data.
+7.2 Discussions and Extensions
+The methodologies in this study — rooted in network and game theory — emphasize real-time interactions among economies without relying on fixed parameters or predefined utility functions.
+These network game approaches can also be applied to other bilateral or dyadic systems, such as traffic networks, social platforms, and global payment infrastructures, as well as analogous conflicts, including currency and technology wars.
+Several critical issues warrant further exploration.
+First, significant trade imbalances can create counterbalance disequilibrium.
+For example, when country $i$ uses borrowed funds to purchase exports from country $j$, $i$’s consumption still influence $j$’s production.
+However, this influence originates not only from $i$ but also from the lender, who exerts some control over $i$’s consumption.
+If the lender is country $j$ itself, then $\pi_{i}$ may be overestimated and $\pi_{j}$ underestimated, potentially leading to a debt trap for $i$.
+Second, unlike economic superpowers, emerging economies may prioritize export growth over competitive advantages.
+To close the competitiveness gap, their economic growth must significantly outpace that of advanced economies, consistent with the Matthew effect in $\pi$.
+Consequently, maximizing competitive advantages may be a low priority.
+For instance, during China’s early capital accumulation in the 1980s and 1990s, earning additional U.S.
+dollars was crucial to finance industrial modernization, such as high-speed rail technology transfer.
+Incorporating the time dimension — currently absent from this research — is essential for developing effective growth strategies.
+Thirdly, key dimensions of economic globalization — namely capital, information, and technology — remain insufficiently integrated into the analysis.
+Although these elements interact with goods and services, their associated data are comparatively less accessible.
+Additionally, several data limitations may hinder empirical assessments, including missing values in service sectors and frequent discrepancies between country $i$’s reported imports from country $j$ and $j$’s reported exports to $i$.
+The integration of regional economies also warrants deeper investigation.
+Theorem 2 and $\mathcal{N}_{i}^{+}$ could be employed to identify suitable candidates for integration; however, the underlying assumptions may require revision.
+For example, if country $i$ participates in multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries, as assumed in Theorem 6.
+Moreover, additional variables could be incorporated into the demand-side counterbalance, particularly from the supply side — such as labor mitigation — which could refine the fractions in $P$.
+For instance, accounting for labor loss in $g_{i}$ would capture factors directly influencing country $i$’s production.
+These factors could be reflected in the $i$th row of $P$.
+While consumption may still predominantly drive these fractions, not all production is consumption-driven.
+The framework offers several avenues for extension.
+First, in the context of a bilateral trade war and rising anti-globalization sentiments, a country may target specific groups of partners without directly affecting others.
+For example, as shown in Table II, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan.
+Second, while the dynamical systems described by Eqs.
+(9)–(11) or (17)–(18) assume a fixed $P$, introducing a temporal dimension could yield valuable insights.
+Incorporating a time horizon would enable analysis of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and improvements in trade deficits.
+This temporal extension could also enhance the model’s forecasting capabilities.
+Intertemporal delays often arise when production and consumption are distributed across multiple years.
+For instance, if production spans two years, imports of intermediate goods from the previous year could be included in the current year’s completed production.
 Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios.
-An industrial or geopolitical analysis could provide valuable insights into this choice.
-For instance, as Japan and South Korea competed in automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was appropriate for $\lambda_{ji}$.
-In the same year, no industry served as a battleground for China and the USA, making $g_{j}/g_{i}$ a natural choice for $\lambda_{ji}$.
-Moreover, the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the importance of the distance between trade partners and the type of merchandise as crucial determinants of trade flows.
-Consequently, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
-Econometric analyses could also provide additional insights.
-In the linear regression of Eq.
-(8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is greater or less than 00, we may infer that $\pi_{i}$ is over- or under-valued, respectively.
-Including additional regressors, such as debt to GDP ratio, in the regression helps to better understand $\pi$ through explanatory data.
-Finally, anticipating the side effects of country $i$’s trade war against country $j$, country $j$ might seek cooperation with third parties that would substantially suffer from the conflict.
-However, adding more countries to both sides could result in a global trade war, akin to the Cold War between 1947 and 1991.
+Industrial or geopolitical analyses may guide this choice.
+For instance, in 2000, Japan and South Korea were competitors in automobile and electronic exports, making $\pi_{j}/\pi_{i}$ a suitable basis for $\lambda_{ji}$.
+In contrast, no major industry served as a battleground between China and the USA that year, suggesting that $g_{j}/g_{i}$ or $\lambda_{ji}^{*}$ might be more appropriate.
+The literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the significance of distance between trade partners and the type of merchandise as key determinants of trade flows.
+Accordingly, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, thereby partially removing locational advantages from $\pi$.
+Econometric analyses could also offer further insights.
+For example, in the linear regression of Eq.
+(8), a positive or negative estimated residual $\hat{\epsilon}_{i}$ may indicate that $\pi_{i}$ is overvalued or undervalued, respectively.
+Including additional regressors, such as the debt-to-GDP ratio, could enhance the explanatory power of $\pi$.
+Finally, in anticipation of the side effects of a trade war between country $i$ and country $j$, the latter might seek cooperation with third parties that would also be adversely affected.
+Country $i$ could also form an alliance with those who would benefit from the conflict.
+However, expanding the number of countries involved on both sides could escalate the situation into a global trade war, reminiscent of the Cold War era (1947–1991).
 Appendix
@@
 We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$.
-Accordingly, $P$ has three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row.
-Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$.
+Accordingly, $P$ has three other simultaneous changes: $\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row.
+Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; 1-1 at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $\lambda_{ji}$ at $(i,i)$.
 Without loss of generality, we assume $1\leq i<j\leq n$.
@@
 $\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&\vec{0}_{i-1}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&-\lambda_{ji}\pi_{i}+\pi_{j}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&\vec{0}_{j-i-1}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\lambda_{ji}\pi_{i}-\pi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&\vec{0}_{n-j}\end{array}\right].$
-Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the $(i+1)$st row, making the $(i+1)$st row of Eq.
+Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the (i+1)(i+1)st row, making the (i+1)(i+1)st row of Eq.
 (A.5) a zero vector.
-After dropping the $(i+1)$st row in Eq.
+After dropping the (i+1)(i+1)st row in Eq.
 (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
 To make the matrix in Eq. (A.6) lower-triangular, we multiply the following non-singular matrix
-$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.7)
+$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ \vec{0}_{n-1}&I_{n-1}\end{array}\right]$ (A.7)
 to the left side of Eq. (A.6) to get:
@@
 Also, by the second row of Eq. (A.8):
-$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
+$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
 and thus:
@@
 Similarly, after we add all rows (except for the first one) in Eq.
-(A.5) to the $(j+1)$st row, the $(j+1)$st row becomes a zero vector.
-After dropping the $(j+1)$st row in Eq.
+(A.5) to the (j+1)(j+1)st row, the (j+1)(j+1)st row becomes a zero vector.
+After dropping the (j+1)(j+1)st row in Eq.
 (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
 We multiply the following non-singular matrix:
-$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
+$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ \vec{0}_{n-1}&I_{n-1}\end{array}\right]$
 to the left side of Eq. (A.9) to get:
@@
 $\Delta P_{ji}>0$ $\pi_{i}\uparrow$ $\Delta\pi_{i}=0$ $\pi_{i}\downarrow$ $\pi_{i}\downarrow\downarrow$ $\pi_{i}\downarrow\downarrow\downarrow$
----- - - --- - -- zero net $+$
+--- -- - zero net ++
 $\Delta P_{ji}<0$ $\pi_{i}\downarrow$ $\Delta\pi_{i}=0$ $\pi_{i}\uparrow$ $\pi_{i}\uparrow\uparrow$ $\pi_{i}\uparrow\uparrow\uparrow$
-+++++++ + + +++++ + $+$ zero net --
-$+$ and -- for net trade surplus and deficit, respectively.
-Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, $+$, or --.
+++++++ ++++ ++ zero net -
+++ and - for net trade surplus and deficit, respectively.
+Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, ++, or -.
 $\Delta\pi_{j}$ and $\Delta\pi_{i}$ have opposite signs and country $i$ and $j$’s net trade balances have opposite signs.
@@
 A10. Proof of Corollary 6
-Country $j$’s direct competition with $i$ for a higher $\pi_{j}$ results in $\lambda_{ji}=\pi_{j}/\pi_{i}$.
-If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect.
-Consequently, by Eq.
-(10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, regardless of trade surplus or deficit.
-Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq.
-(10).
-Thus, $j$ should also not compete with $i$.
+When country $j$ directly competes with country $i$ to increase $\pi_{j}$, the Nash solution yields $\lambda_{ji}=\pi_{j}/\pi_{i}$.
+In this case, $\pi_{j}$ remains unchanged.
+However, if country $j$ strategically aims to improve its trade balance instead, the Nash solution becomes $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is typically larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect.
+Therefore, by Eq.
+(10), country $j$ achieves its original objective of enhancing $\pi_{j}$, as its exports to $i$ increase — regardless of a trade surplus or deficit.
+Similarly, if country $j$ competes with $i$ solely to improve its trade balance, the Nash solution is $\lambda_{ji}=g_{j}/g_{i}$, which results in zero trade surplus.
+But if country $j$ strategically seeks to enhance $\pi_{j}$, the Nash solution again becomes $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is generally smaller than $g_{j}/g_{i}$ due to the Matthew effect.
+Consequently, an increasing in $P_{ji}$ leads to a positive trade surplus, fulfilling country $j$’s original objective.
 A11. Proof of Theorem 6
@@
 As the last column of Eq.
-(A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the $(i+1)$st row to make the $(i+1)$st row a zero vector.
-After dropping the $(i+1)$st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
+(A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the (i+1)(i+1)st row to make the (i+1)(i+1)st row a zero vector.
+After dropping the (i+1)(i+1)st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
 $\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$ (A.18)
@@
 Also, by the second row of Eq. (A.19):
-$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
+$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
 and thus:
@@
 […]
-$\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that $-(1-P_{ii})\pi M$, i.e.,
+$\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that $(1-P_{ii})\pi M$, i.e.,
 $\left(\pi_{1}\lambda_{i1}P_{i1},\cdots,\pi_{i-1}\lambda_{i,i-1}P_{i,i-1},-\pi_{i}(1-P_{ii}),\pi_{i+1}\lambda_{i,i+1}P_{i,i+1},\cdots,\pi_{n}\lambda_{in}P_{in}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right],$
-is a zero vector except for its $i$th element. Thus, for any $t\neq i$, the zero value in the $t$th element implies that:
-$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}P_{it}.$
+is a zero vector except for its $i$th element.
+For any $t\neq i$, zero value of the $t$th element of the above vector implies that:
+$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\frac{\pi_{i}(1-P_{ii})P_{it}}{1-P_{ii}}=0.$
 We sum the above equations over all $t\neq i$ to get:
-$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}\sum_{t\neq i}P_{it},$
+$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}\sum_{t\neq i}P_{it}=0$
 which leads to:
-$\sum\limits_{j\neq i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.21)
-and:
-$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\odot P\right]_{ji}^{\top}=\pi_{i}.$
+$\sum\limits_{j\in\mathcal{N}}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}.$ (A.21)
 When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\odot P\right]^{\top}=\pi$.